### v1

# and even simpler proofs of quantumness

Abstract
Bell non-locality is a fundamental feature of quantum mechanics whereby measurements performed on “spatially separated” quantum systems can exhibit correlations that cannot be understood as revealing predetermined values. This is a special case of the more general phenomenon of “quantum contextuality”, which says that such correlations can occur even when the measurements are not necessarily on separate quantum systems, but are merely “compatible” (i.e. commuting). Crucially, while any non-local game yields an experiment that demonstrates quantum advantage by leveraging the “spatial separation” of two or more devices (and in fact several such demonstrations have been conducted successfully in recent years), the same is not true for quantum contextuality: finding the contextuality analogue of such an experiment is arguably one of the central open questions in the foundations of quantum mechanics.
In this work, we show that an arbitrary contextuality game can be compiled into an operational “test of contextuality” involving a single quantum device, by only making the assumption that the device is computationally bounded. Our work is inspired by the recent work of Kalai et al. (STOC ’23) that converts any non-local game into a classical test of quantum advantage with a single device. The central idea in their work is to use cryptography to enforce spatial separation within subsystems of a single quantum device. Our work can be seen as using cryptography to enforce “temporal separation”, i.e. to restrict communication between sequential measurements.
Beyond contextuality, we employ our ideas to design a “proof of quantumness” that, to the best of our knowledge, is arguably even simpler than the ones proposed in the literature so far.
1 Introduction
One of the most intriguing features of quantum mechanics is that, in general, observable properties of a quantum system, usually referred to as “observables”, do not seem to hold a precise value until they are measured. In technical jargon, quantum mechanics is not a “local hidden variable theory”. While this feature is now well-understood, Einstein, Podolski, and Rosen, in their paper [EPR35], originally conjectured that a local hidden variable explanation of quantum mechanics should exist. It was only years later that Bell [Bel64], and subsequently Clauser, Horne, Shimony, and Holt (CHSH) [CHSH69], in their seminal works, proposed an operational test, i.e. an experiment, capable of ruling out a local hidden variable explanation of quantum mechanics. More precisely, they showed that there exists an experiment involving measurements on “spatially separated” quantum systems such that the outcomes exhibit correlations that cannot be explained by a local hidden variable theory. Such an experiment, usually referred to as a Bell test or a non-local game, has been performed convincingly multiple times [HBD+15, GVW+15, SMSC+15, LWZ+18, RBG+17, SSK+23]. Crucially, a Bell test rules out a local hidden variable theory under the assumption that the devices involved in the experiment are non-communicating (which is usually enforced through “spatial separation”).
[…]
Is there a way to compile an arbitrary contextuality game into an operational “test of contextuality”?
Beyond demonstrating the presence of genuine quantum behaviour, non-local games can be employed to achieve a much more fine-grained control over the behaviour of untrusted quantum devices: for example, they allow a classical user to verify the correctness of full-fledged quantum computations, by interacting with two non-communicating quantum devices [RUV13, CGJV19]. Of course, any guarantee obtained via non-local games hinges on the physical (and non-falsifiable) assumption that the devices involved are non-communicating. To circumvent the need for this assumption, a lot of the attention in recent years has shifted to the computational setting. This exploration was kick-started by Mahadev’s seminal work [Mah18] showing, via cryptographic techniques, that the verification of quantum computations can be achieved with a single quantum device, under the assumption that the device is computationally bounded. She and her collaborators [BCM+21] later proposed what can be thought of as the analogue of a Bell/CHSH experiment with a single device—they proposed a simple test that an efficient quantum device can pass, but that an efficient classical device cannot. Since then, various works have proposed increasingly efficient “proofs of quantumness” in this setting [BKVV20, ACGH20, KMCVY22, KLVY23, AMMW22, BGKM+23a]. The goal of this line of work is to simplify these tests to the point that they can be implemented on current quantum devices. An experimental realisation of such a proof of quantumness would be a milestone for the field of quantum computation, as it would constitute the first efficiently verifiable demonstration of quantum advantage. Towards this goal, the second question that we consider in this work is:
Can contextuality help realise simpler proofs of quantumness?
1.1 Our results
[…]
Theorem 1 (informal, simplified).
Any contextuality game can be compiled into a single prover, 2-round (i.e. 4-message) protocol such that the following two conditions hold (assuming standard cryptographic assumptions):
  • (Completeness) There is a quantum polynomial-time (QPT) prover that wins with probability at least
    $\frac{1}{2}(1+{\rm valQu})-\mathsf{negl}\,.$
  • (Soundness) Any probabilistic polynomial-time (PPT) prover wins with probability at most
    $\frac{1}{2}(1+{\rm valNC})+\mathsf{negl}\,.$
Here $\mathsf{negl}$ are (possibly different) negligible functions of a security parameter.
The bounds in the statement above are for games with contexts of size two (which subsume “$2$-player” non-local games, see LABEL:Fig:whatcompileswhat). The general bounds are similar but depend on the size of the contexts and are stated later. Notably, the number of messages, even in the general case (which subsumes non-local games with any number of players), remains constant (i.e. four). The cryptographic assumptions we make are the existence of (1) a quantum homomorphic encryption (QFHE) scheme [^10] With an assumption on the form of encrypted ciphertexts, which is satisfied by both Mahadev’s and Brakerski’s QFHE schemes, [Bra18, Mah20]; see LABEL:Subsec:techoverviewQFHE. and (2) a new primitive, the oblivious pad, which we introduce below. QFHE can be realised assuming the quantum hardness of the Learning With Errors problem (LWE) [Reg09]. We show how to construct the oblivious pad under the same assumption in the quantum random oracle model—which in turn can be heuristically instantiated using a cryptographic hash function, such as SHA3.111111In the random oracle model [BR93] (ROM), a hash function $f$ is modelled as a uniformly random black-box function: parties can evaluate it by sending a query $x$ and receiving $f(x)$ in return. In the quantum random oracle model (QROM), such queries can also be made in superposition. A proof of security in this model is taken to be evidence for security of the protocol when the black-box is replaced by, for example, a suitable hash function $f$. This is because, informally, any attack on the resulting protocol must necessarily exploit the structure of $f$.
Figure 1: The compiler in this work compiles a much larger set of games compared to the one in [KLVY23].
Our result is motivated by the recent work of Kalai et al. [KLVY23] (KLVY from here on) that converts any non-local game into a test of quantumness with a single device. Consider a non-local game with two players, Alice and Bob. The central idea behind the KLVY compiler is to use cryptography to enforce spatial separation within a single quantum device, i.e. to enforce that Alice and Bob’s measurements occur on separate subsystems. The KLVY compiler relies on the following mechanism to cryptographically enforce spatial separation: Alice’s question and answer are encrypted (using a quantum fully homomorphic encryption (QFHE) scheme), while Bob’s question and answer are in the clear. The referee, who holds the decryption key, can then test the correlation across the two. Unfortunately, this approach does not extend to contextuality, at least not in any direct way, since in a contextuality game there is no notion of Alice and Bob.
In contrast, our work can be seen as using cryptography to enforce “temporal separation”, i.e. to restrict communication between sequential measurements. In a nutshell, our idea is to ask the first question under a homomorphic encryption and the second question in the clear, as in KLVY, but with the following important difference. In KLVY, the quantum device prepares an entangled state, whose first half is encrypted, and used to homomorphically answer the first question, while the second half is not, and is used to answer the second question in the clear. In our protocol, there are no separate subsystems between the two rounds: instead, the encrypted post-measurement state from the first round (which results from the measurement performed to obtain the encrypted answer) is re-used in the second round. The technical barrier is, of course, the following: how can the post-measurement state be re-used if it is still encrypted? The two most natural approaches do not work:
  • (i)
    Providing the decryption key to the quantum device is clearly insecure as it allows the device to learn the first question in the clear.
  • (ii)
    Homomorphically encrypting the second question does not work either because the quantum device can correlate its answers to the two questions “under the hood of the homomorphic encryption”.
We circumvent this barrier by introducing a procedure that allows the prover to obliviously “re-encrypt” the post-measurement state non-interactively. This re-encryption procedure is such that it allows the verifier to achieve the following: the verifier can now reveal some information that allows a quantum prover to access the post-measurement state in the clear, while a PPT prover does not learn the first question. More precisely, the verifier does not directly reveal the original decryption key, which would clearly be insecure, as pointed out earlier. Instead, the verifier expects the prover to “re-encrypt” its state using the new procedure, and then the verifier is able to safely reveal the resulting “overall” decryption key. Crucially, while a classical prover is unable to make use of the additional information to beat the classical value ${\rm valNC}$, we show that there is an efficient quantum prover that can access the post-measurement state in the clear, and proceed to achieve the quantum value ${\rm valQu}$.
The main technical tool that we introduce to formalise this idea, which may find applications elsewhere, is a primitive that we call oblivious (Pauli) pad. The oblivious pad takes as input a state |$\ket{\psi}$ and a public key $\mathsf{pk}$ and produces a Pauli-padded state $X^{x}Z^{z}\ket{\psi}$ together with a string $s$ (which can be thought of as encrypting $x$ and $z$ using $\mathsf{pk}$). The string $s$ can be used to recover $x,z$ given the corresponding secret key $\mathsf{sk}$. The security requirement is modelled as the following distinguishing game between a PPT prover and a challenger:
  • The challenger generates public and secret keys $(\mathsf{pk},\mathsf{sk})$ and sends $\mathsf{pk}$ to the prover.
  • The prover produces a string $s$ and sends it to the challenger.
  • The challenger either returns $(x,z)$ (as decoded using $s$ and $\mathsf{sk}$), or a fresh pad $(\tilde{x},\tilde{z})$ sampled uniformly at random.
We require that no PPT prover can distinguish between the two cases with non-negligible advantage. We show that an oblivious pad can be realised based on ideas from [BCM+21] in the quantum random oracle model. [^12] We do not foresee any obstacle in obtaining a construction in the plain model, at the cost of constantly many additional rounds of interaction. Crucially, while in KLVY the second phase of the protocol happens on a different subsystem, the oblivious pad is what allows us to carry out the second phase on the same subsystem, as depicted in Figure 2.
Figure 2: A schematic comparing the non-local game compiler [KLVY23] with our contextuality game compiler. The key idea in KLVY is to ask the first question of a non-local game under a homomorphic encryption and the second one in the clear, with the prover using two entangled subsystems (one that is encrypted, and one in the clear). In our compiler, the oblivious pad ($\mathsf{OPad}$) allows the prover to “re-encrypt” its post-measurement state, just before Round 2. Upon obtaining information about the “re-encryption” that took place, the verifier can then safely reveal the “overall” decryption key in Round 2, allowing the prover to proceed with the next measurement in the clear.
Our second contribution streamlines the ideas introduced to prove LABEL:Thm:infMain1 in order to obtain a 2-round “proof of quantumness” relying on the classical hardness of the Learning-with-Errors (LWE) problem [Reg09]. Our construction has the main advantage of being simpler than existing ones in the literature, in the sense explained below. Our proof of quantumness makes use of Noisy Trapdoor Claw-Free functions (NTCFs), introduced in [BCM+21] (but it does not require the NTCFs to have an “adaptive hardcore bit” property). It relies on the particular structure of the “encrypted CNOT operation” introduced in Mahadev’s QFHE scheme [Mah20]. We informally state our result, but we defer the construction to the technical overview.
Theorem 2 (Informal).
Assuming the classical hardness of LWE, there exists a 2-round (i.e. 4-message) proof of quantumness with the following properties:
  • It requires only one coherent evaluation of an NTCF, and one layer of single-qubit Hadamard gates.
  • The quantum device only needs to maintain a single qubit coherent in-between the two rounds.
Our proof of quantumness can be seen as combining ideas from [KLVY23] and [KMCVY22, AMMW22]. It is simpler than existing proofs of quantumness in the following ways:
  • Single encrypted CNOT operation. The 2-round proof of quantumness of KLVY is based on a QFHE scheme. Concretely, an implementation of their proof of quantumness based on Mahadev’s QFHE scheme requires performing a homomorphic controlled-Hadamard gate, which requires three sequential applications of the “encrypted CNOT operation”. Crucially, these three operations require computing the NTCF three times in superposition while maintaining coherence all along. Our 2-round proof of quantumness only requires a single application of the “encrypted CNOT” operation. Moreover, in KLVY, the encrypted subsystem, on which Alice’s operations are applied homomorphically needs to remain entangled with a second subsystem, which is used to perform Bob’s operations in the clear. In contrast, in our proof of quantumness, the encrypted CNOT operation and the subsequent operations in the clear happen on a single system (of the same size as Alice’s in KLVY).
  • Single qubit coherent across rounds and 2 rounds of interaction. Compared to the 3-round proof of quantumness of [KMCVY22], ours requires one less round of interaction. However, more importantly than the number of rounds, the protocol from [KMCVY22] requires the quantum device to keep a superposition over preimages of the NTCF coherent in-between rounds, while waiting for the next message. In contrast, both our proof of quantumness and that of KLVY only require the quantum device to keep a single qubit coherent in-between rounds.
  • Simple quantum operations. The 2-round proof of quantumness of [AMMW22] matches ours in that it requires a single coherent application of an NTCF based on LWE, and the quantum device only needs to keep a single qubit coherent in between rounds. However, their protocol requires the prover to perform single-qubit measurements in “rotated” bases coming from a set of a size that scales linearly with the LWE modulus (for which typical parameters are $\approx 10^{2}$ to $10^{3}$). Their completeness-soundness gap also suffers a loss compared to ours that comes from the use of the “rotated” basis measurements.
To the best of our knowledge, the only aspect in which our proof of quantumness compares unfavourably with existing ones, e.g. [KMCVY22], is that, based on current knowledge of constructions of NTCFs, our proof of quantumness requires a construction from LWE (in order to implement the “encrypted CNOT” procedure), whereas [KMCVY22] has the flexibility that it can be instantiated using any TCF, e.g. based on Diffie-Hellman or Rabin’s function. While the latter are more efficient to implement, they also generally require a larger security parameter (inverting Rabin’s function is as hard as factoring, whereas breaking LWE is as hard as worst-case lattice problems). Hence, it is currently still unclear which route is closer to a realisation at scale.
Some existing proofs of quantumness are non-interactive (i.e. 1-round), with security in the random oracle model [BKVV20, ACGH20, YZ22]. Likewise, our proof of quantumness can also be made non-interactive by using the Fiat-Shamir transformation [FS87] (where the computation of the hash function is classical, and does not increase the complexity of the actual quantum computation). The proof of quantumness in [YZ22] has the additional desirable property of being publicly-verifiable, although it currently seems to be more demanding than the others in terms of quantum resources.
1.2 Future Directions
  • Better contextuality compilers. Focusing on compilers for contextuality, the following important aspect remains to be strengthened. The current compiler does not in general achieve quantum soundness (in the sense that there are contextuality games in which a QPT prover can do much better than the “compiled” quantum value from LABEL:Thm:infMain1). Interestingly, recent works show that KLVY does satisfy quantum soundness for certain families of non-local games [NZ23, BGKM+23b, CMM+24].
  • Oblivious Pauli pad. We think that the new functionality of the oblivious Pauli pad has the potential to be useful elsewhere, and we leave this exploration to future work. A related question is to realise the oblivious Pauli pad in the plain model (without increasing the number of rounds).
  • More efficient “encrypted CNOT”. Turning to proofs of quantumness, the broad goal is of course to simplify these even further towards experimental implementations. One concrete direction in which our proof of quantumness could be simplified is the following. Currently, we use an NTCF that supports the “encrypted CNOT” operation from [Mah20]. However, our only requirement is that the NTCF satisfies the potentially weaker property that it hides a bit in the xor of the first bit of a pair of preimages (see LABEL:Def:ntcf-2). This is because we only need to perform the “encrypted CNOT” operation once, and not repeatedly as part of a full-fledged homomorphic computation. It would be interesting to see if the weaker requirement could be achieved by simpler claw-free functions (from LWE or other assumptions, like DDH or factoring).
  • Testing other sources of quantumness. Many other sources of quantumness have been identified in the literature, such as generalised contextuality (which allows arbitrary experimental procedures, not just projective measurements, and a broad class of ontological models, not just deterministic ones) [Spe05] and the Leggett-Garg experiment (which is the time analogue of Bell’s experiment) [LG85, BMKG13]. These all suffer from the same limitation as contextuality, and our results raise the following question: can one construct analogous operational single-device tests for these sources of quantumness as well?
  • Testing indefinite causal order. More ambitiously, one could try to separate quantum mechanics from more general theories such as those with indefinite causal order [CDPV13, OCB12] (i.e. theories that obey causality only locally and that may not admit a definite causal order globally). For instance, a recent result gives evidence (in the black-box model) that indefinite causal order does not yield any relevant advantage over quantum mechanics [AMP23]. Perhaps one can obtain a clear separation under cryptographic assumptions?
Acknowledgements
We thank Ulysse Chabaud, Mauro Morales and Thomas Vidick for their feedback on early drafts of this work. We thank Yusuf Alnawakhtha, Alexandru Gheorghiu, Manasi Mangesh Shingane, and Urmila Mahadev for various helpful discussions. ASA acknowledges support from the U.S. Department of Defense through a QuICS Hartree Fellowship, IQIM, an NSF Physics Frontier Center (GBMF-1250002) and MURI grant FA9550-18-1-0161. Part of the work was carried out while ASA was visiting the Simons Institute for the Theory of Computing. AC acknowledges support from the UK National Quantum Computing Centre through the Quantum Advantage Pathfinder (QAP) research programme (EP/X026167/1).
2 Technical Overview
In this overview, we start by briefly recalling non-local games, and introducing the notion of contextuality with some examples. We then build up towards our compiler for contextuality games, by first introducing the ideas behind the KLVY compiler, and then describing why new ideas are needed to achieve a compiler in the contextuality setting. We then describe our main novel technical tool, the oblivious pad, and how we use it to realise a compiler for contextuality games. We also discuss the main ideas in the proof. Finally, we describe how some of the new ideas can be streamlined to obtain a potentially simpler proof of quantumness. The latter can be understood without any reference to contextuality, and the interested reader may wish to skip directly to it (Section 2.6).
2.1 Non-local Games
Let $A,B,X,Y$ be finite sets. A $2$-player non-local game is specified by a predicate $\mathsf{pred}:A\times B\times X\times Y\to\{0,1\}$ which indicates whether the players win or not, and a probability distribution ${\cal D}_{{\rm questions}}$ over the questions, which specifies $\Pr(x,y)$ for $(x,y)\in X\times Y$. The game consists of a referee and two players, Alice and Bob, who can agree on a strategy before the game starts, but cannote communicate once the game starts. The game proceeds as follows: the referee samples questions $(x,y)\leftarrow{\cal D}_{{\rm questions}}$, sends $x$ to Alice and $y$ to Bob, and receives their answers $a\in A$ and $b\in B$ respectively. Their success probability can be written as
$\Pr({\rm win})=\sum_{x,y,a,b}\mathsf{pred}(a,b,x,y)\Pr(a,b|x,y)\Pr(x,y)\,,$ (1)
where the strategy used by Alice and Bob specifies $\Pr(a,b|x,y)$.
Local Hidden Variable Strategy
The “local hidden variable” model allows Alice and Bob to share a classical (random) variable $r$, and then have their answers be arbitrary, but fixed, functions of their respective questions, i.e.
$\Pr(a,b|x,y)=\sum_{r}P_{A}(a|x,r)P_{B}(b|y,r)P_{R}(r)$
where $P_{A}$, $P_{B}$ and $P_{R}$ are probability distributions that specify Alice and Bob’s strategy. We refer to the optimal winning probability achievable by local hidden variable strategies as the classical value of the game.
Quantum Strategy
A quantum strategy allows Alice and Bob to share a state $\left|\psi\right\rangle_{AB}$, and use local measurements $M^{A}_{x}=\{M_{a|x}^{A}\}_{a}$ and $M_{y}^{B}=\{M_{b|y}^{B}\}_{b}$ to produce their answers (where the measurements can be taken to be projective without loss of generality), so that
$\Pr(a,b|x,y)=\left\langle\psi\right|M_{a|x}^{A}\otimes M_{b|y}^{B}\left|\psi\right\rangle.$
We refer to the optimal winning probability achievable by quantum strategies as the quantum value of the game.
It is well-known that there exist non-local games (like the CHSH game) where the quantum value exceeds the classical value. However, is it necessary to consider spatial separation (i.e. a tensor product structure) to observe such a “quantum advantage”? A partial answer is no: one can consider contextuality.
Figure 3: A two-player non-local game. Alice and Bob get $(x,y)$ from the referee and respond with $(a,b)$. They cannot communicate once the game starts.
2.2 Contextuality
[…]
  • The referee samples $C=\{q_{1},\ldots,q_{k}\}\in C^{{\rm all}}$ according to ${\cal D}_{{\rm contexts}}$, and sends $C$ to the player.
  • The player responds with answers $\{a_{1},\ldots,a_{k}\}\in A^{k}$.
The success probability is
$\Pr({\rm win})=\sum_{a_{1},\ldots,a_{k},C}\mathsf{pred}(a_{1},\ldots,a_{k},C)\Pr(a_{1},\ldots,a_{k}|C)\Pr(C)\,,$ (2)
where the strategy used by the player specifies $\Pr(a_{1},\ldots,a_{k}|C)$.
Non-contextual strategy
This is the analogue of a “local hidden variable” strategy. A deterministic assignment maps each question $q\in Q$ to a fixed answer $a_{q}$. This represents the following strategy: upon receiving the context $C=\{q_{1},\ldots,q_{k}\}\in C^{{\rm all}}$, return $(a_{q_{1}},\ldots,a_{q_{k}})$ as the answer. A strategy that can be expressed as a convex combination of deterministic assignments is referred to as non-contextual, i.e. the answer to a question $q$ is independent of the context in which $q$ is being asked.
Quantum strategy
A quantum strategy is specified by a quantum state |$\ket{\psi}$, and a collection of observables $\mathbf{O}=\{O_{q}\}_{q\in Q}$, such that, for any context $C=\{q_{1},\ldots,q_{k}\}\in C^{{\rm all}}$, the observables $O_{q_{1}},\ldots,O_{q_{k}}$ are compatible (i.e commuting). The strategy is the following: upon receiving context $C=\{q_{1},\ldots,q_{k}\}\in C^{{\rm all}}$, measure observables $O_{q_{1}},\ldots,O_{q_{k}}$ on state |$\ket{\psi}$, and return the respective outcomes $a_{1},\ldots,a_{k}$.
Quantum mechanics is contextual in the sense there are examples of games for which a quantum strategy can achieve a higher winning probability than the best non-contextual strategy. However, crucially, unlike a non-local game, a contextuality game does not directly yield an “operational test” of contextuality. The issue is that there is no clear way for the referee to enforce that the player’s answer to question $q$ is consistent across the different contexts in which $q$ appears!
We informally describe three simple examples: the magic square game (Peres-Mermin) [Per90, Mer90, Cab01, Ara04], non-local games, and the KCBS experiment (the contextuality analogue of the Bell/CHSH experiment) [KCBS08]. We do this by directly specifying a quantum strategy first and then “deriving” the corresponding game (where the observables are just labels for the questions). In doing so, we abuse the notation slightly and identify questions with observables.
Example 1 (Peres-Mermin (Magic Square) [Per90, Mer90]).
Consider the following set of observables
$\ \mathbf{O}:=\\ \\ \\ \\ \end{array}\begin{array}[]{ccccc}\{X\otimes\mathbb{I},&\mathbb{I}\otimes Z,&X\otimes Z,&&\mathbb{I}\\ \mathbb{I}\otimes X,&Z\otimes\mathbb{I},&Z\otimes X,&&\mathbb{I}\\ X\otimes X,&Z\otimes Z,&\ Y\otimes Y\}&&\mathbb{I}\\ \\ \mathbb{I}&\mathbb{I}&-\mathbb{I}\end{array}$
(where $X,Y,Z$ are Pauli matrices). They satisfy the following properties: (a) they take ±$\pm 1$ values (i.e. they have ±$\pm 1$ eigenvalues), (b) operators along any row or column commute, and (c) the product of observables along any row or column equals $\mathbb{I}$, except along ${\rm col}_{3}$, where it equals $-\mathbb{I}$. If we define the set of contexts by $C^{{\rm all}}:=\{{\rm row}_{1},{\rm row}_{2},{\rm row}_{3},{\rm col}_{1},{\rm col}_{2},{\rm col}_{3}\}$, it is not difficult to see that no deterministic assignment can be such that the condition on the products is satisfied along each row and column. For instance, the assignment
$\begin{array}[]{ccc}1&-1&-1\\ 1&-1&-1\\ 1&1&?\end{array}$
satisfies all constraints except one: if the question mark is $1$, then the condition along ${\rm col}_{3}$ fails, and if it is $-1$ then the condition along ${\rm row}_{3}$ fails. To satisfy both, somehow the value assigned to the last “observable” has to depend on the context, ${\rm row}_{3}$ or ${\rm col}_{3}$, in which it appears—the assignment has to be “contextual”. In quantum mechanics, all of the constraints can be satisfied using the observables described. One can in fact measure these observables on an arbitrary state $\left|\psi\right\rangle$ to win with probability $1$. One thus concludes that quantum mechanics is “contextual” in this sense.
2-player non-local games can be viewed as a special case of contextuality games as follows (and similarly for non-local games with more players).
Example 2 (2-player non-local games).
Given a quantum strategy for a non-local game (as in LABEL:Subsec:Non-local-Game), we identify the measurements $M^{A}_{x}$ and $M^{B}_{y}$ with corresponding observables. Then, define $\mathbf{O}:=\{M_{x}^{A}\otimes I\}_{x}\cup\{I\otimes M_{y}^{B}\}_{y}$ and $C^{{\rm all}}:=\{\{M_{x}^{A}\otimes I,I\otimes M_{y}^{B}\}\}_{x,y}$. It is not hard to see that the set of non-contextual strategies is the same as the set of local hidden variable strategies.
Some contextuality games can yield a separation between non-contextual and quantum strategies with even smaller quantum systems than what is possible for non-local games. The following example yields a separation using just a qutrit—a single $3$-dimensional system (in contrast, non-local games require at least two qubits, i.e. dimension $4$, as in the CHSH game). For contextuality $3$ dimensions are necessary and sufficient [KS67]. [^13] There are generalisations of contextuality [Spe05] that can give a separation with dimension $2$, but we do not consider these here.
Example 3 (KCBS [KCBS08]).
Consider a $3$-dimensional vector space spanned by {|$\{\ket{0},\ket{1},\ket{2}\}$. Let |$\ket{\psi}=\ket{0}$ and define five vectors
[…]
indexed by $q\in\{1,\dots,5\}$ where $\phi_{q}=4\pi q/5$ and $\cos^{2}\theta=\cos(\pi/5)/(1+\cos(\pi/5))$. The heads of these vectors form a pentagon with |$\ket{\psi}$ at the centre, and vectors indexed consecutively are orthogonal, i.e. $\left\langle v_{q}|v_{q+1}\right\rangle=0$ (where we take the indices to be periodic) as illustrated in LABEL:Fig:KCBS. Define $\mathbf{O}:=\{\Pi_{q}\}_{q\in 1\dots 5}$ where $\Pi_{q}:=\left|v_{q}\right\rangle\left\langle v_{q}\right|$ and $C^{{\rm all}}:=\left\{\{\Pi_{1},\Pi_{2}\},\{\Pi_{2},\Pi_{3}\}\dots\{\Pi_{5},\Pi_{1}\}\right\}$. The referee asks a context $C\leftarrow C^{{\rm all}}$ uniformly at random from the set of all contexts and the player wins if the answer is either $(0,1)$ or $(1,0)$ (i.e. neighbouring assignments should be distinct). It is not hard to check that a non-contextual strategy wins at most with probability $4/5=0.8$, while the quantum strategy described above wins with probability $\frac{2}{\sqrt{5}}\approx 0.8944$.
Figure 4: An illustration of the optimal quantum strategy corresponding to the KCBS Game defined in LABEL:Exa:KCBSintro. Here the red vector denotes the quantum state |$\ket{\psi}$ and the blue ones |$\ket{v_{q}}$ correspond to projective measurements, $\Pi_{q}=\ket{v_{q}}\bra{v_{q}}$ |. Consecutively indexed blue vectors, i.e. |$\ket{v_{q}},\ket{v_{q+1}}$, are orthogonal (indexing is periodic).
One route towards obtaining an operational test of contextuality is to find a way to enforce that measurements on a single system happen “sequentially”, i.e. they are separated in “time” (as opposed to being separated in “space”, which is the case for non-local games). We describe one folklore attempt at constructing an operational test, which assumes that the device is “memoryless”. This example is not essential to understanding our results, and may be skipped at first read.
(a) Memoryless Device
(b) A memoryless operational test, illustrated using the Peres-Mermin magic square.
Figure 5: The folklore memoryless interpretation of contextuality.
The “memoryless” attempt at an operational test
As denoted in LABEL:Fig:memoryless, consider a device that takes as input a question $q$, produces an answer $a$, and then forgets the question. The referee in this case proceeds as follows:
  1. 1.
    Samples a context $C$ from $C^{{\rm all}}$ with probability $\Pr(C)$.
  2. 2.
    Sequentially asks all the questions $q_{1}\dots q_{k}$ in the context $C$.
  3. 3.
    Accepts the answers $a_{1}\dots a_{k}$ if the constraint corresponding to $C$ holds, i.e. if $\mathsf{pred}(a_{1}\dots a_{k},C)$ is true.
Note that the most general deterministic model for the device is one that encodes a “truth table” $\tau:Q\to A$ that maps questions to answers. Since there is no memory, no previous question can affect the way the device answers the current question. The most general quantum device, on the other hand, starts with an initial $\left|\psi\right\rangle$ and to each question, assigns an observable $O_{q}$, which is measured to obtain an answer $a$ to the question $q$.
Let us work out an example to illustrate the key point. Consider again the Magic Square game from LABEL:Exa:MagicSquare. Suppose the first question the referee asks is the value of the bottom right box of the magic square (see LABEL:Fig:MagicSquareMemoryless). It can then either ask questions completing the corresponding row or column. For a deterministic memoryless device, since a single truth table $\tau$ is being used, it is impossible to satisfy all the constraints of the magic square. However, a deterministic device with memory can satisfy all the constraints. This is because before answering the last question, it can learn whether the third column is being asked or the third row and thus, it can answer the last question to satisfy the constraint. Thus “classical memory”, allows for classical simulation of contextuality in this test, without using any quantum effects. In fact, [KGP+11, CGGX18] even quantifies the amount of classical memory needed to simulate contextuality.
Evidently, the glaring limitation of this attempt is that there is no operational way of ensuring that a device is “memoryless”. Thus, this has remained a barrier despite the numerous attempts [LLS+11, UZZ+13, JRO+16, ZKK+17, MZL+18, LMZ+18, ZXX+19, UZZ+20, WZL+22, HXA+23, LMX+23, BRV+19b, BRV+19a, XSBC24, SSA20, BCG+22] and the fact that contextuality has, in general, been a bustling area of investigation [BCG+22, WZL+22, HXA+23, LMX+23, Liu23, XSBC24].
Our construction follows the same general approach of enforcing separation in “time” instead of in “space”, but is a radical departure from the attempt described above. We only assume that the device is computationally bounded, and we use cryptographic techniques to construct an operational test, which we describe in the next sections.
2.3 Quantum Fully Homomorphic Encryption (QFHE)
We informally introduce fully homomorphic encryption. We start with the classical notion.
Fully Homomorphic Encryption (FHE)
A homomorphic encryption scheme is specified by four algorithms, $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec},\mathsf{Eval})$, as follows:
  • $\mathsf{Gen}$ takes as input a security parameter $1^{\lambda}$, and outputs a secret key $\mathsf{sk}$.
  • $\mathsf{Enc}$ takes as input a secret key $\mathsf{sk}$ and a message $s$, and outputs a ciphertext $c$. We use the notation $\mathsf{Enc}_{\mathsf{sk}}=\mathsf{Enc}(\mathsf{sk},\cdot)$.
  • $\mathsf{Dec}$ takes as input a secret key $sk$, a ciphertext $c$ and outputs the corresponding plaintext $s$. We use the notation $\mathsf{Dec}_{\mathsf{sk}}=\mathsf{Dec}(\mathsf{sk},\cdot)$.
The “homomorphic” property says that a circuit $\mathsf{Circuit}$ can be applied on an encrypted input to obtain an encrypted output, i.e.
  • $\mathsf{Eval}$ takes as input a circuit $\mathsf{Circuit}$, a ciphertext $c$, and an auxiliary input $\mathsf{aux}$, and outputs a ciphertext $c^{\prime}$. The following is satisfied. Let $c\leftarrow\mathsf{Enc}_{\mathsf{sk}}(s)$ and $c^{\prime}\leftarrow\mathsf{Eval}(\mathsf{Circuit},c,\mathsf{aux})$, then $\mathsf{Dec}_{\mathsf{sk}}(c^{\prime})=\mathsf{Circuit}(s,\mathsf{aux})$.
Crucially, note that the secret key $\mathsf{sk}$ is not needed to apply the $\mathsf{Eval}$ algorithm. The security condition is the usual one: that an encryption of $s$ should be indistinguishable from an encryption of $s^{\prime}\neq s$. If $\mathsf{Eval}$ supports evaluation for the class of all polynomial-size circuits (in the security parameter), then the scheme is said to be fully homomorphic. [^14] While homomorphic schemes for restricted families of circuits have been known for some time, a “fully homomorphic scheme” was only discovered somewhat recently in [Gen09].
Quantum Fully Homomorphic Encryption (QFHE)
For this overview, it suffices to take a QFHE scheme to be the same as an FHE scheme, except that it allows messages and auxiliary inputs to be quantum states, and circuits to be quantum circuits. The QFHE schemes that are relevant to our work [Mah20, Bra18] satisfy the following additional properties.
  1. 1.
    Classical encryption/decryption.
    $\mathsf{Gen}$ is a classical algorithm, while $\mathsf{Enc},\mathsf{Dec}$ become classical algorithms when their inputs are classical. In particular, this means that classical inputs are encrypted into classical ciphertexts. This property is essential when the scheme is deployed in protocols involving classical parties, as will be the case here. [^15] The first schemes to satisfy this property appeared recently in the breakthrough works [Mah20, Bra18].
  2. 2.
    Locality is preserved.
    Consider an arbitrary bipartite state |$\ket{\psi}_{AB}$ and let $M^{A}$ and $M^{B}$ be circuits acting on registers $A$ and $B$ respectively with a measurement at the end. The property requires that correlations between the measurement outcomes from $M^{A}$ and $M^{B}$ should be the same in the following two cases:
    (i) Register $A$ is encrypted, $M_{A}$ is applied using $\mathsf{Eval}$, and the result is decrypted. Let $a$ be the decrypted outcome. $M^{B}$ is applied to register $B$. Let $b$ be the outcome.
    (ii) Apply $M^{A}$ on register $A$ and $M^{B}$ on register $B$. Let $a$ and $b$ respectively be the outcomes.
  3. 3.
    Form of Encryption.
    Encryption of a state $\left|\psi\right\rangle$ takes the form $(X^{x}Z^{z}\left|\psi\right\rangle,\widehat{xz})\leftarrow\mathsf{Enc}_{\mathsf{sk}}(\left|\psi\right\rangle)$, where $X^{x}$ applies a Pauli $X$ to the $i$-th qubit based on the value of $x_{i}$, and $Z^{z}$ is defined similarly. $\widehat{xz}$ is an encryption of the pad $xz$.
All known constructions of QFHE schemes satisfying property 1 also satisfy properties 2 and 3. The KLVY compiler, as one might guess, relies on property 2. Our compiler relies on property 3.
2.4 The KLVY Compiler
Consider a two-player non-local game specified by question and answer sets $X,Y,A,B$, a predicate $\mathsf{pred}:A\times B\times X\times Y\to\{0,1\}$ and a distribution over the questions ${\cal D}_{{\rm questions}}$ that specifies $\Pr(x,y)$. The KLVY compiler takes this non-local game as input and produces the following single-prover game, where the verifier proceeds as follows:
  • Round 1. Sample $(x,y)\leftarrow{\cal D}_{{\rm questions}}$, and a secret key $\mathsf{sk}$ for a $\mathsf{QFHE}$ scheme. Send an encryption $c_{x}$ of “Alice’s question”, and get an encrypted answer $c_{a}$.
  • Round 2. Send an encryption $c_{y}$ of “Bob’s question”, and get an encrypted answer $c_{b}$ in the clear. Decrypt the answers using $\mathsf{sk}$, and accept if $\mathsf{pred}(a,b,x,y)=1$.
The honest prover prepares the entangled state |$\ket{\psi}_{AB}$ corresponding to the optimal quantum strategy for the non-local game. It uses $\mathsf{QFHE}$’s $\mathsf{Eval}$ algorithm on subsystem $A$ to answer question $x$ according to Alice’s optimal strategy, and answers question $y$ in the clear using Bob’s optimal strategy on subsystem $B$. More formally, they proceed as follows.
(Honest Quantum) Prover Verifier
Prepare $\left|\psi\right\rangle_{AB}$ corresponding to the non-local game’s optimal quantum strategy.
$\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$
$(x,y)\leftarrow{\cal D}_{{\rm questions}}$ $c_{x}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(x)$
$\overset{c_{x}}{\longleftarrow}$
Homomorphically compute the answer to $x$ using “Alice’s strategy” on system $A$. Let $c_{a}$ be the encrypted answer.
$\overset{c_{a}}{\longrightarrow}$
[…]
Compute the answer to $y$ using “Bob’s strategy” (in the clear) on system $B$. Let the answer be $b$.
[…]
$a=\mathsf{Dec}_{\mathsf{sk}}(c_{a})$ Accept if $\mathsf{pred}(a,b,x,y)=1$.
Theorem ([KLVY23], informal).
Consider a two-player non-local game with classical and quantum values $\omega_{c}$ and $\omega_{q}$ respectively. The corresponding KLVY-compiled single-prover game satisfies the following:
  • (Completeness) The QPT prover described above makes the verifier accept with probability $\omega_{q}-\mathsf{negl}(\lambda)$.
  • (Soundness) Every PPT prover makes the verifier above accept with probability at most $\omega_{c}+\mathsf{negl}(\lambda)$.
Here $\mathsf{negl}$ are (possibly different) negligible functions.
For completeness, we briefly outline the general compiler for non-local games with $k>2$ players. To compile games with $k$ players, KLVY generalises the procedure above as follows: the compiled game consists of $k$ rounds–each round consisting of a question and an answer (so $2k$ messages in total). The first $k-1$ questions are encrypted (using $k-1$ independent random $\mathsf{QFHE}$ keys) and the last question is asked in the clear.
Let us take a moment and build some intuition about the KLVY compiler. At first, one might think that it would be even more secure to also encrypt $y$. However, as we remarked in the introduction, this is not the case: since the prover can compute any desired answer $b$ using $x,a$ and $y$ under the hood of the QFHE, it can ensure that $b$ is such that $\mathsf{pred}(a,b,x,y)=1$ (for any non-trivial choice of $\mathsf{pred}$). Instead, the key observation of KLVY is that the verifier is testing the correlation between encrypted answers and answers in the clear, and it turns out that a quantum prover can produce stronger correlations than any PPT prover can. How is this intuition formalised? The key idea is that, since a PPT prover is classical, one can rewind the PPT prover to obtain answers to all possible second round questions. This is equivalent to obtaining Bob’s entire assignment of answers to questions (corresponding to a fixed encrypted question and answer for Alice from the first round). Suppose for a contradiction that such a PPT prover wins with probability non-negligibly greater than the classical value. Then, it must be that Bob’s entire assignment is non-trivially correlated to “Alice’s encrypted question”. This information can thus be used to obtain a non-negligible advantage in guessing the encrypted question, breaking security of the QFHE scheme.
Now that we understand the KLVY construction and the key insight behind their proof, we will discuss barriers to extending these ideas to contextuality, and our approach to circumvent these.
2.5 Contribution 1 — A Computational Test of Contextuality
For simplicity, let us first restrict to contextuality games with contexts of size $2$.
Attempts at extending KLVY to contextuality.
Let us consider compilers for contextuality games where the verifier proceeds in an analogous way as the verifier in the KLVY compiler:
  • Round 1. Sample a context $C=(q_{1},q_{2})\leftarrow{\cal D}_{{\rm contexts}}$. Send a $\mathsf{QFHE}$ encryption $c_{q_{1}}$ of $q_{1}$ and receive as a response an encryption $c_{a_{1}}$ of $a_{1}$.
  • Round 2. There is no clear way to proceed. Three natural approaches are listed below.
The honest prover’s state after Round 1 is encrypted and has the form $(X^{x}Z^{z}\ket{\psi_{(q_{1},a_{1})}},\widehat{xz})$ where $\widehat{xz}$ denotes a classical encryption of the strings $xz$ (using Property 3 of the $\mathsf{QFHE}$ scheme), as detailed below.
(Honest Quantum) Prover Verifier
Prepare $\left|\psi\right\rangle$ corresponding to the quantum strategy. $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ $C\leftarrow{\cal D}_{{\rm contexts}}$, where $C=(q_{1},q_{2})$ $c_{q_{1}}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(q_{1})$
$\overset{c_{q_{1}}}{\longleftarrow}$
Evaluate the answer to $q_{1}$ under QFHE to get an encrypted answer $c_{a_{1}}$. The post-measurement state is now $\left(X^{x}Z^{z}\left|\psi_{q_{1}a_{1}}\right\rangle,\widehat{xz}\right).$
$\overset{c_{a_{1}}}{\longrightarrow}$
${\longleftarrow}$longrightarrow
Here are three natural approaches for how to proceed, and how they fail.
  1. 1.
    Proceed just as KLVY: Ask question $q_{2}$ in the clear. This does not work because, unlike in the original KLVY setup, there is no analogue of system $B$ which is left in the clear. Here the prover holds an encrypted state so it is unclear how $q_{2}$ can be answered with any non-trivial dependence on the state under the encryption.
  2. 2.
    Ask the second question also under encryption. If the same key is used for encryption, then, as we argued for KLVY, the predicate of the game can be satisfied by computing everything homomorphically. If the keys are independent, then we essentially return to the problem in item 1.
  3. 3.
    Reveal the value of the (classically) encrypted pad $\widehat{xz}$ and ask question $q_{2}$ in the clear. This has a serious issue: the prover can simply ask for the encrypted pad corresponding to $c_{q_{1}}$ and thereby learn $q_{1}$ (or at least some bits of $q_{1}$). Once $q_{1}$ is learned, again, the predicate of the game can be trivially satisfied.
The Oblivious Pauli Pad
As mentioned in Section 1, our compiler relies on a new cryptographic primitive, that we introduce to circumvent the barriers described above. Here, it helps to be a bit more formal. Let $\mathbf{U}:=\{U_{k}\}_{k\in K}$ be a group of unitaries acting on the Hilbert space ${\cal H}$. We take this group to be the set of Paulis $\{X^{x}Z^{z}\}_{xz}$, as this makes the primitive compatible with the the form of the QFHE scheme that we will employ later in our compiler. Nonetheless, we use the general notation $\{U_{k}\}_{k\in K}$, as it simplifies the presentation. We define the oblivious $\mathbf{U}$ pad as follows.
The oblivious $\mathbf{U}$ pad is a tuple of algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ where $\mathsf{Gen}$ and $\mathsf{Dec}$ are PPT. Let $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$ be the public and the secret keys generated by $\mathsf{Gen}$. Encryption takes the form $(U_{k}\left|\psi\right\rangle,s)\leftarrow\mathsf{Enc}_{\mathsf{pk}}(\left|\psi\right\rangle)$, where $k=\mathsf{Dec}_{\mathsf{sk}}(s)$. Notice the similarity with the post-measurement state in the discussion above (we will return to this in a moment). The security requirement is that no PPT algorithm can win the following security game with probability non-negligibly greater than $1/2$.
Prover Verifier
$(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$
[…]
$b\leftarrow\{0,1\}$ $k_{0}\leftarrow K$ $k_{1}=\mathsf{Dec}_{\mathsf{sk}}(s)$
$\overset{k_{b}}{\longleftarrow}$
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b=b^{\prime}$
The security game formalises the intuition that no PPT prover can distinguish between the correct “key” $k_{1}$ and a uniformly random “key” $k_{0}$. We emphasize, in words, the two distinctive features of this primitive:
  • By running $\mathsf{Enc}$, a QPT prover can obtain, given a state |$\ket{\psi}$, an encryption of the form $(U_{k}\left|\psi\right\rangle,s)$, where $k=\mathsf{Dec}_{\mathsf{sk}}(s)$.
  • There is no way for a PPT prover, given $\mathsf{pk}$, to produce an “encryption” $s$, for which it has non-negligible advantage at guessing $\mathsf{Dec}_{\mathsf{sk}}(s)$.
We describe informally how to instantiate the primitive in the random oracle model assuming noisy trapdoor claw-free functions (the detailed description is in LABEL:Alg:ObliviousPauliPad). The construction builds on ideas from [BKVV20].
The key idea is the following. Let $f_{0},f_{1}$ be a Trapdoor Claw-Free function pair. We take $\mathsf{pk}=(f_{0},f_{1})$, and $\mathsf{sk}$ to be the corresponding trapdoor. Then, $\mathsf{Enc}_{\mathsf{pk}}$ is as follows:
  • (i)
[…]
  • (ii)
    Compute the random oracle “in the phase”, to obtain $((-1)^{H(x_{0})}\alpha\ket{0}\ket{x_{0}}+(-1)^{H(x_{1})}\beta\ket{1}\ket{x_{1}})$. Measure the second register in the Hadamard basis. This results in a string $d$, and the leftover qubit state
    |$\ket{\psi_{Z}}=Z^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\ket{\psi}\,.$
  • (iii)
    Repeat steps (i) and (ii) on |$\ket{\psi_{Z}}$, but in the Hadamard basis! This results in strings $y^{\prime}$ and $d^{\prime}$, as well as a leftover qubit state
    |$\ket{\psi_{XZ}}=X^{d^{\prime}\cdot(x_{0}^{\prime}\oplus x_{1}^{\prime})+H(x_{0}^{\prime})+H(x_{1}^{\prime})}Z^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\ket{\psi}\,,$
    where $x_{0}^{\prime}$ and $x_{1}^{\prime}$ are the pre-images of $y^{\prime}$.
Notice that the leftover qubit state |$\ket{\psi_{XZ}}$ is of the form $X^{x}Z^{z}\ket{\psi}$ where $x,z$ have the following two properties: (a) a verifier in possession of the TCF trapdoor can learn $z$ and $x$ given respectively $y,d$ and $y^{\prime},d^{\prime}$, and (b) no PPT prover can produce strings $y,d$ as well as predict the corresponding bit $z$ with non-negligible advantage (and similarly for $x$). Intuitively, this holds because a PPT prover that can predict $z$ with non-negligible advantage must be querying the random oracle at both $x_{0}$ and $x_{1}$ with non-negligible probability. By simulating the random oracle (by lazy sampling, for instance), one can thus extract a claw $x_{0},x_{1}$ with non-negligible probability, breaking the claw-free property.
We expect that, by allowing a larger constant number of rounds of interaction, it should be possible to realise the oblivious pad functionality in the plain model.
Our Compiler
We finally describe our contextuality game compiler. As mentioned in the introduction, our strategy is still to ask the first question under $\mathsf{QFHE}$ encryption and the second question in the clear, with the following crucial difference: the prover is first asked to “re-encrypt” the post-measurement state using the oblivious pad functionality (from here one referred to as $\mathsf{OPad}$), and only after that the verifier reveals to the prover how to “decrypt” the state, in order to proceed to round 2.
The key idea is easy to state, once the notation is clear. To this end, recall that Property 3 of a $\mathsf{QFHE}$ scheme ensures that the encryption of a quantum state |$\ket{\psi^{\prime\prime}}$ has the form
$(U_{k^{\prime\prime}}\ket{\psi^{\prime\prime}},\hat{k}^{\prime\prime})$ (3)
where $\hat{k}^{\prime\prime}$ denotes a classical encryption of $k^{\prime\prime}$ (the reason why we use double primes will become clear shortly), and $U_{k^{\prime\prime}}$ is an element of the Pauli group. Note that using the secret key of the $\mathsf{QFHE}$ scheme, one can recover $k^{\prime\prime}$ from $\hat{k}^{\prime\prime}$. Further, let the optimal quantum strategy for the underlying contextuality game consist of state |$\ket{\psi}$ and observables $\{O_{q}\}$. Finally, denote by |$\ket{\psi_{a,q}}$ the post-measurement state arising from measuring |$\ket{\psi}$ using $O_{q}$ and obtaining outcome $a$.
We are now ready to describe our compiler. We first explain it in words and subsequently give a more formal description for clarity. In both cases, we highlight the conceptually new parts in blue.
  • Round 1 Ask the encrypted question, and have the prover re-encrypt its post-measurement state using $\mathsf{OPad}$.
    The verifier samples keys for the $\mathsf{QFHE}$ scheme and for the $\mathsf{OPad}$. It samples a context $C\leftarrow\cal{D}_{\rm contexts}$ and then uniformly samples questions $q_{1},q_{2}$ from the context $C$ (note that $q_{1}=q_{2}$ with probability $1/2$ since, for simplicity, we are considering contexts of size $2$.)
    • Message 1 The verifier sends the $\mathsf{QFHE}$ encryption $c_{q_{1}}$ of the first question $q_{1}$, together with the public key of the $\mathsf{OPad}$.
    The honest quantum prover obtains the encrypted answer $c_{a_{1}}$ by measuring, under the $\mathsf{QFHE}$ encryption, the state |$\ket{\psi}$ using the observable $O_{q_{1}}$. It now holds a state of the form in LABEL:Eq:Property3Form, with |$\ket{\psi^{\prime\prime}}=\ket{\psi_{a_{1},q_{1}}}$. Using the public key of the $\mathsf{OPad}$, the prover applies $\mathsf{OPad}.\mathsf{Enc}$ to the encrypted post-measurement state, $U_{k^{\prime\prime}}\ket{\psi_{a_{1},q_{1}}}$, to obtain the “re-encrypted” quantum state, $U_{k^{\prime}}U_{k^{\prime\prime}}\ket{\psi^{\prime\prime}}$, where $U_{k^{\prime}}$ was applied by the $\mathsf{OPad}$, together with a classical string $s^{\prime}$ that encodes $k^{\prime}$. This step is critical to the security of the protocol and is discussed in more detail shortly. Crucially, note that both $U_{k^{\prime}}$ (coming from the $\mathsf{OPad}$) and $U_{k^{\prime\prime}}$ (coming from the $\mathsf{QFHE}$) are Paulis.
    • Message 2 The prover sends the $\mathsf{QFHE}$ encrypted answer $c_{a_{1}}$, together with the two strings $\hat{k}^{\prime\prime}$ and $s^{\prime}$.
  • Round 2 Remove the overall encryption, and proceed in the clear. The verifier recovers $k^{\prime}$ from $s^{\prime}$ (using the secret key of the $\mathsf{OPad}$) and $k^{\prime\prime}$ from $\hat{k}^{\prime\prime}$ (using the secret key of the $\mathsf{QFHE}$ scheme). It then computes $k$ satisfying $U_{k}=U_{k^{\prime}}U_{k^{\prime\prime}}$.
    • Message 3 The verifier sends the second question $q_{2}$ together with $k$ as computed above.
    The prover measures its quantum state $U_{k^{\prime}}U_{k^{\prime\prime}}\ket{\psi^{\prime\prime}}=U_{k}\ket{\psi_{a_{1},q_{1}}}$, using observable $U_{k}O_{q_{2}}U_{k}^{\dagger}$ to obtain an outcome $a_{2}$.
    • Message 4 The prover sends $a_{2}$.
    The verifier decrypts $c_{a_{1}}$ to recover $a_{1}$ using the secret key of the $\mathsf{QFHE}$ scheme. If $q_{1}=q_{2}$, it accepts if the answers match, i.e. $a_{1}=a_{2}$. If $q_{1}\neq q_{2}$, it accepts if the predicate is true, i.e. $\mathsf{pred}(a_{1},a_{2},q_{1},q_{2})=1$.
Below, we summarise our compiler. Since we now have ${\mathsf{Gen},\mathsf{Enc},\mathsf{Dec}}$ algorithms for both $\mathsf{OPad}$ and $\mathsf{QFHE}$, we use prefixes such as $\mathsf{OPad}.\mathsf{Enc}$ to refer to $\mathsf{Enc}$ associated with $\mathsf{OPad}$ to avoid confusion.
(Honest Quantum) Prover Verifier
Prepare $\left|\psi\right\rangle$ corresponding to the quantum strategy.
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $C\leftarrow\cal{D}_{\rm contexts}$ $q_{1}\leftarrow C$ $q_{2}\leftarrow C$ $c_{q_{1}}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q_{1})$ $(\mathsf{OPad}.\mathsf{sk},\mathsf{OPad}.\mathsf{pk})\leftarrow\mathsf{OPad}.\mathsf{Gen}(1^{\lambda})$
$\xleftarrow{(c_{q_{1}},\mathsf{OPad}.\mathsf{pk})}$
Evaluate the answer to $q_{1}$ under QFHE to get an encrypted answer $c_{a_{1}}$. The post-measurement state is now $\left(U_{k^{\prime\prime}}\left|\psi_{q_{1}a_{1}}\right\rangle,\widehat{k}^{\prime\prime}\right)$ and now apply the $\mathsf{OPad}$ to the post-measurement state, i.e. $(U_{k^{\prime}}U_{k^{\prime\prime}}\left|\psi_{q_{1}a_{1}}\right\rangle,s^{\prime})\leftarrow$ $\mathsf{OPad}.\mathsf{Enc}_{\mathsf{OPad}.\mathsf{pk}}(U_{k^{\prime\prime}}\left|\psi_{q_{1}a_{1}}\right\rangle)$.
$\xrightarrow{(c_{a_{1}},\hat{k}^{\prime\prime},s^{\prime})}$
Using the secret keys, find $k$ such that $U_{k}=U_{k^{\prime}}U_{k^{\prime\prime}}$
$\xleftarrow{(q_{2},k)}$
Measure $U_{k}O_{q_{2}}U_{k}^{\dagger}$
$\overset{a_{2}}{\longrightarrow}$
Decrypt $c_{a_{1}}$ to learn $a_{1}$. If $q_{1}=q_{2}$, accept if $a_{1}=a_{2}$. If $q_{1}\neq q_{2}$, accept if $\mathsf{pred}(a_{1},a_{2},q_{1},q_{2})=1$.
Our compiler satisfies the following, assuming the underlying $\mathsf{QFHE}$ and oblivious pad are secure. We first state a special case of our general result (which is stated later in Theorem 3).
Theorem (restatement of LABEL:Thm:infMain1).
Consider a contextuality game with contexts of size $2$. Let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game (as described above) satisfies the following:
  • (Completeness) The QPT prover described above wins with probability $\frac{1}{2}(1+{\rm valQu})-\mathsf{negl}(\lambda)\,.$
  • (Soundness) Any PPT prover wins with probability at most $\frac{1}{2}(1+{\rm valNC})+\mathsf{negl}(\lambda)\,.$
Here $\mathsf{negl}$ denote (possibly different) negligible functions.
Proof sketch: Suppose ${\cal A}$ is a PPT algorithm that wins with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$. Observe that one can associate a “deterministic assignment” corresponding to ${\cal A}$, conditioned on some fixed first round messages, as follows: simply rewind ${\cal A}$ to learn answers to all possible second round questions, obtaining an assignment $\tau:Q\to A$, mapping questions to answers. Let us write $\tau_{q_{1}}$ to make the dependence of the assignment on the first question more explicit (note that the assignment depends on the encrypted question $c_{q_{1}}$ as well as the encrypted answer $c_{a_{1}}$). For the purpose of this overview, suppose also that ${\cal A}$ is consistent, i.e. if $q_{1}=q_{2}$, then $a_{1}=a_{2}$ (note that this in particular ensures that, when $q_{1}=q_{2}$, ${\cal A}$ wins with probability 1. One can show that an adversary that is not consistent can be turned into an adversary that is consistent and wins with at least the same probability). Now, to win with probability more than $\frac{1}{2}(1+{\rm valNC})$, it must be that the $\tau_{q_{1}}$’s are different for different $q_{1}$’s. Otherwise, ${\cal A}$’s strategy is just a convex combination of deterministic assignments and this by definition cannot do better than ${\rm valNC}$ when $q_{1}\neq q_{2}$. But if the distribution over $\tau_{q_{1}}$s and $\tau_{q_{1}^{\prime}}$s is different for at least some $q_{1}\neq q_{1}^{\prime}$, then one is able to distinguish $\mathsf{QFHE}$ encryptions of $q_{1}$ from those of $q_{1}^{\prime}$. Thus, as long as the $\mathsf{QFHE}$ scheme is secure, no PPT algorithm can win with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$.
In the above sketch, we glossed over the very important subtlety that, in order to obtain the truth table $\tau$, the reduction needs to provide as input to $\cal A$ the “correct” decryption key $k$ (as the verifier does in the third message of our compiled game, where $k$ is such that $U_{k}=U_{k^{\prime}}U_{k^{\prime\prime}}$). However, the reduction only sees encryptions of $k^{\prime}$ and $k^{\prime\prime}$. So, how does it compute $k$ without the secret keys? Crucially, this is where the $\mathsf{OPad}$ comes into play—it allows the reduction to instead use an independent uniformly random $k$ (not necessarily the “correct” one) when constructing the reduction that breaks the security of the $\mathsf{QFHE}$ scheme. The fact that such a $k$ is computationally indistinguishable from the correct one (from the point of view of the prover $\mathcal{A}$) follows precisely from the security of the $\mathsf{OPad}$.
General Compilers
The compiler we described earlier handles contextuality games with contexts of size $2$. How does one generalise it to contexts of arbitrary size? Unlike for KLVY, it is not entirely clear what the “correct” way is here.
We design two compilers (which seem incomparable). The first compiler applies universally to all contextuality games. The second applies primarily to contextuality games where the quantum value is $1$ (for instance, it works for the magic square but not for KCBS). Notably, both compilers are 4-message protocols.
  • $(|C|-1,1)$ compiler:
    • Round 1: Ask $|C|-1$ questions under $\mathsf{QFHE}$.
    • Round 2: Ask $1$ uniformly random question in the clear. If the question was already asked in Round $1$, check consistency. Otherwise, check the predicate.
  • $(|C|,1)$ compiler
    • Round 1: Ask all $|C|$ questions under $\mathsf{QFHE}$.
    • Round 2: Ask $1$ uniformly random question in the clear, and check consistency with the questions asked in Round $1$.
By now, one would be wary of guessing that asking more questions under $\mathsf{QFHE}$ is going to improve the security of the protocol. Indeed, the $(|C|-1,1)$ compiler (which reduces to the one we discussed above for $|C|=2$) is the universal one. We show the following.
Theorem 3 (informal).
Consider an arbitrary contextuality game, and let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game, obtained via the $(|C|-1,1)$ compiler, satisfies the following:
  • (Completeness) There is a QPT prover that wins with probability $1-\frac{1}{|C|}+\frac{{\rm valQu}}{|C|}-\mathsf{negl}$.
  • (Soundness) PPT provers win with probability at most $1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\mathsf{negl}$.
The compiled game, obtained via the $(|C|,1)$ compiler, satisfies the following:
  • (Completeness) There is a QPT prover that wins with probability ${\rm valQu}$.
  • (Soundness) PPT provers win with probability at most $1-\mathsf{const}_{1}+\mathsf{negl}$, where $\mathsf{const}_{1}=\min_{C\in C^{{\rm all}}}\frac{\Pr(C)}{|C|}$ (this is constant in the sense that it is independent of the security parameter), and $\Pr(C)$ denotes the probability of sampling the context $C$.
Here, $\mathsf{negl}$ are (possibly different) negligible functions.
We make some brief remarks about the two compilers and defer the details to the main text.
  • The $(|C|,1)$ compiler is not universal because, for instance, when applied to KCBS, there is no gap between the PPT and the QPT prover’s winning probabilities. In fact, there is a PPT algorithm [^16] Assuming that $\mathsf{Eval}$ is PPT if the circuit and input are classical. that does better than the honest quantum strategy. Yet, the compiler does apply to the magic square game, for instance, because $\mathsf{const}_{1}<1$ and ${\rm valQu}=1$. In fact, for the magic square game, this compiler gives a better completeness-soundness gap than the $(|C|-1,1)$ compiler.
  • The $(|C|-1,1)$ compiler is universal in the sense that, when applied to any contextuality game with a gap between non-contextual and quantum value, the compiled game will have a constant gap between completeness and soundness. However, the resulting gap is sometimes smaller compared to the previous compiler. It is unclear if one can do better than this, with or without increasing the number of rounds, while preserving universality.
2.6 Contribution 2 — An even simpler proof of quantumness
Our proof of quantumness, like many of the existing ones in the literature, is based on the use of Trapdoor Claw-Free Functions (TCF). In our protocol, these are used to realize an “encrypted CNOT” functionality, which is the central building block of Mahadev’s QFHE scheme [Mah20]. The “encrypted CNOT” functionality allows a prover to homomorphically apply the gate $\mathsf{CNOT}^{a}$, while holding a (classical) encryption of the bit $a$. Formally, our protocol uses Noisy Trapdoor Claw-Free Functions (NTCF, defined formally in Definition 5), but here we describe our scheme using regular TCFs for simplicity.
The proof of quantumness
[…]
    The verifier samples a bit $a$ uniformly at random.
    • Message 1: The verifier sends an appropriate encryption of $a$ to the prover (and holds on to the corresponding secret key).
    The honest prover prepares the two-qubit state |+|$\ket{+}\ket{0}$, along with auxiliary registers required to perform an “encrypted CNOT” operation. It then performs an “encrypted CNOT” operation (from the first qubit to the second), i.e. homomorphically applies $\mathsf{CNOT}^{a}$, followed by a measurement of the second (logical) qubit.
    • Message 2: The prover sends all measurement outcomes to the verifier.
    Since a CNOT gate can be thought of as a deferred measurement in the standard basis, we can equivalently think of the prover’s operations as performing an “encrypted measurement” of the first qubit, where the first qubit is being measured or not based on the value of $a$. Note that, after having performed these operations, the prover holds a single qubit. Thanks to the specific structure of the “encrypted CNOT” operation from [Mah20], the resulting “post-measurement” qubit state is encrypted with a Quantum One-Time Pad, and is either:
    • $\ket{+}$ or $\ket{-}$, if $a=0$ (i.e. no logical CNOT was performed)
    • |$\ket{0}$ or |$\ket{1}$, if $a=1$ (i.e. a logical CNOT was performed)
[…]
  • Round 2: Ask the prover to perform “Bob’s CHSH measurement”.
[…]
    • Message 3: The verifier sends a uniformly random bit $b$ to the prover (corresponding to Bob’s question in a CHSH game).
    The prover performs Bob’s optimal CHSH measurement corresponding to question $b$.
    • Message 4: The prover returns the measurement outcome to the verifier.
    The verifier checks that the corresponding CHSH game is won.
In LABEL:Fig:prover_circuit, we show the circuit for the honest quantum prover in our proof of quantumness.
[…]
Figure 6: The prover’s circuit. Here, $(f_{0},f_{1})$ is a pair of trapdoor claw-free functions (with inputs of size $n$). $U_{f_{0},f_{1}}$ denotes the $(n+1)$-qubit unitary that coherently computes $f_{0}$ in the last $n$ qubits if the first qubit is |$\ket{0}$, and computes $f_{1}$ otherwise. The circuit starts by preparing $\ket{+}$ in the first qubit, and a uniform superposition over all inputs in the next $n$ qubits. The circuit then applies $U_{f_{0},f_{1}}$ (note that we are omitting auxiliary work registers that are required to compute $U_{f_{0},f_{1}}$), followed by a layer of Hadamard gates on the last $n-1$ qubits. Then, the last $n$ qubits are measured. As a result, the leftover qubit |$\ket{\psi}$ in the first register is now a BB84 state (which one it is depends on $f_{0}$, $f_{1}$, and the measurement outcome). As its second message, the verifier sends a bit $b$, and the prover applies the rotation $R_{b}$ defined as follows: |$\ket{0}\overset{R_{b}}{\mapsto}\cos((-1)^{b}\pi/8)\ket{0}+\sin((-1)^{b}\pi/8)\ket{1}$ and |$\ket{1}\overset{R_{b}}{\mapsto}-\sin((-1)^{b}\pi/8)\ket{0}+\cos((-1)^{b}\pi/8)\ket{1}$. Finally, the prover measures the qubit in the standard basis. [^18] The simplified description of the proof of quantumness before this figure is slightly inaccurate: it states that the prover prepares starts by preparing the two-qubit state |+|$\ket{+}\ket{0}$. Technically, the prover only needs to prepare $\ket{+}$, and the role of the second qubit is performed by the first qubit of the pre-image register (which is initialised as a uniform superposition). We refer to Section 4 for a detailed description.
Soundness
An efficient quantum prover can efficiently pass this test with probability $\cos^{2}(\frac{\pi}{8})\approx 0.85$, while an efficient classical prover can pass this test with probability at most $3/4$. The proof of classical soundness is fairly straightforward. In essence, a classical prover can be rewound to obtain answers to both of the verifier’s possible questions in Message 3. If the classical prover passes the test with probability $3/4+\delta$ (which is on average over the two possible questions), then the answers to both questions together must reveal some information about the encrypted bit $a$ (this is a simple consequence of how the CHSH winning conditions is defined). In particular, such a classical prover can be used to guess $a$ with probability $\frac{1}{2}+2\delta$. This breaks the security of the encryption, as long as $\delta$ is non-negligible. We defer the reader to Section 5.2 for more details.
Putting the ideas in perspective
We have already discussed in Section 1.1 how our proof of quantumness compares to existing ones in terms of efficiency. Here, we focus on how our proof of quantumness compares conceptually to [KLVY23] and [KMCVY22]:
  • In [KLVY23], the prover is asked to create an entangled EPR pair, of which the first half is encrypted, and the second half is in the clear. Then, the prover is asked to perform Alice’s ideal CHSH measurement homomorphically on the first half, and Bob’s CHSH measurement in the clear on the second half. Our proof of quantumness departs from this thanks to two observations:
    • By leveraging the structure of the “encrypted CNOT operation” from [Mah20], the post-measurement state from Alice’s homomorphic measurement can be re-used in the clear (precisely because the verifier knows what the state is, but the prover does not). So the initial entanglement is not needed. This idea is also the starting point for our contextuality compiler from Section 2.5, although for the latter we take this idea much further: we find a way to give the prover the ability to decrypt the leftover state without giving up on soundness. Our proof of quantumness is a baby version of this idea: it leverages the fact that the leftover encrypted state has a special form, namely it is a BB84 state.
    • In order to setup a “CHSH-like correlation” between the verifier and the leftover qubit used by the prover in Round 2, one does not need to compile the CHSH game in its entirety. This compilation, even for the simple CHSH game requires the prover to perform an “encrypted controlled-Hadamard” operation (because Alice’s ideal CHSH measurements are in the standard and Hadamard bases). The latter requires three sequential “encrypted CNOT” operations. Instead, our observation is that one can setup this CHSH-like correlation more directly, as we do in Round 1 of our proof of quantumness.
  • From a different point of view, our proof of quantumness can also be viewed as a simplified version of [KMCVY22]. Indeed, observation (ii) is inspired by the proof of quantumness in [KMCVY22], which introduces the idea of a “computational” CHSH test. One can interpret [KMCVY22] as setting up an “encrypted classical operation”, akin to an “encrypted CNOT”, that either entangles two registers or does not, ultimately having the effect of performing an “encrypted measurement”. This is achieved via an additional round of interaction. Our proof of quantumness can be thought of as zooming in on this interpretation, and finding a direct way to achieve this without the additional interaction.
3 Preliminaries
LABEL:Subsec:notation sets up some notation, LABEL:Subsec:FHE-Scheme formally introduces QFHE, and LABEL:Subsec:NTCF formally introduces trapdoor-claw free functions.
3.1 Notation
  • For mixed state $\rho$ and $\sigma$, we write $\rho\approx_{\epsilon}\sigma$ to mean that $\rho$ and $\sigma$ are at most $\epsilon$-far in trace distance, i.e. $\frac{1}{2}{\rm tr}(|\rho-\sigma|)\leq\epsilon$.
  • Let $X$ be the Pauli $X$ matrix. For a string $x\in\{0,1\}^{n}$, we write $X^{x}$ to mean $\otimes_{i=1}^{n}X^{x_{i}}$. We use a similar notation for Pauli $Z$.
  • We denote the spectrum of an observable $O$ by ${\mathsf{Spectr}}(O)$ and the support of a function $f$ by $\mathsf{Supp}(f)$.
  • We use the abbreviations PPT and QPT for probabilistic polynomial time and quantum polynomial time algorithms respectively.
  • Vector/list indexed by a set.
    Let $S$ and $V$ be finite sets. Let $\mathbf{v}$ be a vector/list with entries in $V$, indexed by $S$, i.e. $\mathbf{v}$ contains a value in $V$ for each $s\in S$. We denote by $\mathbf{v}[s]\in V$ the value corresponding to $s$.
    For a subset $C\subseteq S$, we write $\mathbf{v}[C]$ for the vector obtained by restricting the indices of $\mathbf{v}$ to the set $C$. For example, if $\mathbf{v^{\prime}}$ is a vector indexed by $C$, then, by $\mathbf{v}[C]=\mathbf{v^{\prime}}$ we mean that $\mathbf{v}[s]=\mathbf{v^{\prime}}[s]$ for all $s\in C$.
  • Asymptotic notation.
    • Big-O. Let $f,g:\mathbb{N}\to\mathbb{R}$. We write $f\leq O(g)$ if $\exists c,n_{0}$ such that for all $n\geq n_{0}$, $f(n)\leq cg(n)$.
    • Big-O on $\Lambda$. Take $\Lambda\subseteq\mathbb{N}$ to be an infinite subset of $\mathbb{N}$. Then, by $f\leq O(g)$ on $\Lambda$, we mean that there are $c,n_{0}$ such that for all $n\geq n_{0}$ in $\Lambda$, $f(n)\leq cg(n)$.
    We define Big-$\Omega$ on $\Lambda$ as a similar generalisation of the Big-$\Omega$ notation.
3.2 QFHE Scheme
We formally define the notion of Quantum Homomorphic Encryption that we will employ in the rest of the paper. Note that, as in [KLVY23] we only require security to hold against PPT provers.
Definition 4 (Quantum Homomorphic Encryption).
A Quantum Homomorphic Encryption scheme for a class of circuits $\mathcal{C}$ is a tuple of algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec},\mathsf{Eval})$ with the following syntax:
  • $\mathsf{Gen}$ is PPT. It takes a unary input $1^{n}$, and outputs a secret key $\mathsf{sk}$.
  • $\mathsf{Enc}$ is QPT. It takes as input a secret key $\mathsf{sk}$ and a quantum state $\left|\psi\right\rangle$, and outputs a ciphertext |$\ket{c}$. We require that if |$\ket{\psi}$ is classical, i.e. a standard basis state, then $\mathsf{Enc}$ becomes a PPT algorithm. In particular, $\left|c\right\rangle$ is also classical.
  • $\mathsf{Eval}$ is QPT. It takes as input a tuple $(C,\ket{\xi},\ket{c})$, where
    1. 1.
      $C:\mathcal{H}\times(\mathbb{C}_{2})^{\otimes n}\to(\mathbb{C}_{2})^{\otimes m}$ is a quantum circuit in $\mathcal{C}$,
    2. 2.
      |$\ket{\xi}\in\mathcal{H}$ is a quantum state,
    3. 3.
      |$\ket{c}$ is a ciphertext encrypting an $n$-qubit state.
    $\mathsf{Eval}$ computes a quantum circuit $\mathsf{Eval}_{C}(\ket{\xi},\ket{c})$, which outputs a ciphertext |$\ket{c_{\mathsf{out}}}$. We require that, if $C$ has classical output, then $\mathsf{Eval}_{C}$ also has classical output.
  • $\mathsf{Dec}$ is QPT. It takes as input a secret key $\mathsf{sk}$ and a ciphertext |$\ket{c}$, and outputs a state $\left|\phi\right\rangle$. If |$\ket{c}$ is classical, then |$\ket{\phi}$ is also classical.
The correctness and security conditions are as follows:
  1. 1.
    Correctness: For any quantum circuit $C:\mathcal{H}\times(\mathbb{C}_{2})^{\otimes n}\to(\mathbb{C}_{2})^{\otimes m}$ in $\mathcal{C}$, there is a negligible function $\mathsf{negl}$, such that the following holds. For any quantum state |$\ket{\xi}\in\mathcal{H}$, any $n$-qubit state |$\ket{\psi}$, any $\lambda\in\mathbb{N}$, any $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$, and any |$\ket{c}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(\ket{\psi})$, the following two states are $\mathsf{negl}(\lambda)$-close in trace distance.
    • $C(\left|\xi\right\rangle\otimes\left|\psi\right\rangle)$,
    • $\mathsf{Dec}_{\mathsf{sk}}(\mathsf{Eval}_{C}(\left|\xi\right\rangle,\ket{c}))$.
  2. 2.
    Security: For all PPT distinguishers $D$, for all polynomial functions $\mathsf{poly}$, there exists a negligible function $\mathsf{negl}$ such that, for any two strings $x_{0},x_{1}\in\{0,1\}^{\mathsf{poly}(\lambda)}$, for all $\lambda$,
    $\left|\Pr\left[D(\ket{c_{0}})=1:\begin{array}[]{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ \ket{c_{0}}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(x_{0})\end{array}\right]-\Pr\left[D(\ket{c_{1}})=1:\begin{array}[]{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ \ket{c_{1}}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(x_{1})\end{array}\right]\right|\leq\mathsf{negl}(\lambda)\,.$ (4)
A Quantum Fully Homomorphic Encryption scheme ($\mathsf{QFHE}$) is a Quantum Homomorphic Encryption scheme for the class of all poly-size quantum circuits.
Form of encryption.
For the rest of this work, we restrict to QFHE schemes where encryption takes the following form. For an $n$-qubit state |$\ket{\psi}$,
$(U_{k}\ket{\psi},\hat{k})\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(\psi)\,,$ (5)
where $\{U_{k}\}_{k\in K}$ forms a group (potentially up to global phases) and $K(n)$ is a finite set. Furthermore, $\hat{k}$ is a classical encryption of $k$, which can be decrypted using a PPT algorithm (specified by the $\mathsf{QFHE}$ scheme) given $\mathsf{sk}$.
All known $\mathsf{QFHE}$ schemes satisfying Definition 4 [Mah20, Bra18], also satisfy the property above with $U_{k}$ being a Pauli pad and $k=(x,z)\in\{0,1\}^{2n}$ encoding which Pauli operator is applied, i.e. $U_{k}=X^{x}Z^{z}$.
3.3 Noisy Trapdoor Claw-Free Functions
Our work requires trapdoor claw-free functions for two purposes:
  • They are used to instantiate Mahadev’s QFHE scheme [Mah20].
  • They are used to construct the oblivious pad.
However, the only known constructions of TCFs that satisfy the properties needed to instantiate QFHE (e.g. the property of Definition 6) are “noisy” versions (NTCFs), and they are based on the hardness of LWE. We define NTCFs next. Before doing so, we point out that, while an NTCF is needed to instantiate QFHE based on current knowledge, any TCF (even based on quantum insecure assumptions, like Diffie-Hellman) suffices to build the oblivious pad (however, for simplicity, we still use NTCFs in Section 7).
For readers familiar with the area, note that the “adaptive hardcore bit” property is not required for any of the constructions in this work.
Definition 5 (NTCF family; paraphrased from [BCM+21].).
Let $\lambda$ be a security parameter. Let $\mathcal{X}$ and $\mathcal{Y}$ be finite sets and $\mathcal{K}$ be a finite set of keys (these sets implicitly depend on $\lambda$). A family of functions
=$\mathcal{F}=\{f_{\mathsf{pk},b}:\mathcal{X}\to\mathcal{D}_{\mathcal{Y}}\}_{\mathsf{pk}\in\mathcal{K},b\in\{0,1\}}$
is called a noisy trapdoor claw free function (NTCF) family if the following conditions hold (for each $\lambda$):
  1. 1.
    Efficient Function Generation: There exists an efficient probabilistic algorithm $\mathsf{Gen}$ that generates a (public) key in $\mathcal{K}$ together with a trapdoor (secret key) $\mathsf{sk}$:
    $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda}).$
  2. 2.
    Trapdoor Injective Pair:
    1. (a)
      Trapdoor: There exists an efficient deterministic algorithm $\mathsf{Inv}$ such that with overwhelming probability over the choice of $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$, the following holds:
      $\mathsf{Inv}(\mathsf{sk},b,y)=x$
      for all $b\in\{0,1\},x\in\mathcal{X}$ and $y\in\mathsf{Supp}(f_{\mathsf{pk},b}(x))$.
    2. (b)
      Injective pair: For all keys $\mathsf{pk}\in\mathcal{K}$, there exists a perfect matching $\mathcal{R}_{\mathsf{pk}}\subseteq\mathcal{X}\times\mathcal{X}$ such that $f_{\mathsf{pk},0}(x_{0})=f_{\mathsf{pk},1}(x_{1})$ if and only if $(x_{0},x_{1})\in\mathcal{R}_{\mathsf{pk}}$. Such a pair $(x_{0},x_{1})$ is referred to as a “claw”.
  3. 3.
    Efficient Range Superposition. For all keys $\mathsf{pk}\in\mathcal{K}$ and $b\in\{0,1\}$ there exists a function $f^{\prime}_{\mathsf{pk},b}:\mathcal{X}\to\mathcal{D}_{\mathcal{Y}}$ such that the following holds:
    1. (a)
      $\mathsf{Inv}(\mathsf{sk},b,y)=x_{b}$ and $\mathsf{Inv}(\mathsf{sk},b\oplus 1,y)=x_{b\oplus 1}$ for all $(x_{0},x_{1})\in\mathcal{R}_{\mathsf{pk}}$ and $y\in\mathsf{Supp}(f^{\prime}_{\mathsf{pk},b}(x_{b}))$.
    2. (b)
      There is an efficient deterministic procedure $\mathsf{Chk}$ s.t. for all $b\in\{0,1\}$, $\mathsf{pk}\in\mathcal{K}$, $x\in\mathcal{X}$ and $y\in\mathcal{Y}$,
      $\mathsf{Chk}(\mathsf{pk},b,x,y)=\begin{cases}1&\text{if }y\in\mathsf{Supp}(f^{\prime}_{\mathsf{pk},b}(x))\\ 0&\text{else}.\end{cases}$
      Observe that $\mathsf{Chk}$ is not provided the secret trapdoor $\mathsf{sk}$.
    3. (c)
      For each $\mathsf{pk}\in\mathcal{K}$ and $b\in\{0,1\}$, it holds that
      $\mathbb{E}_{x\leftarrow\mathcal{X}}[H^{2}(f_{\mathsf{pk},b}(x),f^{\prime}_{\mathsf{pk},b}(x))]\leq\mathsf{negl}(\lambda)$ (6)
      for some negligible function $\mathsf{negl}$, where $H^{2}$denotes the Hellinger distance.
      Further, there is an efficient procedure $\mathsf{Samp}$ that on input $\mathsf{pk}$ and $b\in\{0,1\}$ prepares the state
      $\frac{1}{\sqrt{|\mathcal{X}|}}\sum_{x\in\mathcal{X},y\in\mathcal{Y}}\sqrt{(f^{\prime}_{\mathsf{pk},b}(x))(y)}\left|x\right\rangle\left|y\right\rangle.$
  4. 4.
    Claw-Free Property. For any PPT adversary $\mathcal{A}$,
    $\Pr\left[(x_{0},x_{1})\in\mathcal{R}_{\mathsf{pk}}:\begin{array}[]{c}(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})\\ (x_{0},x_{1})\leftarrow\mathcal{A}(\mathsf{pk})\end{array}\right]\leq{\mathsf{negl}}(\lambda).$
When ${\cal F}$ is used with other primitives, we refer to the algorithms $(\mathsf{Gen},\mathsf{Inv},\mathsf{Samp},\mathsf{Chk})$ and the various sets $({\cal X},{\cal D}_{{\cal Y}},{\cal Y})$ with ${\cal F}$ prefixed, e.g. $\mathsf{Gen}(1^{\lambda})$ is referred to as .${\cal F}.\mathsf{Gen}(1^{\lambda})$.
We define an additional property of NTCFs, which we will invoke directly in the soundness analysis of our proof of quantumness in Section 5.2.
Definition 6 (“Hiding a bit in the xor”).
Let =$\mathcal{F}=\{f_{\mathsf{pk},b}:\mathcal{X}\to\mathcal{D}_{\mathcal{Y}}\}_{\mathsf{pk}\in\mathcal{K},b\in\{0,1\}}$ be an NTCF family (as in Definition 5). We say that $\mathcal{F}$ “hides a bit in the xor” if $\mathsf{Gen}$ takes an additional bit of input $s$:
$(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda},s)\,,$
such that the following holds (in addition to the properties already satisfied by $\mathsf{pk}$ and $\mathsf{sk}$ in Definition 5). For all $s\in\{0,1\}$, if $(\mathsf{pk},\mathsf{sk})$ is in the support of $\mathsf{Gen}(1^{\lambda},s)$, then $s=x_{0}[1]\oplus x_{1}[1]$ for all $(x_{0},x_{1})\in\mathcal{R}_{\mathsf{pk}}$, where $x_{0}[1]$ and $x_{1}[1]$ denote the first bits of $x_{0}$ and $x_{1}$ respectively, and $\mathcal{R}_{\mathsf{pk}}$ is as in Definition 5. Moreover, for all QPT algorithms $\mathcal{A}$, there exists a negligible function $\mathsf{negl}$ such that, for all $\lambda$,
$\Pr[s\leftarrow\mathcal{A}(\mathsf{pk}):(\mathsf{pk},\mathsf{sk})\leftarrow{\cal F}.\mathsf{Gen}(1^{\lambda},s)\,\,,\,\,s\leftarrow\{0,1\}]\leq\mathsf{negl}(\lambda)\,.$ (7)
The following theorem is implicit in [Mah20].
Theorem 7 ([Mah20]).
There exists an NTCF family with the property of Definition 6, assuming the quantum hardness of the Learning With Errors (LWE) problem.
Part I Even Simpler Proof of Quantumness
4 The Proof of Quantumness
For an informal description of our proof of quantumness, we refer the reader to Section 2.6. Here we provide a formal description. Our proof of quantumness makes use of a NTCF family =$\mathcal{F}=\{f_{\mathsf{pk},b}:\mathcal{X}\to\mathcal{D}_{\mathcal{Y}}\}_{\mathsf{pk}\in\mathcal{K},b\in\{0,1\}}$ (as in Definition 5) that additionally satisfies the property of Definition 6, i.e. a claw-free function pair hides a bit in the xor of the first bit of any claw.
We parse the domain $\mathcal{X}$ as $\mathcal{X}=\{0,1\}\times\mathcal{V}$. Then, a bit more precisely, the property of Definition 6 says the following. Let $s\in\{0,1\}$, $(\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)$. Let $(\mu_{0},v_{0})$ and $(\mu_{1},v_{1})$ be such that $f_{\mathsf{pk},0}(\mu_{0},v_{0})=f_{\mathsf{pk},1}(\mu_{1},v_{1})$. Then $\mu_{0}\oplus\mu_{1}=s$.
Our proof of quantumness invokes the procedure .$\mathcal{F}.\mathsf{Samp}$ from the “Efficient Range Superposition” property in Definition 5. For simplicity, we describe our proof of quantumness assuming that the “Efficient Range Superposition” property holds exactly, i.e. the RHS of Equation (6) is zero. The actual procedure .$\mathcal{F}.\mathsf{Samp}$ is indistinguishable from the exact one, up to a negligible distinguishing advantage in the security parameter.
Let $\lambda\in\mathbb{N}$ be a security parameter.
  • Message 1: The verifier samples $s\leftarrow\{0,1\}$. Then, she samples $(\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)$. The verifier sends $\mathsf{pk}$ to the prover.
  • Message 2: The prover uses .$\mathcal{F}.\mathsf{Samp}$ to create the state
    $\frac{1}{\sqrt{2|\mathcal{X}|}}\sum_{b\in\{0,1\},x\in\mathcal{X},y\in\mathcal{Y}}\sqrt{f_{\mathsf{pk}}(x)(y)}\ket{b}\ket{x}\ket{y}$ (8)
    This state can be created by preparing a qubit in the state $\ket{+}$ (and a sufficiently large auxiliary register), and then running the .$\mathcal{F}.\mathsf{Samp}$ controlled on the first qubit.
    For $b\in\{0,1\}$, let $(\mu^{y}_{b},v_{b}^{y})$ be such that $y\in\mathsf{Supp}(\mu^{y}_{b},v_{b}^{y})$ (this is a unique element by the properties of the NTCF). Then, we can rewrite the state as:
    $\frac{1}{\sqrt{2|\mathcal{X}|}}\sum_{b\in\{0,1\}}\sqrt{f_{\mathsf{pk}}(\mu^{y}_{0},v_{0}^{y})(y)}\ket{b}\ket{\mu^{y}_{b},v_{b}^{y}}\ket{y}$
    $=$ $\frac{1}{\sqrt{2|\mathcal{X}|}}\sum_{b\in\{0,1\}}\sqrt{f_{\mathsf{pk}}(\mu^{y}_{0},v_{0}^{y})(y)}\ket{b}\ket{\mu^{y}_{0}\oplus b\cdot s,v_{b}^{y}}\ket{y}$
    The prover applies Hadamard gates to every qubit in the “$v_{b}^{y}$” register, and then measures all qubits except the first. Let the output be $(\mu,d,y)$. Then, the resulting post-measurement state of the first qubit is (up to global phases):
    {|$\begin{cases}\ket{0}+(-1)^{d\cdot(v^{y}_{0}+v^{y}_{1})}\ket{1}\quad\textnormal{ if }s=0\\ \ket{\mu\oplus\mu_{0}^{y}}\,\,\,\,\,\quad\quad\quad\quad\quad\quad\textnormal{if }s=1\end{cases}$ (9)
    The prover returns $(\mu,d,y)$ to the verifier.
  • Message 3: The verifier samples $c\leftarrow\{0,1\}$, and sends $c$ to the prover.
  • Message 4:
    • If $c=0$, the prover measures the qubit in the basis
      $\{\cos(\frac{\pi}{8})\ket{0}+\sin(\frac{\pi}{8})\ket{1},-\sin(\frac{\pi}{8})\ket{0}+\cos(\frac{\pi}{8})\ket{1}\}\,.$
    • If $c=1$, the prover measures in the basis
      $\{\cos(-\frac{\pi}{8})\ket{0}+\sin(-\frac{\pi}{8})\ket{1},-\sin(-\frac{\pi}{8})\ket{0}+\cos(-\frac{\pi}{8})\ket{1}\}\,.$
    The prover returns the outcome $b$ to the verifier.
  • Verifier’s final computation:
    • If $s=0$, the verifier runs $(\mu_{0}^{y},v_{0}^{y})\leftarrow\mathcal{F}.\mathsf{Inv}(\mathsf{sk},0,y)$ and $(\mu_{1}^{y},v_{1}^{y})\leftarrow\mathcal{F}.\mathsf{Inv}(\mathsf{sk},1,y)$. She sets $a=d\cdot(v_{0}^{y}\oplus v_{1}^{y})$. Finally, she outputs accept if $a\oplus b=c$, and $\mathsf{reject}$ otherwise.
    • If $s=1$, the verifier runs $(\mu_{0}^{y},v_{0}^{y})\leftarrow\mathcal{F}.\mathsf{Inv}(\mathsf{sk},0,y)$. She sets $a=\mu\oplus\mu_{0}^{y}$. Finally, she outputs accept if $a\oplus b=0$, and $\mathsf{reject}$ otherwise.
Algorithm 1 (Proof of Quantumness)
5 Analysis
5.1 Correctness
Correctness as stated below is straightforward to verify.
Theorem 8 (Correctness).
There exists a QPT algorithm $\mathcal{A}$ and a negligible function $\mathsf{negl}$ such that, for all $\lambda$,
[…]
Proof.
The QPT algorithm $\mathcal{A}$ follows the steps of the prover in LABEL:Alg:simplerPoQ. Then, correctness follows from the fact that the state in Equation (9) is one of the four BB84 states, and a straightforward calculation. One can also realize that the prover’s measurement is Bob’s ideal CHSH measurement corresponding to question $c$, and that the verifier is precisely checking the appropriate CHSH winning condition based on $s$. The negligible loss in the correctness probability comes from the fact that, as we mentioned earlier, the procedure .$\mathcal{F}.\mathsf{Samp}$ actually generates a state that is only negligibly close to that of Equation (8). ∎
5.2 Soundness
Theorem 9.
For any PPT prover $\mathcal{A}$, there exists a negligible function $\mathsf{negl}$ such that, for all $\lambda$,
[…]
Proof.
We show this by giving a reduction from a prover $\mathcal{A}$ for LABEL:Alg:simplerPoQ to an adversary $\mathcal{A}^{\prime}$ breaking the property of Definition 6 satisfied by $\mathcal{F}$, i.e. predicting the bit $s$ “hidden in the claw-free pair”. Specifically, we show that if $\mathcal{A}$ wins with probability $\frac{3}{4}+\delta$ for some $\delta\geq 0$, then $\mathcal{A}^{\prime}$ can guess the the bit $s$ from Equation (7) with probability at least $\frac{1}{2}+2\delta$. $\mathcal{A}^{\prime}$ proceeds as follows, for security parameter $\lambda$:
  • (i)
    $\mathcal{A}^{\prime}$ receives $\mathsf{pk}$ (where $\mathsf{pk}$ is sampled according to $(\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)$ for some uniformly random $s$).
  • (ii)
    $\mathcal{A}^{\prime}$ runs $\mathcal{A}$ on “first message” $\mathsf{pk}$. Let $z$ be the output.
  • (iii)
    Continue running $\mathcal{A}$ on “second message” $c=0$. Let $b_{0}$ be the received output.
  • (iv)
    Rewind $\mathcal{A}$ to just before step (iii). Run $\mathcal{A}$ on “second message” $c=1$. Let $b_{1}$ be the received output.
  • (v)
    Output the guess $s^{\prime}=b_{0}\oplus b_{1}$.
We show that $\Pr[\mathcal{A}^{\prime}\textnormal{ wins}]\geq\frac{1}{2}+2\delta$. Using the notation introduced above, first notice that when $b_{0}$ is a response that a Verifier from LABEL:Alg:simplerPoQ would accept, we have, by construction, that
$a\oplus b_{0}=0\,,$ (10)
where $a$ is defined as in LABEL:Alg:simplerPoQ (given $\mathcal{A}$’s first response $z$, and $\mathsf{sk}$). We refer to such a $b_{0}$ as “valid”. Similarly, if $b_{1}$ is valid, then, by construction,
$a\oplus b_{1}=s\,.$ (11)
Summing Equations (10) and (11) together gives $b_{0}\oplus b_{1}=s$ (where $s$ is the bit that was actually sampled in (i)). Since $\mathcal{A^{\prime}}$ outputs precisely $s^{\prime}=b_{0}\oplus b_{1}$, we have that
$\Pr[\mathcal{A}^{\prime}\textnormal{ wins}]=\Pr[s^{\prime}=s]\geq\Pr[b_{0}\textnormal{ and }b_{1}\textnormal{ are valid}]\,.$ (12)
Now, for fixed $\mathsf{pk}$ and $z$, define
$p^{\mathsf{pk},z}_{win,0}:=\Pr[b_{0}\textnormal{ is valid}\,|\,\mathsf{pk},z]\,,$
and define $p^{\mathsf{pk},z}_{win,1}$ analogously using $b_{1}$. Then, we have
$\Pr[\mathcal{A}^{\prime}\textnormal{ wins}]$ $\geq\Pr[b_{0}\textnormal{ and }b_{1}\textnormal{ are valid}]$
$=\mathbb{E}_{\begin{subarray}{c}s\leftarrow\{0,1\}\\ (\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)\\ z\leftarrow\mathcal{A}(\mathsf{pk})\end{subarray}}\,\,\left[p^{\mathsf{pk},z}_{win,0}\cdot p^{\mathsf{pk},z}_{win,1}\right]$
$\geq\mathbb{E}_{\begin{subarray}{c}s\leftarrow\{0,1\}\\ (\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)\\ z\leftarrow\mathcal{A}(\mathsf{pk})\end{subarray}}\left[p^{\mathsf{pk},z}_{win,0}+p^{\mathsf{pk},z}_{win,1}-1\right]$ (13)
$=2\cdot\mathbb{E}_{\begin{subarray}{c}s\leftarrow\{0,1\}\\ (\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)\\ z\leftarrow\mathcal{A}(\mathsf{pk})\end{subarray}}\left[\frac{1}{2}\left(p^{\mathsf{pk},z}_{win,0}+p^{\mathsf{pk},z}_{win,1}\right)\right]-1\,,$ (14)
where (13) follows from the inequality $x\cdot y\geq x+y-1$ for all $x,y\in[0,1]$ (subtracting $x$ from both sides makes this inequality apparent).
Now, notice that
$\mathbb{E}_{\begin{subarray}{c}s\leftarrow\{0,1\}\\ (\mathsf{pk},\mathsf{sk})\leftarrow\mathcal{F}.\mathsf{Gen}(1^{\lambda},s)\\ z\leftarrow\mathcal{A}(\mathsf{pk})\end{subarray}}\left[\frac{1}{2}\left(p^{\mathsf{pk},z}_{win,0}+p^{\mathsf{pk},z}_{win,1}\right)\right]=\Pr[\mathcal{A}\textnormal{ wins}]\,.$ (15)
Thus, plugging this into (14), gives
$\Pr[\mathcal{A}^{\prime}\textnormal{ wins}]\geq 2\cdot\Pr[\mathcal{A}\textnormal{ wins}]-1=2\cdot\left(\frac{3}{4}+\delta\right)-1=\frac{1}{2}+2\delta\,,$ (16)
as desired.
Part II A Computational Test of Contextuality—for size 2 contexts
6 Contextuality Games
We start by defining contextuality games. This section uses the list/vector notation from LABEL:Subsec:notation.
Definition 10 (Contextuality game, strategy, value).
A contextuality game $\mathsf{G}$ is specified by a tuple $(Q,A,C^{{\rm all}},{\rm pred},{\cal D})$ where
  • $Q$ is a set of questions.
  • $A$ is a set of answers.
  • $C^{{\rm all}}$ is a set of subsets of $Q$. We refer to an element of $C^{{\rm all}}$ as a context.
  • ${\cal D}$ is a distribution over contexts in $C^{{\rm all}}$, and
  • ${\rm pred}$ is a binary-valued function. It takes as input pairs of the form $(\mathbf{a},C)$, where $C\in C^{{\rm all}},$ and $\mathbf{a}\in A^{|C|}$. We think of $\mathbf{a}$ as being indexed by the questions in $C$, and we write $\mathbf{a}[q]\in A$ to represent the answer to question $q\in C$.
$\mathsf{G}$ can be thought of as a 2-message game between a referee and a player that proceeds as follows:
  1. (i)
    The referee samples a context $C\leftarrow{\cal D}$ and sends $C$ to the prover.
  2. (ii)
    The player responds with $\mathbf{a}\in A^{|C|}$.
  3. (iii)
    The referee accepts if ${\rm pred}(\mathbf{a},C)=1$.
A strategy for the game $\mathsf{G}$ is specified by a family $P$ of probability distributions (one for each $C\in C^{{\rm all}}$), where $P[\mathbf{a}|C]$ can be thought of as the probability of answering $\mathbf{a}$ on questions in $C$.
We define the value of $\mathsf{G}$ with respect to a strategy $P$ to be
${\rm val}(P):=\sum_{C\in C^{{\rm all}},\mathbf{a}\in A^{|C|}}{\rm pred}(\mathbf{a},C)\cdot P(\mathbf{a}|C)\cdot\Pr_{\cal D}[C].$
where $\Pr_{\cal D}[C]$ denotes the probability assigned to $C$ by $\cal D$.
Remark 11.
One can assume that all contexts $C\in C^{{\rm all}}$ have the same size without loss of generality by adding additional observables, and having the predicate $\mathsf{pred}$ remain unchanged (i.e. the predicate does not depend on the values taken by the additional observables).
Let us consider the magic square as a contextuality game to clarify the notation.
Example 4 (Magic Square Contextuality Game).
The set of questions is
$Q:=\{(1,1),(1,2),(1,3),(2,1),(2,2),(2,3),(3,1),(3,2),(3,3)\}\,,$
$A:=\{+1,-1\}$ and $C^{{\rm all}}$ consists of all rows and columns of this matrix, i.e. subsets of the form $\{(r,1),(r,2),(r,3)\}$ for $r\in\{1,2,3\}$ and $\{(1,c),(2,c),(3,c)\}$ for $c\in\{1,2,3\}$. The distribution ${\cal D}$ uniformly selects an element $C\leftarrow C^{{\rm all}}$, i.e. it samples either the row or column uniformly. The predicate ${\rm pred}(\mathbf{a},C)$ is 00 unless the following holds:
$\prod_{q\in C}\mathbf{a}[q]=\begin{cases}-1&{\text{if }C=\{(1,3),(2,3),(3,3)\}}\\ 1&\text{else}.\end{cases}$
Call this game, $\mathsf{G}_{{\rm masq}}:=(Q,A,C^{{\rm all}},{\rm pred},{\cal D})$, the magic square game.
The game only becomes meaningful when the strategies that the players follow are somehow restricted. Thus, to consider non-contextual and quantum values associated with a contextuality game $\mathsf{G}$, we must first define the corresponding strategies.
For the following three definitions, let $\mathsf{G}=(Q,A,C^{{\rm all}})$ be a contextuality game.
Definition 12 (Classical/Non-Contextual Strategy for $\mathsf{G}$${\rm Strat}_{NC}(\mathsf{G})$).
We write a joint probability distribution over answers to all questions in $Q$ as
$P_{{\rm joint}}[\mathbf{a_{joint}}]\in[0,1]$
where $\mathbf{a_{joint}}$ is a vector of length $|Q|$, indexed by $Q$. A strategy $P$ is a non-contextual strategy if $P[\mathbf{a}|C]$ is derived from some joint probability $P_{{\rm joint}}[\mathbf{a_{joint}}]$, i.e.
$P(\mathbf{a}|C)=\sum_{\mathbf{a_{joint}}:\mathbf{a_{joint}}[C]=\mathbf{a}}P_{{\rm joint}}(\mathbf{a_{joint}})$
where by $\mathbf{a_{joint}}[C]=\mathbf{a}$ we mean that $\mathbf{a_{joint}}[\tilde{q}]=\mathbf{a}[\tilde{q}]$ for all $\tilde{q}\in C$ [^19] Note that $\mathbf{a_{joint}}$ is a vector of length $|Q|$ while $\mathbf{a}$ is a vector of length $|C|$. The set of all non-contextual strategies is denoted by ${\rm Strat}_{NC}(\mathsf{G})$.
We now define a quantum strategy for $\mathsf{G}$. It would be helpful to first define a “con-instance” corresponding to a contextuality game $\mathsf{G}$.
Definition 13 (Con-instance of $\mathsf{G}$).
A con-instance is simply a triple $({\cal H},\left|\psi\right\rangle,\mathbf{O})$, where ${\cal H}$ is a Hilbert space, $\left|\psi\right\rangle\in{\cal H}$ is a quantum state and $\mathbf{O}$ is a list of Hermitian operators (observables), indexed by the questions $Q$ (i.e. $\mathbf{O}[q]$ is an observable for each $q\in Q$) acting on ${\cal H}$. We call the triple $({\cal H},\left|\psi\right\rangle,\mathbf{O})$ a con-instance of $\mathsf{G}$ if $\mathbf{O}$ satisfies the following:
(i) ${\mathsf{Spectr}}[\mathbf{O}[q]]\subseteq A$ for all $q\in Q$, i.e. the values returned by the observables correspond to potential answers in the game.
(ii) $[\mathbf{O}[q],\mathbf{O}[q^{\prime}]]=0$ for all $q,q^{\prime}\in C$, for each $C\in C^{{\rm all}}$, i.e. operators corresponding to the same context are compatible.
Let $\mathbf{C}:=\{\{\mathbf{O}[q]\}_{q\in C}\}_{C\in C^{{\rm all}}}$ denote the set of contexts, except this time, we consider the operators corresponding to the questions. We refer to the list $({\cal H},\left|\psi\right\rangle,\mathbf{O},\mathbf{C})$ as an (explicit) con-instance of $\mathsf{G}$.
To be concrete, consider the Magic square as an example of a con-instance.
Example 5 (A con-instance for the magic square).
An explicit con-instance corresponding to LABEL:Exa:MagicSquare for the magic square is given by (,|$(\cal{H},\ket{\psi},\mathbf{O},\mathbf{C})$ where ${\cal{H}}:=\mathbb{C}^{2}\times\mathbb{C}^{2}$, |$\ket{\psi}\in\cal{H}$ is any fixed state, say |$\ket{\psi}=\ket{00}$, and $\mathbf{O}$ is specified by the following (indexed by the questions $Q$ in some canonical way):
$\begin{array}[]{ccc}\mathbb{I}\otimes X&Z\otimes\mathbb{I}&Z\otimes X\\ X\otimes\mathbb{I}&\mathbb{I}\otimes Z&X\otimes Z\\ X\otimes X&Z\otimes Z&XZ\otimes XZ\end{array}$ (17)
and ${\mathbf{C}}$ (set of “contexts”, i.e. commuting operators) is a set containing rows and columns of these operators, i.e. $\{\{\mathbb{I}\otimes X,X\otimes\mathbb{I},X\otimes X\},\ \{Z\otimes\mathbb{I},\mathbb{I}\otimes Z,Z\otimes Z\},\ \dots\{X\otimes X,Z\otimes Z,XZ\otimes XZ\}\}\in{\mathbf{C}}$.
The quantum strategy for a contextuality game $\mathsf{G}$ is defined as follows.
Definition 14 (Quantum Strategy for $\mathsf{G}$${\rm Strat}_{Qu}(\mathsf{G})$).
A strategy is a quantum strategy, i.e. $P\in{\rm Strat}_{Qu}(G)$, iff there is a con-instance $({\cal H},\left|\psi\right\rangle,\mathbf{O})$ of $\mathsf{G}$ satisfying
$P[\mathbf{a}|C]=\Pr\left[\text{Start with $\left|\psi\right\rangle$ measure $\mathbf{O}[C]$ \text{and obtain $\mathbf{a}$}}\right]\quad\forall\ \mathbf{a},C.$ (18)
Note that this is not the most general strategy, i.e. one could have different operators for each context and the operators in a given context may not even commute. Why then, do we restrict to the strategies defined above? This is because, as explained briefly in the introduction, the appeal of a contextuality test is that even by only measuring commuting observables at any given time, one can find scenarios where the idea of pre-determined values of observables becomes untenable.
While conceptually appealing, the main obstacle to testing contextuality using the contextuality game is that it is unclear how the provers’ strategies can be restricted to the ones above. As noted earlier, when one considers a Bell game as a contextuality game, spatial separation automatically restricts the provers’ strategies in this way.
Assuming that the provers follow only these restricted strategies, one can nevertheless define the classical/non-contextual and quantum value of the contextuality game.
Definition 15 (Non-Contextual and Quantum Value of a Contextuality Game $\mathsf{G}$).
Let $\mathsf{G}=(Q,A,C^{{\rm all}},{\rm pred},{\cal D})$ be a contextuality game (see LABEL:Def:ContextualityGame and recall the definition of ${\rm val}$). The non-contextual value of $\mathsf{G}$ is given by
${{\rm valNC}}:=\max_{P\in{\rm Strat}_{NC}(\mathsf{G})}{\rm val}(P)\quad\text{and similarly}\quad{{\rm valQu}}:=\max_{P\in{\rm Strat_{Qu}(\mathsf{G})}}{\rm val}(P).$
is the quantum value of $\mathsf{G}$.
Before moving further, we quickly note what these values are for the magic square.
Example 6 (Magic Square Contextuality Game (cont.)).
Let $\mathsf{G}_{{\rm masq}}$ be as in LABEL:Exa:masq and note that
${\rm valNC}=5/6,\quad\text{while}\quad{{\rm valQu}}=1$
using the Magic-Square con-instance as in LABEL:Exa:con-instance-The-Magic-Square.
The KCBS example (LABEL:Exa:KCBSintro) mentioned in the introduction, in this notation, can be expressed as follows.
Theorem 16 (KCBS).
There exists a contextuality game $\mathsf{G}$ with ${{\rm valNC}}=0.8<{{\rm valQu}}=\frac{2}{\sqrt{5}}\approx 0.8944$ where the quantum strategy can be realised using qutrits and five binary valued observables (i.e. the corresponding con-instance $({\cal H},\left|\psi\right\rangle,\mathbf{O})$ is such that $\dim{\cal H}=3$, $|Q|=|\mathbf{O}|=5$ and $A=\{\pm 1\}$).
7 OPad — Oblivious ${\mathbf{U}}$-Pad
Before describing our compiler, we introduce the oblivious pad primitive, and show how to construct it.
7.1 Definition
Let $\mathbf{U}:=\{U_{k}\}_{k\in K}$ be a set of unitaries acting on a Hilbert space $\cal{H}$, where $K$ is a finite set.
Definition 17.
An Oblivious $\mathbf{U}$-Pad (or an OPad) is a tuple of algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ as follows:
  • $\mathsf{Gen}$ is a PPT algorithm with the following syntax:
    • Input: $1^{\lambda}$ (a security parameter in unary).
    • Output: $(\mathsf{pk},\mathsf{sk})$.
  • $\mathsf{Enc}$ is A QPT algorithm with the following syntax:
    • Input: $\mathsf{pk},$ a state $\rho$ on $\cal H$.
    • Output: a state $\sigma$ (also on $\cal H$) and a string $s$.
  • $\mathsf{Dec}$ is a classical polynomial-time deterministic algorithm with the following syntax:
    • Input: $\mathsf{sk},s$.
    • Output: $k\in K$.
We require the following.
  • Correctness: There exists a negligible function $\mathsf{negl}$ such that, for all $\rho$ on $\mathcal{H}$, $\lambda\in\mathbb{N}$, the following holds with probability at least $1-\mathsf{negl}(\lambda)$ over sampling $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$, and $(\sigma,s)\leftarrow\mathsf{Enc}(\mathsf{pk},\rho)$:
    $\sigma\approx_{\mathsf{negl}(\lambda)}U_{k}\rho U_{k}^{\dagger}\,,$
    where $k=\mathsf{Dec}(\mathsf{sk},s)$.
  • Soundness: For any PPT prover $P$, there exists a negligible function $\mathsf{negl}$ (not necessarily equal to the previous one) such that, for all $\lambda\in\mathbb{N}$, $P$ wins in the following game with probability at most $1/2+\mathsf{negl}(\lambda)$.
Prover Challenger
$(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$
$\overset{\mathsf{pk}}{\longleftarrow}$
(The PPT prover can proceed arbitrarily)
[…]
$b\leftarrow\{0,1\}$ $k_{0}\leftarrow K$ $k_{1}=\mathsf{Dec}(\mathsf{sk},s)$
$\overset{k_{b}}{\longleftarrow}$
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
7.2 Instantiating the OPad — Oblivious Pauli Pad
LABEL:Alg:ObliviousPauliPad below shows how to instantiate an oblivious $\mathbf{U}$-pad in the random oracle model, where $\mathbf{U}=\{X^{x}Z^{z}\}_{x,z}$ where $x,z$ are strings. We refer to this as an oblivious Pauli pad. The instantiation leverages ideas from the proof of quantumness protocol in [BKVV20]. For convenience, we restate the informal description from LABEL:Subsec:Contributions. We describe the encryption procedure for a single qubit, and using a TCF pair $f_{0},f_{1}$ (rather than an NTCF as in LABEL:Alg:ObliviousPauliPad). To encrypt a multi-qubit state, one simply applies the following encryption procedure to each qubit.
We take $\mathsf{pk}=(f_{0},f_{1})$, and $\mathsf{sk}$ to be the corresponding trapdoor. Then, $\mathsf{Enc}_{\mathsf{pk}}$ is as follows:
  • (i)
[…]
  • (ii)
    Compute the random oracle “in the phase”, to obtain $(-1)^{H(x_{0})}\alpha\ket{0}\ket{x_{0}}+(-1)^{H(x_{1})}\beta\ket{1}\ket{x_{1}}$. Measure the second register in the Hadamard basis. This results in a string $d$, and the leftover qubit state
    |$\ket{\psi_{Z}}=Z^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\ket{\psi}\,.$
  • (iii)
    Repeat steps (i) and (ii) on |$\ket{\psi_{Z}}$, but in the Hadamard basis! This results in strings $y^{\prime}$ and $d^{\prime}$, as well as a leftover qubit state
    |$\ket{\psi_{XZ}}=X^{d^{\prime}\cdot(x_{0}^{\prime}\oplus x_{1}^{\prime})+H(x_{0}^{\prime})+H(x_{1}^{\prime})}Z^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\ket{\psi}\,,$
    where $x_{0}^{\prime}$ and $x_{1}^{\prime}$ are the pre-images of $y^{\prime}$.
Notice that the leftover qubit state |$\ket{\psi_{XZ}}$ is of the form $X^{x}Z^{z}\ket{\psi}$ where $x,z$ have the following two properties: (a) a verifier in possession of the TCF trapdoor can learn $z$ and $x$ given respectively $y,d$ and $y^{\prime},d^{\prime}$, and (b) no PPT prover can produce strings $y,d$ as well as predict the corresponding bit $z$ with non-negligible advantage (and similarly for $x$). Intuitively, this holds because a PPT prover that can predict $z$ with non-negligible advantage must be querying the random oracle at both $x_{0}$ and $x_{1}$ with non-negligible probability. By simulating the random oracle (by lazy sampling, for instance), one can thus extract a claw $x_{0},x_{1}$ with non-negligible probability, breaking the claw-free property.
Let
  • ${\cal F}$ be an NTCF family (see LABEL:Def:NTCF).
  • $\left|\psi\right\rangle\in{\cal H}$ be a state acting on $J$ qubits.
  • $H:\{0,1\}^{*}\to\{0,1\}$ denote the random oracle.
Define:
  • $\mathsf{Gen}(1^{\lambda})$:
    • Execute $(\mathsf{pk},\mathsf{sk})\leftarrow{\cal F}.\mathsf{Gen}(1^{\lambda})$ and output $(\mathsf{pk},\mathsf{sk})$.
  • $\mathsf{Enc}(\mathsf{pk},\text{$\rho$})$:
    • For each $j\in\{1\dots J\}$,
      • execute $\mathsf{qubitEnc}$ (as described in LABEL:Alg:Oblivious-Pauli-Pad-procedure) on qubit $j$ of $\rho$ using the public key $\mathsf{pk}$, and
      • use $(k_{j},s_{j})$ to denote $(k^{\prime},s^{\prime})$ as in LABEL:Eq:kPrime.
    • Let $k=(k_{1}\dots k_{J})$, and $s=(s_{1}\dots s_{J})$. Denote the resulting state by $\sigma_{k}$.
    • Return $\sigma_{k}$ and $s$.
  • $\mathsf{Dec}(\mathsf{sk},s)$:
    • For each $j\in\{1\dots J\}$
      • denote by $k_{j}$ the output of $\mathsf{qubitDec}(\mathsf{sk},s_{j})$ (as described in LABEL:Alg:qubitDec). [^20] Note that $\mathsf{qubitDec}$ does not output a qubit, but just a classical string! Nonetheless, we choose to call it $\mathsf{qubitDec}$, since the output are the quantum one-time pad keys associated to the $\mathsf{qubitEnc}$ procedure.
    • Return $k=(k_{1}\dots k_{j})$.
Algorithm 2 Oblivious Pauli Pad
$\mathsf{qubitEnc}(\mathsf{pk},\left|\psi\right\rangle)$ where $\left|\psi\right\rangle$ is a single qubit state (extended to mixed states by linearity)
  • Without loss of generality, suppose the state of the qubit is given by $\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle$.
  • Proceed as follows
    |$\ket{\psi}=\alpha\left|0\right\rangle+\beta\left|1\right\rangle$
    $\mapsto\sum_{x\in{\cal X},y\in{\cal Y}}\alpha\sqrt{({\cal F}.f^{\prime}_{\mathsf{pk},0}(x))(y)}\left|0\right\rangle\left|x\right\rangle\left|y\right\rangle+\beta\sqrt{({\cal F}.f^{\prime}_{\mathsf{pk},1}(x))(y)}\left|1\right\rangle\left|x\right\rangle\left|y\right\rangle$ Using $\mathsf{Samp}$
    $\approx_{\epsilon}\sum_{x\in{\cal X},y\in{\cal Y}}\alpha\sqrt{({\cal F}.f{}_{\mathsf{pk},0}(x))(y)}\left|0\right\rangle\left|x\right\rangle\left|y\right\rangle+\beta\sqrt{({\cal F}.f_{\mathsf{pk},1}(x))(y)}\left|1\right\rangle\left|x\right\rangle\left|y\right\rangle$ where $\text{where }\epsilon\text{ is negligible}$
[…]
    where $y$ is s.t.
    .${\cal F}.f_{\mathsf{pk},0}(x_{0})={\cal F}.f_{\mathsf{pk},1}(x_{1})=y$.
    $\mapsto\left((-1)^{H(x_{0})}\alpha\left|0\right\rangle\left|x_{0}\right\rangle+(-1)^{H(x_{1})}\beta\left|1\right\rangle\left|x_{1}\right\rangle\right)\ \left|y\right\rangle$ Applying the random oracle
    $H$ in the phase.
    $\mapsto\left((-1)^{d\cdot x_{0}+H(x_{0})}\alpha\left|0\right\rangle+(-1)^{d\cdot x_{1}+H(x_{1})}\beta\left|1\right\rangle\right)\ \left|d\right\rangle\left|y\right\rangle$ Measuring the second register
    in the Hadamard basis to get some $d$.
[…]
  • Relabel $(d,y,x_{0},x_{1})$ to $(d_{Z},y_{Z},x_{Z,0},x_{Z,1})$.
  • Denote by $\left|\phi\right\rangle=\alpha^{\prime}\left|+\right\rangle+\beta^{\prime}\left|-\right\rangle$ the state of the first qubit in LABEL:Eq:postMeasuredState. Continue as follows
    $\alpha^{\prime}\left|+\right\rangle+\beta^{\prime}\left|-\right\rangle$
    $\mapsto\sum_{x\in{\cal X},y\in{\cal Y}}\alpha\sqrt{({\cal F}.f^{\prime}_{\mathsf{pk},0}(x))(y)}\left|+\right\rangle\left|x\right\rangle\left|y\right\rangle+\beta\sqrt{({\cal F}.f^{\prime}_{\mathsf{pk},1}(x))(y)}\left|-\right\rangle\left|x\right\rangle\left|y\right\rangle$ By applying $\mathsf{Samp}$.
    $\vdots$
    proceed as above to obtain
    $\vdots$
    $=X^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\left|\phi\right\rangle\left|d\right\rangle\left|y\right\rangle$ Up to a global phase. (20)
  • Relabel $(d,y,x_{0},x_{1})$ to $(d_{X},y_{X},x_{X,0},x_{X,1})$.
  • Now note that the final state in LABEL:Eq:obliviousPauliPad_second can be written as
    $X^{\mathsf{phase}(d_{X},x_{X,0},x_{X,1})}Z^{\mathsf{phase}(d_{Z},x_{Z,0},x_{Z,1})}\ket{\psi}$ (21)
    where ${\rm\mathsf{phase}}(d,x_{0},x_{1}):=d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})$.
  • Define
    $k^{\prime}$ $:=({\mathsf{phase}(d_{X},x_{X,0},x_{X,1})},{\mathsf{phase}(d_{Y},x_{Y,0},x_{Y,1})}),\,\text{and}\,s^{\prime}:=(d_{X},y_{X},\ d_{Z},y_{Z}).$ (22)
  • Return the state in LABEL:Eq:correctnessOpad, and $s^{\prime}$.
Algorithm 3 $\mathsf{qubitEnc}$
$\mathsf{qubitDec}(\mathsf{sk},\underbrace{(d_{X},y_{X},\,d_{Z},y_{Z})}_{s^{\prime}})$
  • From $y_{X}$, compute $x_{X,0}=\mathsf{Inv}(\mathsf{sk},0,y_{X})$ and $x_{X,1}=\mathsf{Inv}(\mathsf{sk},1,y_{X})$.
  • Similarly, from $y_{Z}$, compute $x_{Z,0}=\mathsf{Inv}(\mathsf{sk},0,y_{Z}),x_{Z,1}=\mathsf{Inv}(\mathsf{sk},1,y_{Z})$.
  • Compute $k^{\prime}$ as in LABEL:Eq:kPrime, i.e.
    $k^{\prime}=({\mathsf{phase}(d_{X},x_{X,0},x_{X,1})},{\mathsf{phase}(d_{Y},x_{Y,0},x_{Y,1})})$ (23)
    where ${\rm\mathsf{phase}}(d,x_{0},x_{1}):=d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})$
  • Return $k^{\prime}$.
Algorithm 4 $\mathsf{qubitDec}$
Lemma 18 (Correctness).
Suppose ${\cal F}$ is an NTCF as in LABEL:Def:NTCF. Then, LABEL:Alg:ObliviousPauliPad satisfies the correctness requirement in LABEL:Def:Oblivious-U-pad.
Proof.
This is straightforward to verify given the properties of the NTCF.
Lemma 19 (Soundness).
Suppose ${\cal F}$ is an NTCF as in LABEL:Def:NTCF. Then, LABEL:Alg:ObliviousPauliPad satisfies the soundness requirement in LABEL:Def:Oblivious-U-pad.
Proof.
We first analyse the simpler game ${\cal G}$ (described below) and then observe that the reasoning carries over to the soundness of LABEL:Alg:ObliviousPauliPad.
Prover $({\cal P})$ Challenger $({\cal C})$
$(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$
$\overset{\mathsf{pk}}{\longleftarrow}$
[…]
$b\leftarrow\{0,1\}$
$h_{0}\leftarrow\{0,1\}$
$h_{1}=H(x_{0})\oplus H(x_{1})$ where $x_{0}={\cal F}.\mathsf{Inv}(\mathsf{sk},0,y)$ and $x_{1}={\cal F}.\mathsf{Inv}(\mathsf{sk},1,y)$.
$\overset{h_{b}}{\longleftarrow}$
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
Intuitively, it is clear that no PPT prover ${\cal P}$ wins with probability more than $1/2+{\rm negl}$ because if a PPT prover can distinguish $h_{0}$ from $h_{1}$, it must know $H$ at both $x_{0}$ and $x_{1}$ with non-negligible probability. By simulating the random oracle, one can then construct a PPT algorithm to extract preimages $x_{0},x_{1}$ of $y$. This violates the claw-free property of ${\cal F}$.
Formally, suppose that ${\cal P}$ succeeds with probability at most $1/2+\eta$ for some non-negligible function $\eta$. We show below that the following straightforward reduction extracts preimages $x_{0},x_{1}$ of $y$ with non-negligible probability:
Prover $({\cal P})$, queries to $H$ are simulated by ${\cal R}$ Reduction ${\cal R}$ that uses ${\cal P}$ Challenger .${\cal F}.{\cal C}$, for the claw-free property of ${\cal F}$
Answer queries to $H$ made by ${\cal P}$ by maintaining a database $D$. If the query to $H$ is at $x$, check if $D$ contains $x$, otherwise sample ${\rm val}\leftarrow\{0,1\}$, respond with ${\rm val}$ and append $(x,{\rm val})$ to $D$. $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$
$\overset{\mathsf{pk}}{\longleftarrow}$ $\overset{\mathsf{pk}}{\longleftarrow}$
[…]
$h_{0}\leftarrow\{0,1\}$
$\overset{h_{0}}{\longleftarrow}$
$\overset{b^{\prime}}{\longrightarrow}$
Ignore $b^{\prime}$.
Check if $D$ has two inputs $(x_{0},x_{1})$ with the same output $y$ under ${\cal F}$ (i.e. .${\cal F}.\mathsf{Chk}(\mathsf{pk},0,x_{0},y)=$${\cal F}.\mathsf{Chk}(\mathsf{pk},1,x_{1},y)=1$). Otherwise, set $x_{0}=x_{1}=0$.
$\xrightarrow{(x_{0},x_{1},y)}$
We lower bound the success probability of ${\cal R}$. Assume, without loss of generality, that ${\cal P}$ does not repeat queries. Let $E$ be the following event, during the execution of ${\cal P}$, interacting with ${\cal C}$: ${\cal P}$ makes a query at $x\in\{x_{0},x_{1}\}$ such that after this query, both $x_{0}$ and $x_{1}$ have been queried.
Note that $\neg E$ means that, by an information theoretic argument, ${\cal P}$ cannot distinguish the random variable $H(x_{0})\oplus H(x_{1})$ from a uniformly random bit.
Suppose ${\cal P}$ interacts with ${\cal C}$. Then,
[…]
NB: $\Pr[b=0|\neg E]=\Pr[b=1|\neg E]=\frac{1}{2}$ because $\Pr[b=0|\neg E]=\Pr[\neg E|b=0]\Pr[b=0]/\Pr[\neg E]$ and $\Pr[\neg E|b=0]=\Pr[\neg E|b=1]$ because until $E$ happens, $b=0$ and $b=1$ cannot make any difference in the execution of the protocol.
Continuing,
[…]
$\ \ \Pr[E]\Pr[{\cal P}\text{ wins}|E]$
$=\Pr[\neg E]\cdot\frac{1}{2}+\Pr[E]\cdot\Pr[{\cal P}\text{ wins}|E]$
$=\frac{1}{2}+\Pr[E]\cdot\left(\Pr[{\cal P}\text{ wins}|E]-\frac{1}{2}\right)$
[…]
To use this result, we first need to generalise to the case of multiple $y$s. This is also quite straightforward. Consider the following game ${\cal G}^{\prime}$:
Prover $({\cal P})$ Challenger $({\cal C})$
$(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$
$\overset{\mathsf{pk}}{\longleftarrow}$
$\xrightarrow{y_{1}\dots y_{2J}}$
$b\leftarrow\{0,1\}$
$h_{0}\leftarrow\{0,1\}^{2J}$ $h_{1}=h_{11},h_{21}\dots h_{2J,1}$ where $h_{j1}=H(x_{j0})\oplus H(x_{j1})$ where in turn $x_{j,0}={\cal F}.\mathsf{Inv}(\mathsf{sk},0,y_{j})$ and $x_{j,1}={\cal F}.\mathsf{Inv}(\mathsf{sk},1,y_{j})$.
$\overset{h_{b}}{\longleftarrow}$
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
Also consider the modified reduction ${\cal R}^{\prime}$ which is the same as the reduction ${\cal R}$, except that it receives $y_{1}\dots y_{2J}$ and at the last step, where it checks $D$ for any two inputs corresponding to any one of the $y_{1}\dots y_{2J}$. The success probability of ${\cal R}^{\prime}$ can be lower bounded as before, except that the event $E$ now becomes the following: ${\cal P}$ makes a query $x\in\{x_{j,0},x_{j,1}\}_{j\in[2J]}$ such that all for at least some $j$, both $x_{j,0}$ and $x_{j,1}$ have been queried.
The last step is to relate the game ${\cal G}^{\prime}$ with the soundness game in LABEL:Def:Oblivious-U-pad. The main difference between the two is that instead of $h_{0}$ and $h_{1}$, the soundness game uses $k_{0}$ and $k_{1}$ where, for $\mathsf{Dec}$ given by the oblivious pauli pad (i.e. LABEL:Alg:ObliviousPauliPad), $k_{1}$ has the form (for $j$ odd)
$k_{j1}=(d_{j}\cdot(x_{j0}\oplus x_{j1})+H(x_{j,0})+H(x_{j,1}),\ d_{j+1}\cdot(x_{j+1,0}\oplus x_{j+1,1})+H(x_{j+1,0})+H(x_{j+1,1})).$
The reasoning in computing the probability of $E$ goes through as above. The reduction ${\cal R}^{\prime}$ remains unchanged (it anyway was not computing $k_{1}$). Assuming that the claw-free property of ${\cal F}$ holds, we conclude that LABEL:Alg:ObliviousPauliPad satisfies the soundness condition. ∎
In the next section, we show how to use the oblivious pad, along with a $\mathsf{QFHE}$ scheme, to obtain a contextuality compiler. In order for this to be possible, the oblivious pad needs to be “compatible” with the $\mathsf{QFHE}$ scheme in the following sense.
Definition 20 ($\mathsf{OPad}$ compatible with $\mathsf{QFHE}$).
Suppose a $\mathsf{QFHE}$ scheme satisfies LABEL:Def:QFHEscheme with the form of encryption of $n$-qubit states specified by $\{U_{k}\}_{k\in K}$ as in LABEL:Eq:MBform. Then, an oblivious $\mathbf{U}$-Pad (or an $\mathsf{OPad}$) as in LABEL:Def:Oblivious-U-pad is compatible with the $\mathsf{QFHE}$ scheme if $\mathbf{U}=\{U_{k}\}_{k\in K}$.
8 Construction of the $(1,1)$ compiler
[…]
Our compiler takes as input the following:
  • A contextuality game $\mathsf{G}=:(Q,A,C^{{\rm all}},{\rm pred},{\cal D})$ with contexts of size $2$, i.e. satisfying $|C|=2$ for all $C\in C^{{\rm all}}$.
  • A con-instance $\mathbb{I}=:({\cal H},\left|\psi\right\rangle,\mathbf{O})$ for $\mathsf{G}$ (one that achieves the quantum value of $\mathsf{G}$; used to describe the honest prover)
  • A $\mathsf{QFHE}$ scheme as in LABEL:Def:QFHEscheme and a security parameter $\lambda$.
    • Let $\mathbf{U}=\{U_{k}\}_{k\in K}$ be the group (up to global phases) of unitaries acting on ${\cal H}$, as in LABEL:Eq:MBform.
    • Recall we use $\hat{k}$ to denote a classical encryption of $k$ under the secret key of $\mathsf{QFHE}$ as in LABEL:Subsec:FHE-Scheme.
  • An Oblivious $\mathbf{U}$-Pad scheme, $\mathsf{OPad}$, as in LABEL:Def:Oblivious-U-pad.
The compiler produces the following compiled game $\mathsf{G}^{\prime}$ between a verifier and prover (summarised in LABEL:Alg:The-1-1-compiler). We describe $\mathsf{G}^{\prime}$, along with the actions of an honest prover that achieves the completeness guarantee.
  1. 1.
    The verifier proceeds as follows:
    1. (a)
      Sample a secret key $\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda}),$ and a context $C\leftarrow{\cal D}$ according to the distribution specified by the game ${\cal D}$, picks a question $q^{\prime}$ from this context at random $q^{\prime}\leftarrow C$. Evaluate $c_{q^{\prime}}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(q^{\prime})$.
    2. (b)
      Sample a public key/secret key pair $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$.
    Send $(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})$ to the prover.
  2. 2.
    The honest prover prepares, under the $\mathsf{QFHE}$ encryption, the state $\left|\psi\right\rangle$ and subsequently measure $O_{q^{\prime}}$ to obtain an answer $a^{\prime}$. It then applies an oblivious $\mathbf{U}$-pad and returns the classical responses.
    1. (a)
      Since the operations happen under the $\mathsf{QFHE}$ encryption, the prover ends up with a $\mathsf{QFHE}$ encryption of $a^{\prime}$ which we denote by
      $c_{a^{\prime}}.$
      Further, it holds a $\mathsf{QFHE}$ encryption of the post-measurement state which (by the assumption on the form of the $\mathsf{QFHE}$ encryption) looks like
      $(U_{k^{\prime\prime}}\left|\psi_{q^{\prime},a^{\prime}}\right\rangle,\hat{k}^{\prime\prime}).$ (24)
      Here $\left|\psi_{q^{\prime},a^{\prime}}\right\rangle$ is the post-measurement state when $\left|\psi\right\rangle$ is measured using $O_{q^{\prime}}$ and the outcome $a^{\prime}$ is obtained.
    2. (b)
      The honest prover applies an oblivious $\mathbf{U}$-pad (see LABEL:Alg:ObliviousPauliPad) to obtain
      $(U_{k^{\prime}}U_{k^{\prime\prime}}\left|\psi_{q^{\prime},a^{\prime}}\right\rangle,s^{\prime})\leftarrow\mathsf{OPad}.\mathsf{Enc}(\mathsf{OPad}.\mathsf{pk},U_{k^{\prime\prime}}\left|\psi_{q^{\prime},a^{\prime}}\right\rangle).$
    3. (c)
      The prover returns $(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})$.
  3. 3.
    The verifier proceeds as follows:
    1. (a)
      Compute $a^{\prime}:=\mathsf{QFHE}.\mathsf{Dec}_{\mathsf{sk}}(c_{a^{\prime}})$, $k^{\prime\prime}$ from $\hat{k}^{\prime\prime}$ (the latter is by the form of QFHE) and $k^{\prime}=\mathsf{OPad}.\mathsf{Dec}(\mathsf{OPad}.\mathsf{sk},s^{\prime})$.
    2. (b)
      Find $k$ such that212121up to a global phase $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$ (such a $k$ always exists because $\mathbf{U}$ form a group).
    3. (c)
      Samples a new question $q\leftarrow C$.
    Send $(q,k)$ in the clear to the prover.
  4. 4.
    The honest prover measures observable $O_{q}$ conjugated by $U_{k}$, i.e. $U_{k}O_{q}U_{k}^{\dagger}$ and returns the corresponding answer $a$.
  5. 5.
    There are two cases, either the questions are the same or they are different (since $|C|=2$). The verifier proceeds as follows:
    1. (a)
      If $q=q^{\prime}$, accept if $a=a^{\prime}$;
    2. (b)
      If $q\neq q^{\prime}$, accept if ${\rm pred}((a,a^{\prime}),C)=1$.
We say that a prover wins $\mathsf{G}^{\prime}$ if the verifier accepts.
Honest Prover (${\cal A}$) Challenger (${\cal C}$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $C\leftarrow{\cal D}$ $q^{\prime}\leftarrow C$ $c_{q^{\prime}}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$
Under the QFHE encryption, measures $O_{q^{\prime}}$ and obtains an encrypted answer $c_{a^{\prime}}$ and post-measurement state $(U_{k^{\prime\prime}}\left|\psi_{q^{\prime}a^{\prime}}\right\rangle,\hat{k}^{\prime\prime})$. Applies an oblivious $\mathbf{U}$-pad to this state to obtain  $\left(U_{k^{\prime}}U_{k^{\prime\prime}}\left|\psi_{q^{\prime}a^{\prime}}\right\rangle,s^{\prime}\right)$$\leftarrow$ $\mathsf{OPad}.\mathsf{Enc}(\mathsf{OPad}.\mathsf{pk},U_{k^{\prime\prime}}\left|\psi_{q^{\prime}a^{\prime}}\right\rangle).$
$\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
Using the secret keys $\mathsf{sk},\mathsf{OPad}.\mathsf{sk}$, finds the $k$ such that $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$, samples $q\leftarrow C$
$\xleftarrow{(q,k)}$
Measures $U_{k}O_{q}U_{k}^{\dagger}$ and obtains $a$
[…]
Computes $a^{\prime}=\mathsf{Dec}_{\mathsf{sk}}(c_{a^{\prime}})$. If $q=q^{\prime}$, accept if $a=a^{\prime}$ If $q\neq q^{\prime}$, accept if ${\rm pred}((a,a^{\prime}),C)=1$.
Algorithm 5 Game $\mathsf{G}^{\prime}$ produced by the $(1,1)$-compiler for any contextuality game $\mathsf{G}$ with contexts of size two.
8.1 Compiler Guarantees
The compiler satisfies the following.
Theorem 21 (Guarantees of the $(1,1)$ compiled contextuality game $\mathsf{G}^{\prime}$).
Let $\mathsf{G}$ be a contextuality game with ${\rm valNC}<1$ and $|C|=2$ for all contexts $C\in C^{{\rm all}}$. Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:The-1-1-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
  • (Completeness) There is a negligible function $\mathsf{negl}$, such that, for all $\lambda\in\mathbb{N}$, the honest QPT prover from LABEL:Alg:The-1-1-compiler wins $\mathsf{G}^{\prime}_{\lambda}$ with probability at least
    $\frac{1}{2}\left(1+{{\rm valQu}}\right)-\mathsf{negl}(\lambda).$
  • (Soundness) For every PPT adversary $\cal A$, there is a negligible function $\mathsf{negl}^{\prime}$ such that, for all $\lambda\in\mathbb{N}$, the probability that $\cal A$ wins $\mathsf{G}^{\prime}_{\lambda}$ is at most
    $\frac{1}{2}\left(1+{\rm valNC}\right)+\mathsf{negl}^{\prime}(\lambda)\,,$
    assuming $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad).
Completeness is straightforward to verify. We prove soundness in Section 9.
9 Soundness Analysis
The security notion we considered in the definition of our $\mathsf{QFHE}$ scheme (see LABEL:Def:QFHEscheme) says that encryptions of any two distinct messages is indistinguishable from one another. The corresponding security game is often referred to as $2$-IND. What if one considers $\ell$-IND, i.e. $\ell$-many different encryptions? We prove that $2$-IND implies a version of $\ell$-IND which we call $\cal D$-IND’. $\cal D$-IND’ is more general since it allows the $\ell$ messages to be sampled from an arbitrary distribution $\cal D$ (instead of being limited to uniform). $\cal D$-IND’ is also more restricted in that (1) the $\ell$ messages are fixed and known to the challenger in advance and (2) we only consider classical algorithms. [^22] It appears that proving the general version in the quantum case may not be completely straightforward because of the inability to apply rewinding directly.
But why are we suddenly talking about $\ell$-IND and $\cal D$-IND’? It turns out that the proof technique used to show $2$-IND implies $\cal D$-IND’ can be applied, albeit it takes some care, to prove the soundness of the compiler.
Thus, as a warmup, LABEL:Subsec:warmup shows that $2$-IND implies $\cal D$-IND’. Then LABEL:Subsec:The-Reduction shows how to apply these ideas to show that the $2$-IND of $\mathsf{QFHE}$ encryptions implies soundness of the compiled game $\mathsf{G}^{\prime}$ (using LABEL:Alg:The-1-1-compiler) as stated in LABEL:Thm:CompiledGameIsSecure (assuming that the $\mathsf{OPad}$ is secure). We emphasise that we do not use the result in LABEL:Subsec:warmup directly—we only use the proof technique.
9.1 Warm up — $2$-IND implies ${\cal D}$-IND’
This subsection first recalls the $2$-IND game (which is essentially a restatement of LABEL:Eq:distinguisher-1). Then it defines the ${\cal D}$-IND’ game (the prime emphasises that the number of messages is constant and known apriori [^23] Proving the implication for the general case may be non-trivial in the quantum setting.). It ends by showing that $2$-IND implies ${\cal D}$-IND’.
Claim 22 ($2$-IND Game for QFHE).
Let $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ correspond to a $\mathsf{QFHE}$ encryption scheme. The scheme satisfies LABEL:Eq:distinguisher-1 if and only if the following holds:
In the game below, for every PPT distinguisher, there is a negligible function ${\rm negl}$ such that the challenger accepts with probability at most $\frac{1}{2}+{\mathsf{negl}}(\lambda).$
Distinguisher Challenger
$\xrightarrow{(x_{0},x_{1})}$
$\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ $b\leftarrow\{0,1\}$ $c_{b}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(x_{b})$
$\overset{c_{b}}{\longleftarrow}$
Objective: guess $b$
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$.
Definition 23 (${\cal D}$-IND’ Game (where ${\cal D}$ is a distribution over a fixed message set) for QFHE).
Let
  • $\lambda$ be a security parameter, $\ell=O(1)$ be a fixed constant (relative to $\lambda$),
  • $M=\{m_{1}\dots m_{\ell}\}$ be a set of distinct messages where $m_{1},\dots m_{\ell}\in\{0,1\}^{O(1)}$ are of constant size,
  • ${\cal D}$ be a probability distribution over $M$ and let $q_{{\rm guess}}=\max_{i}q_{i}$ where $q_{i}$ is the probability assigned to $m_{i}$ by by ${\cal D}$.
The ${\cal D}$-IND’ Security Game is a two-party game, for $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ as specified by $\mathsf{QFHE}$, is as follows:
Distinguisher Challenger
$\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ $m\leftarrow{\cal D}$ $c_{m}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(m)$
$\overset{c_{m}}{\longleftarrow}$
Objective: guess $m$
$\overset{m^{\prime}}{\longrightarrow}$
Accept if $m^{\prime}=m$.
Proposition 24.
A $\mathsf{QFHE}$ scheme that satisfies LABEL:Eq:distinguisher-1, implies that every PPT distinguisher for the ${\cal D}$-IND’ game above, wins with probability at most $q_{{\rm guess}}+{\mathsf{negl}}(\lambda)$.
Proof.
We can assume, without loss of generality, that the messages are given by $m_{i}=i$. This is because they can be uniquely indexed, and our arguments go through unchanged (as one can check).
Goal: we want to show that if there is a PPT distinguisher [^24] Note that ${\cal A}_{\ell}$ and $C_{\ell}$ can depend on ${\cal D}$; the use of the subscript $\ell$ is just notational convenience and not meant to convey all the dependencies. ${\cal A}_{\ell}$ that wins against $C_{\ell}$ in the ${\cal D}$-IND’ game above with probability $q_{{\rm guess}}+\eta^{\prime}$ for some non-negligible function $\eta^{\prime}$, then one can construct a PPT distinguisher ${\cal A}_{2}$ that wins against $C_{2}$ in the $2$-IND game (see LABEL:Claim:2-IND_QFHE) with probability at least $\frac{1}{2}+\eta^{\prime\prime}$ for some other non-negligible function $\eta^{\prime\prime}$.
To this end, denote by $p_{ij}$ the probability that ${\cal A}_{\ell}$ outputs $j$ when given the encryption of $i$ as input, i.e.
$p_{ij}:=\Pr\left[j\leftarrow{\cal A}_{\ell}(c_{i}):\begin{array}[]{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ c_{i}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(i)\end{array}\right].$
Since the message space is of constant size, ${\cal A}_{2}$ can compute $p_{ij}$ to inverse polynomial errors. For now, we assume ${\cal A}_{2}$ knows $p_{ij}$ exactly and handle the precision issue at the end. To proceed, consider the following observations:
  1. 1.
    $\Pr[\text{accept}\leftarrow\left\langle{\cal A}_{\ell},C_{\ell}\right\rangle]=\sum_{i\in\{1\dots\ell\}}q_{i}p_{ii}$ where recall that $q_{i}$ is the probability assigned to $i$ by the distribution ${\cal D}$.
  2. 2.
    There exist $k^{*}\neq i^{*}$ such that $\sum_{j}\left|p_{i^{*}j}-p_{k^{*}j}\right|\geq\eta$ for some non-negligible function $\eta$.
The first observation follows directly from the definition of ${\cal D}$-IND’ and $p_{ij}$. The second follows from the assumption that ${\cal A}_{\ell}$ wins with probability at least $q_{{\rm guess}}+\eta^{\prime}$ for some non-negligible function $\eta^{\prime}$. To see this, proceed by contradiction: Suppose for all $i^{*}\neq k^{*}$, it is the case that $\sum_{j}\left|p_{i^{*}j}-p_{k^{*}j}\right|\leq{\rm negl}$ for some negligible function, then one could write
$\Pr[\text{accept}\leftarrow\left\langle{\cal A}_{\ell},C_{\ell}\right\rangle]$ $=\sum_{i\in\{1\dots\ell\}}q_{i}p_{ii}$
$=\sum_{i\in\{1\dots\ell\}}q_{i}(p_{i^{*}i}+{\rm negl})$ by the assumption above
[…]
$\leq q_{{\rm guess}}+{\rm negl}.$
But this cannot happen because we assumed that ${\cal A}_{\ell}$ wins with probability non-negligibly greater than $q_{{\rm guess}}$. We therefore conclude that observation 2 must hold. Since the message space is constant, a PPT ${\cal A}_{2}$ can determine the following:
  • The indices $i^{*}\neq k^{*}$
  • The disjoint sets $J_{0},J_{1}\subseteq\{1\dots\ell\}$ such that
    • for $j\in J_{0}\subseteq\{1\dots\ell\}$, $p_{i^{*}j}\geq p_{k^{*}j}$, and
    • for $j\in J_{1}\subseteq\{1\dots\ell\}$, $p_{i^{*}j}<p_{k^{*}j}$.
Once these two indices, and the sets $J_{0},J_{1}$ are known, ${\cal A}_{2}$’s remaining actions are as follows:
${\cal A}_{3}$ (adversary for ${\cal D}$-IND’) ${\cal A}_{2}$ (the reduction) ${\cal C}_{2}$ (challenger for $2$-IND)
Let $k^{*}\neq i^{*}\in\{1\dots\ell\}$ be as above.
$\xrightarrow{(k^{*},i^{*})=:(x_{0},x_{1})}$
$\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ $b\leftarrow\{0,1\}$ $c_{b}\leftarrow\mathsf{Enc}_{\mathsf{sk}}(x_{b})$
$\overset{c_{b}}{\longleftarrow}$ $\overset{c_{b}}{\longleftarrow}$
$\xrightarrow{j_{{\rm guess}}}$
$b^{\prime}=\begin{cases}0&\text{if }j_{{\rm guess}}=J_{0}\\ 1&\text{if }j_{{\rm guess}}=J_{1}\end{cases}$ $\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$.
Now,
$\Pr[\text{accept}\leftarrow\left\langle{\cal A}_{2},{\cal C}_{2}\right\rangle]=$ $\frac{1}{2}\cdot\sum_{j\in J_{0}}\Pr[{\cal A}_{\ell}\text{ outputs }j|i^{*}\text{ was encrypted}]+$
$\frac{1}{2}\cdot\sum_{j\in J_{1}}\Pr[{\cal A}_{\ell}\text{ outputs }j|k^{*}\text{ was encrypted}]$
$=$ $\frac{1}{2}+\frac{1}{2}\text{$\sum_{j\in J_{0}}\left(p_{i^{*}j}-p_{k^{*}j}\right)$}$ By def of $\text{By def of }p_{ij}\text{ and normalisation}$
$=$ $\frac{1}{2}+\frac{1}{4}\sum_{j}|p_{i^{*}j}-p_{k^{*}j}|$ $\left\|a-b\right\|_{1}=2\sum_{i:a_{i}>b_{i}}a_{i}-b_{i}$ for $a,b$ distrib. (25)
[…]
Since the $\mathsf{QFHE}$ scheme satisfies LABEL:Eq:distinguisher-1, this is a contradiction (via LABEL:Claim:2-IND_QFHE) and thus the claim follows—up to the precision issue which we now address.
Handling the precision issue.
Denote by ${\cal A}_{2}$ the algorithm above where $p_{ij}$ are known exactly. Suppose that
$\Pr[\text{accept}\leftarrow\langle{\cal A}_{\ell},C_{\ell}\rangle]\geq q_{{\rm guess}}+\epsilon$ (26)
where $\epsilon$ is a non-negligible function. For concreteness, suppose [^25] For every non-negligible function $\epsilon$, there is an infinite subset $\Lambda$ of its domain where $\epsilon(\lambda)\geq 1/\lambda^{c}$. One can restrict the entire argument to this domain and still reach the same conclusion. $\epsilon(\lambda)=1/\lambda^{c}$ for some constant $c>0$.
Proposition 25.
Given that LABEL:Eq:prAccept holds, consider an algorithm $\hat{{\cal A}}_{2}$ that uses an estimate for $\hat{p}_{ij}$ up to precision $O(\epsilon^{3})$ (i.e. $|\hat{p}_{ij}-p_{ij}|\leq O(\epsilon^{3})$). Then,
$\Pr[{\rm accept}\leftarrow\langle\hat{{\cal A}}_{2},C_{2}\rangle]$ $\geq\Pr[{\rm accept}\leftarrow\langle{\cal A}_{2},C_{2}\rangle]-O(\epsilon^{3}).$ (27)
To prove LABEL:Prop:finitePrecisionD-IND_prime, we first show (see LABEL:Claim:epsilonsquare below) that $\sum_{j}|p_{i^{*}j}-p_{k^{*}j}|$ is at least $\Omega(\epsilon^{2})$ given that LABEL:Eq:prAccept holds. Using this and LABEL:Eq:p_istarj-p_kstarj, it follows that $\Pr[{\rm accept}\leftarrow\langle{\cal A}_{2},C_{2}\rangle]$ is at least one half plus $\Omega(\epsilon^{2})$ which means $\hat{{\cal A}}_{2}$—the finite-precision variant of ${\cal A}_{2}$—succeeds with one half plus non-negligible probability in the $2$-IND game of the $\mathsf{QFHE}$ scheme (see LABEL:Claim:2-IND_QFHE).
Claim.
If LABEL:Eq:prAccept holds then there exist $i^{*}\neq k^{*}$ such that $\left\|p_{i^{*}}-p_{j^{*}}\right\|_{1}:=\sum_{j}\left|p_{i^{*}j}-p_{k^{*}j}\right|\geq\Omega(\epsilon^{2})$.
Proof of LABEL:Claim:epsilonsquare.
We start with an elementary fact: Let $f,g:\Lambda\to\mathbb{R}$, where $\Lambda\subset\mathbb{N}$ is an infinite subset of $\mathbb{N}$. Let $P$ be the proposition that $f\leq O(g)$ on the set $\Lambda$. Then $\neg P$ implies that $f\geq\Omega(g)$ on an infinite set $\Lambda^{\prime}\subseteq\Lambda$.
Using this fact, deduce that the negation of $\sum_{j}|p_{i^{*}j}-p_{k^{*}j}|\geq\Omega(\epsilon^{2})$ implies there is some infinite set $\Lambda\subseteq\mathbb{N}$ over which $\sum_{j}|p_{i^{*}j}-p_{k^{*}j}|\leq O(\epsilon^{2})$. We prove the claim by contradiction. Consider to the contrary that for all $i^{*}\neq k^{*}$, $\sum_{j}|p_{i^{*}j}-p_{k^{*}j}|<O(\epsilon^{2})$. Then, it holds that
$\Pr[\text{accept}\leftarrow\langle{\cal A}_{\ell},C_{\ell}\rangle]$ $=\sum_{i\in\{1\dots\ell\}}q_{i}p_{ii}$
$\leq\sum_{i\in\{1\dots\ell\}}q_{i}(p_{i^{*}i}+O(\epsilon^{2}))$
[…]
$=q_{{\rm guess}}+O(\epsilon^{2})$
but this violates LABEL:Eq:prAccept. ∎
We can now prove LABEL:Prop:finitePrecisionD-IND_prime.
Proof of LABEL:Prop:finitePrecisionD-IND_prime.
Since $\hat{{\cal A}}_{2}$ estimates $p_{ij}$ to precision $O(\epsilon^{3})$, it follows that it will find $i^{*}\neq k^{*}$ and a $j$ such that $|p_{i^{*}j}-p_{k^{*}j}|\geq\Omega(\epsilon^{2})$; recall that by the definition, $\ell=O(1)$ so $\left\|p_{i^{*}}-p_{j^{*}}\right\|_{1}\geq\Omega(\epsilon^{2})$ implies there is some $j$ for which $|p_{i^{*}j}-p_{k^{*}j}|\geq\Omega(\epsilon^{2})$.
Further, the only $j$s for which $\hat{{\cal A}}_{2}$ makes an error in deciding whether to have $j\in J_{0},$ or $j\in J_{1}$, are those for which $|p_{i^{*}j}-p_{k^{*}j}|\leq O(\epsilon^{3})$. (Note that it cannot be that for all $j$s, $|p_{i^{*}j}-p_{k^{*}j}|\leq O(\epsilon^{3})$ because then $\left\|p_{i^{*}}-p_{k^{*}}\right\|_{1}$ cannot be $\geq\Omega(\epsilon^{2})$.) Thus, the error in computing $\Pr[{\rm accept}\leftarrow\langle\hat{{\cal A}}_{2},C_{\ell}\rangle]$ using $\Pr[{\rm accept}\leftarrow\langle{\cal A}_{2},C_{\ell}\rangle]=\frac{1}{2}+\frac{1}{2}\sum_{j\in J_{0}}\left(p_{i^{*}j}-p_{k^{*}j}\right)$ is at most $O(\epsilon^{3})$ (again, using $\ell=O(1)$). This yields LABEL:Eq:epsiloncubeerrorinA_2hat as asserted. ∎
This completes the proof. ∎
9.2 The Reduction
We are going to essentially adapt the proof of LABEL:Prop:DINDprime to our setting. Start by considering the honest prover ${\cal A}$ in LABEL:Alg:The-1-1-compiler. We reduce ${\cal A}$ to a distinguisher ${\cal A}_{2}$ for the $2$-IND security game of $\mathsf{QFHE}$.
  • Notation: Let $(Q,A)$ denote the set of questions and answers in the contextuality game that was compiled.
  • Phase 1: Learning “$p_{ij}$”.
    • Consider the following procedure ${\cal A}_{2}.\mathsf{TruthTable}$ that
      takes $q^{\prime}\in Q$ as input and
      produces a truth table $\tau:Q\to A.$
      1. 1.
        ${\cal A}_{2}$ simulates the challenger ${\cal C}$ in LABEL:Alg:The-1-1-compiler as needed (see LABEL:Alg:ATruthTable), except that it takes $q^{\prime}$ as input (instead of uniformly sampling it in the beginning) and uses $k\leftarrow K$ (in the third step). It uses this simulation to generate the first message $(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})$ and feeds it to ${\cal A}$.
      2. 2.
        ${\cal A}_{2}$ receives $(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})$ from ${\cal A}$. For each $q\in Q$, ${\cal A}_{2}$ sends $(q,k)$ to ${\cal A}$ (where recall $k$ is sampled uniformly at random from $K$) and receives an answer $a$. Denote by $\tau$ the truth table, i.e. the list of answers indexed by the questions.
    • ${\cal A}_{2}$ repeats the procedure ${\cal A}_{2}\cdot\mathsf{TruthTable}(q^{\prime})$ above for each $q^{\prime}\in Q$ to estimate the probability $p_{q^{\prime}\tau}$ of the procedure outputting $\tau$ on input $q^{\prime}$ (the randomness is also over the encryption procedure etc).
      Here $p_{q^{\prime}\tau}$ is analogous to $p_{ij}$.
    • It finds questions $q_{*0}\neq q_{*1}$ such that
      $\left\|p_{q_{*0}}-p_{q_{*1}}\right\|:=\sum_{\tau}|p_{q_{*0}\tau}-p_{q_{*1}\tau}|\geq\eta$ (28)
      for some non-negligible function $\eta$.
      (We defer the proof that questions (or indices) $q_{*0}\neq q_{*1}$ exist if the $\mathsf{OPad}$ is secure and ${\cal A}$ wins the compiled contextuality game $\mathsf{G}^{\prime}$, with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$. )
    • It defines the disjoint sets $T_{0}$ and $T_{1}$ as follows: $\tau\in T_{0}$ if $p_{q_{*0}\tau}\geq p_{q_{*1}\tau}$ and $\tau\in T_{1}$ if $p_{q_{*1}\tau}>p_{q_{*0}\tau}$.
    ${\cal A}$ ${\cal A}_{2}.\mathsf{TruthTable}(q^{\prime})$
    $\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c_{q^{\prime}}\leftarrow\mathsf{QFHE}.\mathsf{Enc}(q^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
    $\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$
    $\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
    Samples $k\leftarrow K$
[…]
    Outputs $\tau$
    Algorithm 6 The procedure ${\cal A}_{2}.\mathsf{TruthTable}$ takes as input a question $q^{\prime}$ and produces a truth table $\tau$ corresponding to it. Note that this is a randomised procedure (depends on the $\mathsf{QFHE}$ encryption procedure) so for the same $q^{\prime}$ the procedure may output different $\tau$s. The goal is to learn the probabilities of different $\tau$s appearing for each question $q^{\prime}$.
  • Phase 2: Interaction with $C_{2}$ of the $2$-IND game.
    • ${\cal A}_{2}$ sends $q_{*0},q_{*1}$ to $C_{2}$ and $C_{2}$ returns $c_{b}$, the QFHE encryption of $q_{*b}$ (where $C_{2}$ picks $b\leftarrow\{0,1\}$).
    • ${\cal A}_{2}$ simulates the $\mathsf{OPad}$ itself and forwards $c_{b}$ to ${\cal A}$, learns the truth table $\tau$ corresponding to $c_{b}$ and outputs $b^{\prime}\in\{0,1\}$ such that $\tau\in T_{b^{\prime}}$.
    ${\cal A}$ ${\cal A}_{2}$
    Phase 1
[…]
    Use $p_{q^{\prime}\tau}$ to learn the questions $q_{*0},q_{*1}$ and the disjoint sets $T_{0},T_{1}$ of truth tables.
    ${\cal A}$ Phase 2 $C_{2}$
    $\xrightarrow{(q_{*0},q_{*1})}$
    $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$ $\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $b\leftarrow\{0,1\}$ $c_{b}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q_{*b})$
    $\xleftarrow{(c_{b},\mathsf{OPad}.\mathsf{pk})}$ $\overset{c_{b}}{\longleftarrow}$
    $\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
    Samples $k\leftarrow K$
[…]
    Set $b^{\prime}=0$ if $\tau\in T_{0}$, and $b^{\prime}=1$ if $\tau\in T_{1}$ $\overset{b^{\prime}}{\longrightarrow}$
    Accept if $b^{\prime}=b$
    Algorithm 7 The algorithm ${\cal A}_{2}$ uses the adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ to break the $2$-IND security game for the $\mathsf{QFHE}$ scheme.
    The intuition is that given an encryption of $q_{*0}$, on an average, the above procedure would output $b^{\prime}=0$ more often than $b^{\prime}=1$, essentially by definition of $T_{0}$ and $T_{1}$, and LABEL:Eq:non-negl-ptau.
Remark 26.
There are three subtleties that are introduced, aside from the rewinding needed to learn $\tau$, in analysing $\mathsf{G}^{\prime}$ as opposed to the ${\cal D}$-IND’ game. We glossed over these above.
  1. 1.
    Random $k$: When interacting with the prover ${\cal A}$, ${\cal A}_{2}$ above is feeding in a uniformly random $k$ instead of the correct $k$ which depends on $\hat{k}^{\prime\prime}$ and $s^{\prime}$ (see LABEL:Alg:The-1-1-compiler where ${\cal A}$ interacts with ${\cal C}$ and compare it to the interaction of ${\cal A}$ with ${\cal A}_{2}$).
    • However, the guarantees about ${\cal A}$ are for the correct $k$.
    • We need to formally show that the security of $\mathsf{OPad}$ allows us to work with random $k$s directly.
    • This is straightforward enough but we need to use this a few times; we’ll see.
  2. 2.
    ${\cal A}$ is consistent: The other subtlety has to do with the proof that ${\cal A}$ winning with probability non-negligibly more than $\frac{1}{2}(1+{\rm valNC})$ implies $\left\|p_{q_{*0}}-p_{q_{*1}}\right\|_{1}$ is non-negligible. In the proof, we use the assumption that ${\cal A}$ is consistent, i.e. if it answers with an encryption of $a^{\prime}$ upon being asked an encryption of $q^{\prime}$ in the first interaction, it will respond consistently if the same question is asked in the clear, i.e. it answers $a^{\prime}$ upon being asked the same question $q^{\prime}$ in the next round. We will show that given an ${\cal A}$ that is non-consistent, one can still bound it’s success probability by treating it as though it is consistent.
  3. 3.
    Precision of $p_{q^{\prime}\tau}$: We also completely skipped the precision issue, i.e. we assumed ${\cal A}_{2}$ can learn $p_{q^{\prime}\tau}$ exactly. In practice, this is not possible, of course. However, $p_{q^{\prime}\tau}$ can be learnt to enough precision to make the procedure work, just as we did in the proof of LABEL:Prop:DINDprime.
9.3 Proof Strategy
Let ${\cal A}$ be any PPT algorithm interacting with ${\cal C}$ in the compiled contextuality game $\mathsf{G}^{\prime}$.
The following allows us to assume that we can give a uniformly random $k$ as input to ${\cal A}$ without changing the output distribution of the interaction in LABEL:Alg:random-k-or-not (as alluded to in point 1 of LABEL:Rem:RandomK_ConsistentA_PrecisionOfP_itau).
Lemma 27 (Uniformly random $k$ is equivalent to the correct $k$).
Let $B_{0}$ (resp. $B_{1}$) be a PPT algorithm that takes $q^{\prime}\in Q$ as an input, interacts with ${\cal A}$ and outputs a bit, as described in LABEL:Alg:random-k-or-not. Then there is a negligible function $\mathsf{negl}$ such that $\left|\Pr[0\leftarrow\left\langle B_{0},{\cal A}\right\rangle]-\Pr[0\leftarrow\left\langle B_{1},{\cal A}\right\rangle]\right|\leq\mathsf{negl}$.
${\cal A}$ $B_{0}(q^{\prime})$ (resp. $B_{1}(q^{\prime})$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c_{q^{\prime}}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
$B_{0}$ computes $k^{\prime}:=\mathsf{OPad}.\mathsf{Dec}(\mathsf{OPad}.\mathsf{sk},s^{\prime})$ and uses the secret key $\mathsf{sk}$ to compute $k^{\prime\prime}$ and then finds the $k$ satisfying $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$. (resp. $B_{1}$ samples a uniform $k\leftarrow K$).
$\overset{(q,k)}{\circlearrowleft}$ Potentially rewinds ${\cal A}$ to this step and queries with $(q,k)$ for arbitrary $q\in Q$.
Runs an arbitrary procedure to compute a bit $b^{\prime}$.
Algorithm 8 Whether a PPT adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ is used with the correct $k$ or a uniformly random $k$ does not affect the outcome of this interaction more than negligibly.
The following allows us to treat ${\cal A}$ as though it is consistent (as anticipated in point 2 of LABEL:Rem:RandomK_ConsistentA_PrecisionOfP_itau).
Lemma 28 (Consistency only helps).
Let
  • ${\cal C}_{k\leftarrow K}$ be exactly the same as the challenger ${\cal C}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ except that it samples a uniformly random $k$ instead of computing it correctly, let
  • ${\cal A}$ be any PPT algorithm that is designed to play the complied contextuality game $\mathsf{G}^{\prime}$ and makes ${\cal C}_{k\leftarrow K}$ accept with probability $p$, i.e. $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle]=p$, and denote by
  • $p_{q^{\prime}\tau}$ the probability that on being asked $q^{\prime}$, the truth table ${\cal A}$ produces is $\tau$ (as defined in LABEL:Subsec:The-Reduction).
Then
$p\leq\sum_{C}\Pr(C)\cdot\frac{1}{2}\left(1+\frac{1}{|C|}\sum_{q^{\prime}\in C}\sum_{\tau}p_{q^{\prime}\tau}\mathsf{pred}(\tau[C],C)\right)$
where $\Pr(C)$ denotes the probability with which ${\cal C}$ samples the context $C$.
The two terms in the sum above, correspond to the consistency test ($q=q^{\prime}$) and the predicate test ($q^{\prime}\neq q$). Finally, the following allows us to neglect the precision issue (as detailed in point 3 of LABEL:Rem:RandomK_ConsistentA_PrecisionOfP_itau).
Lemma 29 (Precision is not an issue).
Suppose
  • ${\cal A}$ wins with probability $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq\frac{1}{2}(1+{\rm valNC})+\epsilon$ for some non-negligible function $\epsilon$
  • ${\cal A}_{2}$ is as in LABEL:Alg:A_2_reduction, i.e. it is a PPT algorithm except for the time it spends in learning $p_{q^{\prime}\tau}$ exactly
  • Denote by ${\cal A}_{2,\epsilon}$ a PPT algorithm that is the same as ${\cal A}_{2}$ except that it computes and uses an estimate $\hat{p}_{q^{\prime}\tau}$ satisfying $|\hat{p}_{q^{\prime}\tau}-p_{q^{\prime}\tau}|\leq O(\epsilon^{3})$, in place of $p_{q^{\prime}\tau}$.
Then, the PPT algorithm ${\cal A}_{2,\epsilon}$ wins with essentially the same probability as ${\cal A}_{2}$, i.e.
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2,\epsilon},C_{2}\right\rangle]\geq\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2},C_{2}\right\rangle]-O(\epsilon^{3}).$
We prove the following contrapositive version of the soundness guarantee in LABEL:Thm:CompiledGameIsSecure.
Theorem 30 (Soundness condition restated from LABEL:Thm:CompiledGameIsSecure).
Suppose
  • ${\cal A}$ is any PPT algorithm that wins with probability $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq\frac{1}{2}(1+{\rm valNC})+\epsilon$ for some non-negligible function $\epsilon$, and
  • the $\mathsf{OPad}$ used is secure, then
there is a PPT algorithm ${\cal A}_{2,\epsilon}$ that wins the $2$-IND security game of the $\mathsf{QFHE}$ scheme with probability
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2,\epsilon},C_{2}\right\rangle]\geq\frac{1}{2}+\mathsf{nonnegl},$
where $\mathsf{nonnegl}$ is a non-negligible function that depends on $\epsilon$.
In LABEL:Subsec:Step1, we prove LABEL:Thm:MainSoundness assuming LABEL:Lem:1guessK=00003DgoodK,_2EstimateAssumingConsistent,_3PrecisionNonIssue. In LABEL:Subsec:Step2, we prove the lemmas.
9.4 Proof assuming the lemmas (Step 1 of 2)
Proof.
This part is analogous to the proof of LABEL:Prop:DINDprime. First, recall the challenger ${\cal C}$ of the complied game $\mathsf{G}^{\prime}$ and let ${\cal C}_{k\leftarrow K}$ denote the the same challenger, except that it samples $k$ uniformly at random. From LABEL:Lem:1guessK=00003DgoodK, one can conclude that $|\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]-\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle]|\leq\mathsf{negl}$.
Recall the definition of $p_{q^{\prime}\tau}$ from the discussion in LABEL:Subsec:The-Reduction. Using LABEL:Lem:2EstimateAssumingConsistent, one can write
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]-\mathsf{negl}$ $\leq\sum_{C}\Pr(C)\cdot\frac{1}{2}\left(1+\frac{1}{|C|}\sum_{q^{\prime}\in C}\sum_{\tau}p_{q^{\prime}\tau}\mathsf{pred}(\tau[C],C)\right)$
$=\frac{1}{2}\left(1+\sum_{C}\Pr(C)\frac{1}{|C|}\sum_{q^{\prime}\in C}\sum_{\tau}p_{q^{\prime}\tau}\mathsf{pred}(\tau[C],C)\right)$ (29)
Observe also that there exist $q_{*0}\neq q_{*1}$ such that
$\left\|p_{q_{*0}}-p_{q_{*1}}\right\|_{1}:=\sum_{\tau}|p_{q_{*0}\tau}-p_{q_{*1}\tau}|\geq\eta$ (30)
for some non-negligible function $\eta$. This is a consequence of the assumption that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq\frac{1}{2}(1+{\rm valNC})+\epsilon$ (31)
for some non-negligible function $\epsilon$. To see this, proceed by contradiction: Suppose that for all $q_{*0}\neq q_{*1}$, it is the case that $\sum_{\tau}|p_{q_{*0}\tau}-p_{q_{*1}\tau}|\leq\mathsf{negl}$ for some negligible function, then one could write, using LABEL:Eq:pr_acc_A_C,
[…]
$=\frac{1}{2}\left(1+\sum_{\tau}p_{q_{*0}\tau}\sum_{C}\Pr(C)\mathsf{pred}(\tau[C],C)\right)+\mathsf{negl}^{\prime}$
$\leq\frac{1}{2}\left(1+{\rm valNC}\right)+\mathsf{negl}^{\prime}$ (32)
where in the first step, we used $p_{q_{*0}\tau}$ instead of $p_{q^{\prime}\tau}$ at the cost of a $\mathsf{negl}^{\prime}$ additive error, in the second step, we rearranged the sum, and in the final step, we observe that the expression is just a convex combination of values achieved using non-contextual strategies, and this is at most ${\rm valNC}$. However, LABEL:Eq:valNC_at_most contradicts LABEL:Eq:A_wins_against_C_w_non-negl that says, by assumption, ${\cal A}$ wins with probability non-negligibly more than $\frac{1}{2}\left(1+{\rm valNC}\right)$.
So far, we have established LABEL:Eq:p_q_star0tau_minus_p_q_star1tau holds for some distinct questions $q_{*0}\neq q_{*1}$. We assume that ${\cal A}_{2}$ can learn $p_{q^{\prime}\tau}$ exactly and invoke LABEL:Lem:3PrecisionNonIssue to handle the fact that these can only be approximated to inverse polynomial errors but that does not change the conclusion. Therefore, ${\cal A}_{2}$ can learn $q_{*0}\neq q_{*1}$ and construct the sets $T_{0},T_{1}$ of truth tables (recall: $\tau$ is in $T_{0}$ if $p_{q_{*0}\tau}\geq p_{q_{*1}\tau}$ and in $T_{1}$ otherwise). Proceeding as in the proof of LABEL:Prop:DINDprime, and focusing on Phase 2 of the interaction, it holds that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2},C_{2}\right\rangle]$ $=\frac{1}{2}\cdot\sum_{\tau\in T_{0}}\Pr[{\cal A}\text{ outputs }\tau|q_{*0}\text{ was encrypted}]+$
$\ \ \frac{1}{2}\cdot\sum_{\tau\in T_{1}}\Pr[{\cal A}\text{ outputs }\tau|q_{*1}\text{ was encrypted}]$
$=\frac{1}{2}+\frac{1}{2}\sum_{\tau\in T_{0}}\left(p_{q_{*0}\tau}-p_{q_{*1}\tau}\right)$
$=\frac{1}{2}+\frac{1}{4}\left\|p_{q_{*0}}-p_{q_{*1}}\right\|_{1}$
$=\frac{1}{2}+\frac{\eta}{4}.$
Since the $\mathsf{QFHE}$ scheme satisfies LABEL:Eq:distinguisher-1, we have a contradiction (via LABEL:Claim:2-IND_QFHE) which means our assumption that ${\cal A}$ wins with probability non-negligibly more than $\frac{1}{2}\left(1+{\rm valNC}\right)$ is false, completing the proof. ∎
9.5 Proof of the lemmas (Step 2 of 2)
LABEL:Lem:1guessK=00003DgoodK, about using a uniformly random $k$ instead of the correct one, is almost immediate but we include a brief proof.
Proof of LABEL:Lem:1guessK=00003DgoodK.
Consider the adversary ${\cal A}_{\mathsf{OPad}}$ for $\mathsf{OPad}$ as defined in LABEL:Alg:red_for_lemma1.
${\cal A}$ ${\cal A}_{\mathsf{OPad}}$ $C_{\mathsf{OPad}}$
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c_{q^{\prime}}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$ $\xleftarrow{\mathsf{OPad}.\mathsf{sk}}$
$\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$ $\overset{s^{\prime}}{\longrightarrow}$
$b\leftarrow\{0,1\}$, $k_{0}\leftarrow K$, $k_{1}=\mathsf{OPad}.\mathsf{Dec}(\mathsf{OPad}.\mathsf{sk},s^{\prime})$
$\overset{k_{b}}{\longleftarrow}$
$B$ defines $k^{\prime}:=k_{b}$ and uses the secret key $\mathsf{sk}$ to compute $k^{\prime\prime}$ and then finds the $k$ satisfying $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$.
[…]
Runs an arbitrary procedure to compute a bit $b^{\prime}$.
$\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
Algorithm 9 Whether a PPT adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ is used with the correct $k$ or a uniformly random $k$, it makes no difference, if all algorithms involved are PPT.
Observe that
$\Pr\left[\mathsf{accept}\leftarrow\left\langle{\cal A}_{\mathsf{OPad}},C_{\mathsf{OPad}}\right\rangle\right]=$ $\frac{1}{2}\Pr[{\cal A}_{\mathsf{OPad}}\text{ outputs }b^{\prime}=0|b=0]+$
$\frac{1}{2}\Pr[{\cal A}_{\mathsf{OPad}}\text{ outputs }b^{\prime}=1|b=1]$
$=$ $\frac{1}{2}\Big{(}\Pr[{\cal A}_{\mathsf{OPad}}\text{ outputs }b^{\prime}=0|b=0]-$
$\ \ \Pr[{\cal A}_{\mathsf{OPad}}\text{ outputs }b^{\prime}=0|b=1]\Big{)}+\frac{1}{2}$
$=$ $\frac{1}{2}\left(\Pr[0\leftarrow\left\langle B_{0},{\cal A}\right\rangle]-\Pr[0\leftarrow\left\langle B_{1},{\cal A}\right\rangle]\right)+\frac{1}{2}$ (33)
where $B_{0}$ and $B_{1}$ interact with ${\cal A}$ as in LABEL:Alg:random-k-or-not. The last equality holds because for a random $k^{\prime}$, $k$ also becomes random.
Let ${\cal A}_{\mathsf{OPad}}^{\prime}$ be ${\cal A}_{\mathsf{OPad}}$ except that it outputs $b^{\prime}\oplus 1$ instead of $b^{\prime}$. Using the analogue of LABEL:Eq:securityLemma1 for ${\cal A}^{\prime}_{\mathsf{OPad}}$, LABEL:Eq:securityLemma1 itself and the security of $\mathsf{OPad}$, it follows that there is a negligible function $\mathsf{negl}$ such that
$\left|\Pr[0\leftarrow\left\langle B_{0},{\cal A}\right\rangle]-\Pr[0\leftarrow\left\langle B_{1},{\cal A}\right\rangle]\right|\leq\mathsf{negl}.$
We now look at the proof of LABEL:Lem:2EstimateAssumingConsistent which crucially relies on the fact that the consistency test and the predicate test happen with equal probability.
Proof of LABEL:Lem:2EstimateAssumingConsistent.
Consider the adversary ${\cal A}^{\prime}$ in LABEL:Alg:red_for_lemma2 that uses ${\cal A}$ to interact with ${\cal C}_{k\leftarrow K}=:{\cal C}^{\prime}$.
${\cal A}$ ${\cal A}^{\prime}$ ${\cal C}_{k\leftarrow K}=:{\cal C}^{\prime}$
$\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$ $\xleftarrow{(c_{q^{\prime}},\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{a^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
$\tilde{k}\leftarrow K$
$\overset{(\tilde{q},\tilde{k})}{\circlearrowleft}$ Rewinds ${\cal A}^{\prime}$ to learn the truth table $\tau$.
Evaluate $c_{\tau(q^{\prime})}\leftarrow\mathsf{QFHE}.\mathsf{Eval}_{\tau}(c_{q^{\prime}}).$ $\xrightarrow{(c_{\tau(q^{\prime})},\hat{k}^{\prime\prime},s^{\prime})}$
$\xleftarrow{(q,k)}$
$\xrightarrow{a:=\tau(q)}$ Recall: ($a^{\prime}=\mathsf{Dec}_{\mathsf{sk}}(c_{\tau(q^{\prime})})$ If $q=q^{\prime}$, accept if $a=a^{\prime}$. If $q\neq q^{\prime}$, accept if ${\rm pred}((a,a^{\prime}),C)=1$.
Algorithm 10 ${\cal A}^{\prime}$, a potentially QPT algorithm, uses the PPT algorithm ${\cal A}$ to play the compiled contextuality game $\mathsf{G}^{\prime}$. Its winning probability upper bounds that of ${\cal A}$ and can be computed in terms of $p_{q^{\prime}\tau}$ for ${\cal A}$.
We show that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}^{\prime}\right\rangle]$ $\leq\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}^{\prime},{\cal C}^{\prime}\right\rangle]$ (34)
$=\sum_{C}\Pr(C)\cdot\frac{1}{2}\left(1+\frac{1}{|C|}\sum_{q^{\prime}\in C}\sum_{\tau}p_{q^{\prime}\tau}\mathsf{pred}(\tau[C],C)\right).$ (35)
We note that ${\cal A}^{\prime}$ may be a QPT algorithm because it runs $\mathsf{QFHE}.\mathsf{Eval}$. Indeed, existing $\mathsf{QFHE}$ schemes don’t produce a classical procedure for $\mathsf{QFHE}.\mathsf{Eval}$, even when the ciphertext and the logical circuit are classical. However, we only use ${\cal A}^{\prime}$ to compute an upper bound on the performance of ${\cal A}$ in terms of $p_{q^{\prime}\tau}$ and therefore ${\cal A}^{\prime}$ being QPT (as opposed to PPT) is not a concern here.
We first derive LABEL:Eq:Aprime_does_just_as_well_or_better. Consider the interactions $\left\langle{\cal A},{\cal C}^{\prime}\right\rangle$ and $\left\langle{\cal A}^{\prime},{\cal C}^{\prime}\right\rangle$. Note that for any given $c_{q^{\prime}}$, the challenger ${\cal C}^{\prime}$ either asks $q=q^{\prime}$ or $q\neq q^{\prime}$ with equal probabilities. Conditioned on $c_{q^{\prime}}$, there are two cases: (1) ${\cal A}$ is consistent, in which case the answers to $q,q^{\prime}$ given by ${\cal A}^{\prime}$ and ${\cal A}$ are identical. (2) ${\cal A}$ is inconsistent, in which case, ${\cal C}^{\prime}$ accepts ${\cal A}$ with probability at most $1/2$ while ${\cal C}^{\prime}$ accepts ${\cal A}^{\prime}$ with probability at least $1/2$.
LABEL:Eq:Aprime_wins_with_atmost follows because ${\cal C}^{\prime}$ selects a context $C$ with probability $\Pr(C)$, the $1/2$ denotes whether a consistency test ($q=q^{\prime}$) is performed or a predicate test ($q\neq q^{\prime}$) is performed. By construction of ${\cal A}^{\prime}$, the consistency test passes with probability $1$. The $1/|C|$ factor indicates that either of the two questions in $C$ could have been asked as the first question $q^{\prime}$ (under the $\mathsf{QFHE}$ encryption; and $|C|=2$ here). Again, by construction of ${\cal A}^{\prime}$, the probability of clearing the predicate test is the weighted average of $\mathsf{pred}(\tau[C],C)$ where the weights are given by $p_{q^{\prime}\tau}$. This completes the proof. ∎
LABEL:Lem:3PrecisionNonIssue follows by proceeding as in the proof of LABEL:Prop:finitePrecisionD-IND_prime.
Part III General Computational Test of Contextuality—beyond size-$2$ contexts
So far, we looked at contextuality games with size $2$ contexts. This part explains how to generalise our compiler to games with contexts of arbitrary size. In fact, we consider two generalisations of our compiler, both of which produce single prover, 4-message (2-round) compiled games, irrespective of the size of contexts in the original contextuality game.
The $(|C|,1)$ compiler.
The compiled game asks all $|C|$ questions from one context under the $\mathsf{QFHE}$ encryption in the first round, and asks one question in the clear in the second round. It guarantees that no PPT algorithm can succeed with probability more than
$1-\mathsf{const}_{1}+\mathsf{negl}$
[…]
As stated in the introduction, while this compiler works for games with perfect completeness, the bound on the classical value is sometimes more than the honest quantum value—so this compiler fails to give a separation between quantum and classical for some games (including the KCBS game of LABEL:Exa:KCBSintro). However, for other games (like the magic square game of LABEL:Exa:MagicSquare, it yields a larger completeness-soundness gap than the following universal compiler.
The $(|C|-1,1)$ compiler.
This compiler addresses the limitation of the one above and is truly universal—in the sense that for any contextuality game $\mathsf{G}$ with ${\rm valNC}<{\rm valQu}$, the compiled game $\mathsf{G}^{\prime}$ will also be such that every PPT algorithm wins with probability strictly smaller than a QPT algorithm. More precisely, PPT algorithms cannot succeed with probability more than
$\frac{1}{|C|}\left(|C|-1+{\rm valNC}\right)+\mathsf{negl}$
while there is a QPT algorithm that wins with probability at least
$\frac{1}{|C|}\left(|C|-1+{\rm valQu}\right)-\mathsf{negl}.$
Note that for $|C|=2$, we recover
$\frac{1}{2}\left(1+{\rm valNC}\right)+\mathsf{negl},\quad\frac{1}{2}(1+{\rm valQu})-\mathsf{negl}$
respectively, which are the bounds for our $(1,1)$ compiler.
The idea behind the construction is the following:
  • Sample a context $C$ and pick a question $q_{\mathsf{skip}}\leftarrow C$ uniformly at random.
  • Ask all questions in $C$ except $q_{\mathsf{skip}}$, i.e. ask $C\backslash q_{\mathsf{skip}}$, under $\mathsf{QFHE}$ encryption
  • Sample $q\leftarrow C$ and ask $q$ in the clear.
  • Now,
    1. 1.
      if $q=q_{\mathsf{skip}}$, test the predicate using the decrypted answers to $C\backslash q_{\mathsf{skip}}$ and to $q$;
    2. 2.
      if $q\neq q_{\mathsf{skip}}$, and so $q\in C\backslash q_{\mathsf{skip}}$; check that the corresponding answers are consistent.
The proof is based on the following idea.
  • The key observation is that, just as in the $(1,1)$ compiler, given an adversary ${\cal A}$, one can consider an adversary ${\cal A}^{\prime}$ such that it is “consistent” and wins with at least as much probability as ${\cal A}$ and using this, one can obtain a bound on the success probability of ${\cal A}$ in terms of ${\rm valNC}$.
  • The compiled game was purposefully designed to ensure the following: being “consistent” can only help. As we saw, this was not the case for the $(|C|,1)$ game—so selecting fewer questions in the first round is what, perhaps surprisingly, allows us to construct a universal test (in the sense described above).
Before moving to formal descriptions and proofs, we remark that it would be nice to have the success probability be independent of $|C|$, but this somehow seems hard to avoid. Why? Because the dependence on $|C|$ comes essentially from the fact the we are only asking one question in the clear in the second round. So, it seems that to avoid this dependence one would have to ask more than one question in the clear. However, this complicates the task of extracting a non-contextual assignment by rewinding the PPT adversary, because, for instance, the prover can now assign different values to $q_{1}$ depending on whether it was asked with $q_{2}$ or $q_{3}$. Thus, obtaining a better compiler seems to require new ideas.
10 Construction of the $(|C|,1)$ compiler
We define the compiler formally first.
Honest Prover (${\cal A}$) Challenger (${\cal C}$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $C\leftarrow{\cal D}$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C)$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
Under the QFHE encryption, measures $\{O_{q}\}_{q\in C}$ and obtains an encrypted answers $\{c_{\mathbf{a}}\}$ (where $\mathbf{a}$ are the answers indexed by $C$) and post-measurement state $(U_{k^{\prime\prime}}\left|\psi_{\mathbf{a}C}\right\rangle,\hat{k}^{\prime\prime})$. Applies an oblivious $\mathbf{U}$-pad to this state to obtain  $\left(U_{k^{\prime}}U_{k^{\prime\prime}}\left|\psi_{\mathbf{a}C}\right\rangle,s^{\prime}\right)$$\leftarrow$ $\mathsf{OPad}.\mathsf{Enc}(\mathsf{OPad}.\mathsf{pk},U_{k^{\prime\prime}}\left|\psi_{\mathbf{a}C}\right\rangle).$
$\xrightarrow{(c_{\mathbf{a}},\hat{k}^{\prime\prime},s^{\prime})}$
Using the secret keys $\mathsf{sk},\text{ \& }\mathsf{OPad}.\mathsf{sk}$, finds the $k$ such that $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$, samples $q\leftarrow C$
$\xleftarrow{(q,k)}$
Measures $U_{k}O_{q}U_{k}^{\dagger}$ and obtains $a$
[…]
Computes $\mathbf{a}=\mathsf{Dec}_{\mathsf{sk}}(c_{\mathbf{a}})$. Accepts if both (1) and (2) hold: (1) $\mathsf{pred}(\mathbf{a},C)=1$, and (2) $\mathbf{a}[q]=a$.
Algorithm 11 Game $\mathsf{G}^{\prime}$ produced by the $(|C|,1)$-compiler on input a contextuality game $\mathsf{G}$, and security parameter $\lambda$.
10.1 Compiler Guarantees
The compiler satisfies the following.
Theorem 31 (Guarantees of the $(|C|,1)$ compiled contextuality game $\mathsf{G}^{\prime}$).
Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$. Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:C-1-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
  • (Completeness) There is a negligible function $\mathsf{negl}$, such that, for all $\lambda\in\mathbb{N}$, the honest QPT prover from LABEL:Alg:C-1-compiler wins $\mathsf{G}^{\prime}_{\lambda}$ with probability at least
    ${\rm valQu}-\mathsf{negl}(\lambda).$
  • (Soundness) For every PPT adversary $\cal A$, there is a negligible function $\mathsf{negl}^{\prime}$ such that, for all $\lambda\in\mathbb{N}$, the probability that $\cal A$ wins $\mathsf{G}^{\prime}_{\lambda}$ is at most
    $1-\mathsf{const}_{1}+\mathsf{negl}^{\prime}(\lambda)\,,$
    where $\mathsf{const}_{1}=\min_{C\in C^{{\rm all}}}\Pr(C)/|C|=O(1)$, assuming $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad).
Completeness is straightforward to verify. We prove soundness in Section 11.
Note that in games where all questions are asked uniformly at random, $\mathsf{const}_{1}=1/(|C^{{\rm all}}|\cdot|C|)$, yielding the simple bound $1-1/(|C^{{\rm all}}|\cdot|Q|)+\mathsf{negl}$ for PPT provers (that we quoted earlier).
Two brief remarks about the theorem are in order before we give the proof. First, we note that unlike the completeness value, which is ${\rm valQu}-\mathsf{negl}$, the soundness does not correspondingly depend on ${\rm valNC}$. Second, as mentioned earlier, this compiler is not universal: for KCBS the classical value in $\mathsf{G}^{\prime}$ is $0.9$ which is greater than the honest quantum value which is approximately $0.8944$. However, the compiler is still outputs a non-trivial game, for instance, when given as input the GHZ game, because the latter has perfect completeness.
11 Soundness Analysis of the $(|C|,1)$ compiler
The proof is analogous to that of the $(1,1)$ compiler with some crucial differences and, thus, we skip the intuition and details for parts that are essentially unchanged.
11.1 The Reduction
[…]
The analogue of LABEL:Lem:2EstimateAssumingConsistent is slightly different. Instead of having the adversary be consistent (in the sense that the encrypted answers and the clear answers are always the same by construction), we require the complementary property—the adversary always ensures that the encrypted answer satisfies the predicate, but may not be consistent. Since we now actually rely on the encrypted answers to check whether or not the predicate is satisfied in the analysis, we cannot simply drop the encrypted answer.
Here is the “TruthTable” algorithm and the actual reduction ${\cal A}_{2}$.
${\cal A}$ ${\cal A}_{2}.\mathsf{TruthTable}(C)$
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}(C)$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{\mathbf{a}},\hat{k}^{\prime\prime},s^{\prime})}$
Samples $k\leftarrow K$
[…]
Outputs $\mathbf{a}:=\mathsf{QFHE}.\mathsf{Dec}_{\mathsf{sk}}(\mathbf{a})$ and $\tau$.
Algorithm 12 Analogue of LABEL:Alg:ATruthTable, except that it explicitly outputs $\mathbf{a}$, in addition to $\tau$.
  • The definition of $C_{*0},C_{*1}\in C^{{\rm all}}$ is now “indices” such that
    $\sum_{\tau}\left|\sum_{\mathbf{a}}p_{C_{*0}\mathbf{a}\tau}-\sum_{\mathbf{a}}p_{C_{*1}\mathbf{a}\tau}\right|\geq\eta$
    where $\eta$ is non-negligible.
  • $T_{0}$ and $T_{1}$ are defined as $\tau\in T_{0}$ if $\sum_{\mathbf{a}}p_{C_{*0}\mathbf{a}\tau}\geq\sum_{\mathbf{a}}p_{C_{*1}\mathbf{a}\tau}$ and $\tau\in T_{1}$ otherwise, i.e. $\sum_{\mathbf{a}}p_{C_{*0}\mathbf{a}\tau}<\sum_{\mathbf{a}}p_{C_{*1}\mathbf{a}\tau}$.
${\cal A}$ ${\cal A}_{2}$
Phase 1
[…]
Use $p_{C\mathbf{a}\tau}$ to learn the contexts $C_{*0},C_{*1}$ and the disjoint sets $T_{0},T_{1}$ of truth tables.
${\cal A}$ Phase 2 $C_{2}$
$\xrightarrow{(C_{*0},C_{*1})}$
$(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$ $\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $b\leftarrow\{0,1\}$ $c_{b}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C_{*b})$
$\xleftarrow{(c_{b},\mathsf{OPad}.\mathsf{pk})}$ $\overset{c_{b}}{\longleftarrow}$
$\xrightarrow{(c_{\mathbf{a}},\hat{k}^{\prime\prime},s^{\prime})}$
Samples $k\leftarrow K$
[…]
Set $b^{\prime}=0$ if $\tau\in T_{0}$, and $b^{\prime}=1$ if $\tau\in T_{1}$ $\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
Algorithm 13 The algorithm ${\cal A}_{2}$ uses the adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$, to break the $2$-IND security game for the $\mathsf{QFHE}$ scheme.
11.2 Proof Strategy
Lemma 32 ([Analogue of LABEL:Lem:1guessK=00003DgoodK] Uniformly random $k$ is equivalent to the correct $k$).
Let $B_{0}$ (resp. $B_{1}$) be a PPT algorithm that takes $q^{\prime}\in Q$ as an input, interacts with ${\cal A}$ and outputs a bit, as described in LABEL:Alg:random-k-or-not. Then there is a negligible function $\mathsf{negl}$ such that $\left|\Pr[0\leftarrow\left\langle B_{0},{\cal A}\right\rangle]-\Pr[0\leftarrow\left\langle B_{1},{\cal A}\right\rangle]\right|\leq\mathsf{negl}$.
${\cal A}$ $B_{0}(C)$ (resp. $B_{1}(C)$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C)$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{\mathbf{a}},\hat{k}^{\prime\prime},s^{\prime})}$
$B_{0}$ computes $k^{\prime}:=\mathsf{OPad}.\mathsf{Dec}(\mathsf{OPad}.\mathsf{sk},s^{\prime})$ and uses the secret key $\mathsf{sk}$ to compute $k^{\prime\prime}$ and then finds the $k$ satisfying $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$. (resp. $B_{1}$ samples a uniform $k\leftarrow K$).
$\overset{(q,k)}{\circlearrowleft}$ Potentially rewinds ${\cal A}$ to this step and queries with $(q,k)$ for arbitrary $q\in Q$.
Runs an arbitrary procedure to compute a bit $b^{\prime}$.
Algorithm 14 Whether a PPT adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ is used with the correct $k$ or a uniformly random $k$, it makes no difference, if all algorithms involved are PPT.
We are finally ready to write the first step which is conceptually different from the previous analysis. It is the analogue of LABEL:Lem:2EstimateAssumingConsistent—except that:
  • Earlier, we obtained the bound by essentially treating ${\cal A}$ as being consistent (with the answers under the encryption and those in the clear) where by virtue of being consistent, ${\cal A}$ could not be feasible i.e. ${\cal A}$ could not satisfy all the predicates.
  • Now, we treat ${\cal A}$ as being feasible (satisfies all predicates) , but by virtue of being feasible, it cannot be consistent (can’t have the encrypted answers be consistent with a global truth assignment).
Lemma 33 ([Analogue of LABEL:Lem:2EstimateAssumingConsistent] Feasibility only helps).
Let
  • $\mathsf{G}^{\prime}$ be the compiled game as in LABEL:Thm:C_1_CompiledGameIsSecure, let ${\cal C}$ be the challenger in $\mathsf{G}^{\prime}$, and let
  • ${\cal C}_{k\leftarrow K}$ be exactly the same as the challenger ${\cal C}$ except that it samples a uniformly random $k$ instead of computing it correctly,
  • ${\cal A}$ be any PPT algorithm that is designed to play the complied contextuality game $\mathsf{G}^{\prime}$ and makes ${\cal C}_{k\leftarrow K}$ accept with probability $p$, i.e. $\Pr\left[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle\right]=p$ and denote by
  • $p_{C\mathbf{a}\tau}$ be the probability that on being asked the context $C$, the answers given are $\mathbf{a}$ and the truth table produced is $\tau$ (as defined in LABEL:Subsec:C_1_The-Reduction)
  • $p^{\prime}_{C\mathbf{a}\tau}$ be a probability distribution derived from $p_{C\mathbf{a}\tau}$ to be such that $p^{\prime}$ is always feasible and whenever $p$ has support on $C,\mathbf{a}$ that satisfy the corresponding predicate, $p^{\prime}$ and $p$ agree. More precisely, for any $(C,\mathbf{a},\tau)$ satisfying $p_{C\mathbf{a}\tau}>0$ it holds that
    • ($p^{\prime}$ matches $p$ exactly when $p$ is feasible) if $p_{C\mathbf{a}\tau}>0\land\mathsf{pred}(\mathbf{a}[C],C)=1$,
        $p_{C\mathbf{a}\tau}^{\prime}=p_{C\mathbf{a}\tau}$ and
    • ($p^{\prime}$ is always feasible) if $\mathsf{pred}(\mathbf{a}[C],C)=0$,
        $p^{\prime}_{C\mathbf{a}\tau}=0$
    • (when $p$ is infeasible, $p^{\prime}$ preserves probabilities but makes it feasible) if $p_{C\mathbf{a}\tau}>0\land\mathsf{pred}(\mathbf{a}[C],C)=0$, there is an [^27] The first property implies $\mathsf{pred}(\mathbf{a}^{\prime}[C],C)=1$. $\mathbf{a}^{\prime}$ such that
        $p^{\prime}_{C\mathbf{a}^{\prime}\tau}=p_{C\mathbf{a}\tau}$.
Then
$p\leq\sum_{C\mathbf{a}\tau}p_{C\mathbf{a}\tau}^{\prime}\Pr(C)\frac{1}{|C|}\sum_{q\in C}\delta_{\mathbf{a}[q],\tau(q)}$
where $\Pr(C)$ denotes the probability with which ${\cal C}$ samples the context $C$.
The analogue of LABEL:Lem:3PrecisionNonIssue should go through with almost no changes.
Lemma 34 ([Analogue of LABEL:Lem:3PrecisionNonIssue] Precision is not an issue).
Exactly as LABEL:Lem:3PrecisionNonIssue except that ${\cal C}$ is the challenger for the game compiled using the $(|C|,1)$ compiler.
As before, we prove the contrapositive of the soundness condition.
Theorem 35 (Soundness condition restated from LABEL:Thm:C_1_CompiledGameIsSecure).
Suppose
  • ${\cal A}$ is any PPT algorithm that wins with probability
    $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq 1-\mathsf{const}_{1}+\epsilon$
    for some non-negligible function $\epsilon$ where $\mathsf{const}_{1}=\min_{C\in C^{{\rm all}}}\Pr(C)/|C|=O(1)$, and
  • the $\mathsf{OPad}$ used is secure, then
there is a PPT algorithm ${\cal A}_{2,\epsilon}$ that wins the $2$-IND security game of the $\mathsf{QFHE}$ scheme with probability
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2,\epsilon},C_{2}\right\rangle]\geq\frac{1}{2}+\mathsf{nonnegl},$
where $\mathsf{nonnegl}$ is a non-negligible function that depends on $\epsilon$.
In LABEL:Subsec:C1_proof_step1, we prove LABEL:Thm:C1_MainSoundness assuming LABEL:Lem:1_C1_guessk=00003Dgoodk,_2EstimateAssumingConsistent,_3PrecisionNonIssue. In LABEL:Subsec:C1_proof_step2 we prove the lemmas.
11.3 Proof assuming the lemmas (Step 1 of 2)
We introduce some notation. Recall the definition of $p_{C\mathbf{a}\tau}$ from LABEL:Subsec:C_1_The-Reduction and of $p_{C\mathbf{a}\tau}^{\prime}$ from LABEL:Lem:2_C1_feasibility_only_helps.
  • We use the notation $p_{\mathbf{a}\tau|C}:=p_{C\mathbf{a}\tau}$ since $C$ was given as input to the procedure ${\cal C}_{2}\cdot\mathsf{TruthTable}$ and it produced outputs $\mathbf{a},\tau$. Similarly define $p^{\prime}_{\mathbf{a}\tau|C}:=p^{\prime}_{C\mathbf{a}\tau}$.
  • We also use
    $p_{\mathbf{a}\tau|C}=p_{\mathbf{a}|\tau C}\cdot p_{\tau|C}\text{ and }p_{\mathbf{a}\tau|C}^{\prime}=p_{\mathbf{a}|\tau C}^{\prime}\cdot p_{\tau|C}^{\prime}$ (36)
    to denote conditionals. [^28] Using $p(a,b|c)=p(a|b,c)\cdot p(b|c)=\frac{p(a,b,c)}{p(b,c)}\cdot\frac{p(b,c)}{p(c)}$.
  • Finally, we use the convention of denoting marginals by dropping the corresponding index, i.e.
    $p_{\tau|C}:=\sum_{\mathbf{a}}p_{\mathbf{a}\tau|C}\text{ and }p_{\tau|C}^{\prime}:=\sum_{\mathbf{a}}p_{\mathbf{a}\tau|C}^{\prime}.$
  • Note that by definition of $p^{\prime}_{\mathbf{a}\tau|C}$, it holds that
    $p^{\prime}_{\tau|C}=p_{\tau|C}.$ (37)
    Proof.
    From LABEL:Lem:1_C1_guessk=00003Dgoodk, one concludes that
    $\left|[\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]-\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle]\right|\leq\mathsf{negl}.$
    Recall the definition of $p_{C\mathbf{a}\tau}$ from LABEL:Subsec:C_1_The-Reduction. Using LABEL:Lem:2EstimateAssumingConsistent, one can write
    $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\leq$ $\sum_{C\mathbf{a}\tau}p_{\mathbf{a}\tau|C}^{\prime}\Pr(C)\frac{1}{|C|}\sum_{q\in C}\delta_{\mathbf{a}[q],\tau[q]}+\mathsf{negl}$ (38)
    $=$ $\sum_{C\tau}p{}_{\tau|C}\sum_{\mathbf{a}}p_{\mathbf{a}|\tau C}^{\prime}\Pr(C)\frac{1}{|C|}\sum_{q\in C}\delta_{\mathbf{a}[q],\tau[q]}+\mathsf{negl}$ using (36) & (37) (39)
    Observe also that there exist $C_{*0}\neq C_{*1}$ such that
    $\sum_{\tau}\left|p_{\tau|C_{*0}}-p_{\tau|C_{*1}}\right|\geq\eta$ (40)
    for some non-negligible function $\eta$. This is a consequence of the assumption that
    $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq 1-\mathsf{const}_{1}+\epsilon$ (41)
    for some non-negligible function $\epsilon$. To see this, proceed by contradiction: Suppose that for all $C_{*0}\neq C_{*1}$, it is the case that
    $\sum_{\tau}|p_{\tau|C_{*0}}-p_{\tau|C_{*1}}|\leq\mathsf{negl}$ (42)
    then, setting $p_{\tau}:=p_{\tau|C_{*0}}$ (for instance), it holds that (using LABEL:Eq:C1_pr_acc_further)
    $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]$ $\leq\sum_{\tau}p{}_{\tau}\sum_{C}\Pr(C)\sum_{\mathbf{a}}p_{\mathbf{a}|\tau C}^{\prime}\sum_{q\in C}\frac{\delta_{\mathbf{a}[q],\tau[q]}}{|C|}+\mathsf{negl}^{\prime}$
    $\leq\sum_{\tau}p_{\tau}\Bigg{(}\underbrace{\sum_{C\neq C_{\tau}}\Pr(C)\cdot 1}_{\text{Term 1}}+\underbrace{\Pr(C_{\tau})\cdot\left(1-\frac{1}{|C|}\right)}_{\text{Term 2}}\Bigg{)}+\mathsf{negl}^{\prime}$
    $\leq\sum_{\tau}p_{\tau}\left(1-\frac{\Pr(C_{\tau})}{|C|}\right)+\mathsf{negl}^{\prime}\leq 1-\frac{\min_{C\in C^{{\rm all}}}\Pr(C)}{|C|}+\mathsf{negl}^{\prime}=1-\mathsf{const}_{1}+\mathsf{negl}^{\prime}$ (43)
    where the first line uses LABEL:Eq:C1_p_tau_C_negligible to substitute $p_{\tau|C}$ with $p_{\tau}$ at the cost a negligible factor but the second line needs some explanation. Observe that (a) for each truth table $\tau$, there is a context $C_{\tau}\in C^{{\rm all}}$ such that $\mathsf{pred}(\tau[C],C)=0$ because ${\rm valNC}<1$. Observe also that (b) by assumption on $p^{\prime}_{\mathbf{a}\tau|C}$ (recall LABEL:Lem:2_C1_feasibility_only_helps), all $\mathbf{a}$ with non-zero weight are feasible, i.e. $\mathsf{pred}(\mathbf{a}[C_{\tau}],C_{\tau})=1$ for any $p^{\prime}_{\mathbf{a}|\tau C}>0$. This implies that there is at least one question $q\in C_{\tau}$ such that $\mathbf{a}[q]\neq\tau[q]$ (i.e. $\delta_{\mathbf{a}[q],\tau[q]}=0$)—else $\tau$ would also have satisfied the predicate on $C_{\tau}$ which it does not. Using (a) and (b), in Term 1, we simply upper bound the remaining sum by $1$ while in term $2$, we upper bound the sum by concluding that for at least one question in $C_{\tau}$, the delta function vanishes. The last inequality follows by setting term 1 to be $1-\Pr(C_{\tau})$ and rearranging and minimising. Now, since LABEL:Eq:C1_atmost_1-const_1 contradicts LABEL:Eq:C1_A_wins_against_C-w_non-negl we conclude that our assumption LABEL:Eq:C1_p_tau_C_negligible must be false, establishing LABEL:Eq:p_tau_C_star_0_minus_p_tau_C_star_1.
    The remaining analysis goes through almost unchanged from the $(1,1)$ compiler case.
    $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2},C_{2}\right\rangle]=$ $\frac{1}{2}\cdot\sum_{\tau\in T_{0}}\Pr[{\cal A}\text{ outputs }\tau|C_{*0}\text{ was encrypted}]+$
    $\frac{1}{2}\cdot\sum_{\tau\in T_{1}}\Pr[{\cal A}\text{ outputs }\tau|C_{*1}\text{ was encrypted}]$
    $=$ $\frac{1}{2}+\frac{1}{2}\sum_{\tau\in T_{0}}(p_{\tau|C_{*0}}-p_{\tau|C_{*1}})$
    $=$ $\frac{1}{2}+\frac{1}{4}\sum_{\tau}\left|p_{\tau|C_{*0}}-p_{\tau|C_{*1}}\right|=\frac{1}{2}+\frac{\eta}{4}.$
    This, together with LABEL:Lem:3_C1_precision_non_issue, yields a contradiction with the security of the $\mathsf{QFHE}$ scheme. We therefore conclude that LABEL:Eq:C1_A_wins_against_C-w_non-negl is false, which means $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\leq 1-\mathsf{const}_{1}+\mathsf{negl}$. ∎
11.4 Proof of the lemmas (Step 2 of 2)
The proofs of LABEL:Lem:1_C1_guessk=00003Dgoodk and LABEL:Lem:3_C1_precision_non_issue are analogous to those of LABEL:Lem:1guessK=00003DgoodK and LABEL:Lem:3PrecisionNonIssue. Here, we prove LABEL:Lem:2_C1_feasibility_only_helps.
Proof of LABEL:Lem:2_C1_feasibility_only_helps.
${\cal A}$ ${\cal A}^{\prime}$ ${\cal C}_{k\leftarrow K}=:{\cal C}^{\prime}$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$ $\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$ Recall that $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C)$.
$\xrightarrow{(c_{\mathbf{a}},\hat{k}^{\prime\prime},s^{\prime})}$
$\tilde{k}\leftarrow K$
$\overset{(\tilde{q},\tilde{k})}{\circlearrowleft}$ Rewinds ${\cal A}^{\prime}$ to learn the truth table $\tau$.
Under the QFHE encryption, with probability $p_{\mathbf{a}|\tau C}$ outputs $c_{\mathbf{a}^{\prime}}\leftarrow\mathsf{QFHE}.\mathsf{Eval}_{\mathbf{a}}(c)$ where $\mathbf{a}^{\prime}=\mathbf{a}$ if $\mathsf{pred}(\mathbf{a}[C],C)=1$, otherwise $\mathbf{a}^{\prime}=\mathbf{a}_{C}$. $\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
$\xleftarrow{(q,k)}$
$\xrightarrow{a:=\tau(q)}$ (Recall:) Accept if both (1) and (2) hold: (1) ${\rm pred}(\mathbf{a}^{\prime},C)=1$ and (2) $\mathbf{a}^{\prime}[q]=a$.
Algorithm 15 ${\cal A}^{\prime}$ uses ${\cal A}$ (a PPT algorithm—crucial because it is rewound), to play the compiled contextuality game $\mathsf{G}^{\prime}$. Its winning probability upper bounds that of ${\cal A}$ and can be computed in terms of $p_{\mathbf{a}\tau|C}$ for ${\cal A}$.
Let $p_{\mathbf{a}\tau|c}$ be as in LABEL:Subsec:C1_proof_step1 and recall that $p_{\mathbf{a}\tau|C}=p_{\mathbf{a}|\tau C}p_{\tau|C}$. For each $C\in C^{{\rm all}}$, denote by $\mathbf{a}_{C}$ answers such that $\mathsf{pred}(\mathbf{a}_{C},C)=1$. Consider the adversary ${\cal A}^{\prime}$ as in LABEL:Alg:C1_red_for_lemma2. Since this is just a way to compute an upper bound, the running time of ${\cal A}^{\prime}$ does not matter. We show that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}^{\prime}\right\rangle]\leq$ $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}^{\prime},{\cal C}^{\prime}\right\rangle]$
$=$ $\sum_{C}\Pr(C)\sum_{\tau}p_{\tau|C}\sum_{\mathbf{a}^{\prime}}p^{\prime}_{\mathbf{a}^{\prime}|\tau C}\sum_{q\in C}\delta_{\mathbf{a}^{\prime}[q],\tau[q]}$
where the first line holds because, by construction, ${\cal A}^{\prime}$ can do no worse than ${\cal A}$. More specifically, ${\cal A}^{\prime}$ behaves exactly like ${\cal A}$ when $\mathbf{a}$ satisfies the predicate, i.e. $\mathbf{a}^{\prime}=\mathbf{a}$, and when $\mathbf{a}$ fails the predicate, ${\cal A}^{\prime}$ only increases its probability of success by responding with $\mathbf{a}^{\prime}\neq\mathbf{a}$ such that $\mathsf{pred}(\mathbf{a}^{\prime},C)=1$. The second line is straightforward. Denote by $p^{\prime}_{\tau|C}$ the probability with which ${\cal A}^{\prime}$ responds with $\tau$ on being asked $C$. Similarly, let $p^{\prime}_{\mathbf{a}^{\prime}|\tau C}$ be the probability that ${\cal A}^{\prime}$ responds with $\mathbf{a}^{\prime}$ given $\tau$ and $C$. Then, the challenger ${\cal C}^{\prime}$ asks a context $C$ with probability $\Pr(C)$, to which the prover ${\cal A}^{\prime}$ responds with $\tau$ with probability $p^{\prime}_{\tau|C}=p_{\tau|C}$ and given $\tau C$, it responds with $\mathbf{a}^{\prime}$ with probability $p^{\prime}_{\mathbf{a}^{\prime}|\tau C}$. It follows that $p^{\prime}_{\mathbf{a}^{\prime}\tau|C}:=p^{\prime}_{\mathbf{a}^{\prime}|\tau C}p^{\prime}_{\tau|C}$ satisfies the asserted properties: it has no support over $(\mathbf{a}^{\prime},C)$ that are not feasible (don’t satisfy the predicate for the corresponding context $C$) and whenever $(\mathbf{a},C)$ is feasible, $\mathbf{a}^{\prime}=\mathbf{a}$ by construction. ∎
12 Construction of the $(|C|-1,1)$ Compiler
We define the compiler formally.
Honest Prover (${\cal A}$) Challenger (${\cal C}$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $C\leftarrow{\cal D}$ $q_{\mathsf{skip}}\leftarrow C$ $C^{\prime}:=C\backslash q_{\mathsf{skip}}$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
Under the QFHE encryption, measures $\{O_{q^{\prime}}\}_{q^{\prime}\in C^{\prime}}$ and obtains encrypted answers $c_{\mathbf{a}^{\prime}}$ (where $\mathbf{a}^{\prime}$ are the answers, indexed by $C^{\prime}$) and the post-measurement state $(U_{k^{\prime\prime}}\left|\psi_{C^{\prime}\mathbf{a}^{\prime}}\right\rangle,\hat{k}^{\prime\prime})$. Applies an oblivious $\mathbf{U}$-pad to this state to obtain  $\left(U_{k^{\prime}}U_{k^{\prime\prime}}\left|\psi_{C^{\prime}\mathbf{a}^{\prime}}\right\rangle,s^{\prime}\right)$$\leftarrow$ $\mathsf{OPad}.\mathsf{Enc}(\mathsf{OPad}.\mathsf{pk},U_{k^{\prime\prime}}\left|\psi_{q^{\prime}a^{\prime}}\right\rangle).$
$\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
Using the secret keys $\mathsf{sk},\mathsf{OPad}.\mathsf{sk}$, finds the $k$ such that $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$, samples $q\leftarrow C$
$\xleftarrow{(q,k)}$
Measures $U_{k}O_{q}U_{k}^{\dagger}$ and obtains $a$
[…]
Computes $\mathbf{a}^{\prime}=\mathsf{Dec}_{\mathsf{sk}}(c_{\mathbf{a}^{\prime}})$. If $q\in C^{\prime}$,    accept if $a=\mathbf{a}^{\prime}[q]$ If $q\notin C^{\prime}$ (i.e. $q=q_{\mathsf{skip}}$),    accept if ${\rm pred}(\mathbf{a}^{\prime}\cup a,C)=1$    where $\mathbf{a}^{\prime}\cup a$ denotes answers indexed by $C$.
Algorithm 16 Game $\mathsf{G}^{\prime}$ produced by the $(1,1)$-compiler for any contextuality game $\mathsf{G}$ with contexts of size two.
12.1 Compiler Guarantees
The compiler satisfies the following. Without loss of generality, we restrict to contextuality games where all contexts have the same size (see Remark 11).
Theorem 36 (Guarantees of the $(|C|-1,1)$ compiled contextuality game $\mathsf{G}^{\prime}$).
Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$ where all contexts are of the same size (i.e. $|C|=|C^{\prime}|$ for all $C,C^{\prime}\in C^{{\rm all}}$). Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:The-Cminus11-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
  • (Completeness) There is a negligible function $\mathsf{negl}$, such that, for all $\lambda\in\mathbb{N}$, the honest QPT prover from LABEL:Alg:The-Cminus11-compiler wins $\mathsf{G}^{\prime}_{\lambda}$ with probability at least
    $1-\frac{1}{|C|}+\frac{{\rm valQu}}{|C|}-\mathsf{negl}(\lambda).$
  • (Soundness) For every PPT adversary $\cal A$, there is a negligible function $\mathsf{negl}^{\prime}$ such that, for all $\lambda\in\mathbb{N}$, the probability that $\cal A$ wins $\mathsf{G}^{\prime}_{\lambda}$ is at most
    $1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\mathsf{negl}^{\prime}(\lambda)\,,$
    assuming $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad).
Completeness is straightforward to verify. We prove soundness in Section 13.
13 Soundness Analysis of the $(|C|-1,1)$ compiler
13.1 The Reduction
Denote by $\mathsf{C^{\prime all}}:=\{C\backslash q_{\mathsf{skip}}\}_{C\in C^{{\rm all}},q_{\mathsf{skip}}\in C}$ the set consisting of contexts with exactly one question removed.
The reduction is very similar to that in LABEL:Subsec:The-Reduction, so we do not repeat the accompanying high-level explanations. Let ${\cal A}_{2}\cdot\mathsf{TruthTable}$ be as defined in LABEL:Alg:Cminus11_A_2_reduction. Define $p_{C^{\prime}\tau}$ to be the probability that the procedure ${\cal A}_{2}\cdot\mathsf{TruthTable}(C^{\prime})$ outputs $\tau$, where $C^{\prime}\in\mathsf{C^{\prime all}}$ is a context with exactly one question removed.
${\cal A}$ ${\cal A}_{2}.\mathsf{TruthTable}(C^{\prime})$
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}(C^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
Samples $k\leftarrow K$
[…]
Outputs $\tau$
Algorithm 17 The procedure ${\cal A}_{2}.\mathsf{TruthTable}$ takes as input a context with one question excluded, $C^{\prime}=C\backslash q_{\mathsf{skip}}$. It produces a truth table $\tau$ corresponding to it. Note that this is a randomised procedure (depends on the $\mathsf{QFHE}$ encryption procedure) so for the same $C^{\prime}$ the procedure may output different $\tau$s. The goal is to learn the probabilities of different $\tau$s appearing for each question $C^{\prime}$.
${\cal A}$ ${\cal A}_{2}$
Phase 1
[…]
Use $p_{C^{\prime}\tau}$ to learn the questions $C^{\prime}_{*0},C^{\prime}_{*1}$ and the disjoint sets $T_{0},T_{1}$ of truth tables.
${\cal A}$ Phase 2 $C_{2}$
$\xrightarrow{(C^{\prime}_{*0},C^{\prime}_{*1})}$
$(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$ $\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $b\leftarrow\{0,1\}$ $c_{b}\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(q_{*b})$
$\xleftarrow{(c_{b},\mathsf{OPad}.\mathsf{pk})}$ $\overset{c_{b}}{\longleftarrow}$
$\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
Samples $k\leftarrow K$
[…]
Set $b^{\prime}=0$ if $\tau\in T_{0}$, and $b^{\prime}=1$ if $\tau\in T_{1}$ $\overset{b^{\prime}}{\longrightarrow}$
Accept if $b^{\prime}=b$
Algorithm 18 The algorithm ${\cal A}_{2}$ uses the adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$, to break the $2$-IND security game for the $\mathsf{QFHE}$ scheme.
13.2 Proof Strategy
Lemma 37 (Uniformly random $k$ is equivalent to the correct $k$).
Let $B_{0}$ (resp. $B_{1}$) be a PPT algorithm that takes $C^{\prime}\in\mathsf{C^{\prime all}}$ as an input, interacts with ${\cal A}$ and outputs a bit, as described in LABEL:Alg:Cminus11_random-k-or-not. Then, there is a negligible function $\mathsf{negl}$ such that $\left|\Pr[0\leftarrow\left\langle B_{0},{\cal A}\right\rangle]-\Pr[0\leftarrow\left\langle B_{1},{\cal A}\right\rangle]\right|\leq\mathsf{negl}$.
${\cal A}$ $B_{0}(C^{\prime})$ (resp. $B_{1}(C^{\prime})$)
$\mathsf{sk}\leftarrow\mathsf{QFHE}.\mathsf{Gen}(1^{\lambda})$ $c\leftarrow\mathsf{QFHE}.\mathsf{Enc}_{\mathsf{sk}}(C^{\prime})$ $(\mathsf{OPad}.\mathsf{pk},\mathsf{OPad}.\mathsf{sk})\leftarrow\mathsf{OPad}(1^{\lambda})$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$
$\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
$B_{0}$ computes $k^{\prime}:=\mathsf{OPad}.\mathsf{Dec}(\mathsf{OPad}.\mathsf{sk},s^{\prime})$ and uses the secret key $\mathsf{sk}$ to compute $k^{\prime\prime}$ and then finds the $k$ satisfying $U_{k}=U_{k^{\prime\prime}}U_{k^{\prime}}$. (resp. $B_{1}$ samples a uniform $k\leftarrow K$).
$\overset{(q,k)}{\circlearrowleft}$ Potentially rewinds ${\cal A}$ to this step and queries with $(q,k)$ for arbitrary $q\in Q$.
Runs an arbitrary procedure to compute a bit $b^{\prime}$.
Algorithm 19 Whether a PPT adversary ${\cal A}$ for the compiled contextuality game $\mathsf{G}^{\prime}$ is used with the correct $k$ or a uniformly random $k$, it makes no difference, if all algorithms involved are PPT.
The following we will check carefully at the end. It says that one can treat ${\cal A}$ as though it is consistent.
Lemma 38 (Consistency only helps).
Let
  • ${\cal C}_{k\leftarrow K}$ be exactly the same as the challenger ${\cal C}$ for $\mathsf{G}^{\prime}$ except that it samples $k\leftarrow K$ uniformly, instead of computing it correctly, let
  • ${\cal A}$ be any PPT algorithm that plays $\mathsf{G}^{\prime}$ and $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle]=p$ and denote by
  • $p_{C^{\prime}\tau}$ be as described above.
Then
$p\leq\left(1-\frac{1}{|C|}\right)+\sum_{C\in C^{{\rm all}}}\Pr(C)\sum_{q_{\mathsf{skip}}\in C}\frac{1}{|C|}\sum_{\tau}p_{C^{\prime},\tau}\mathsf{pred}(\tau[C],C)$
where $C^{\prime}=C\backslash q_{\mathsf{skip}}$, and $\Pr(C)$ denotes the probability with which ${\cal C}$ samples the context $C$.
The first term captures the probability that the consistency test passes and the second one captures the probability that the predicate test passes.
Lemma 39.
Exactly the same as LABEL:Lem:3PrecisionNonIssue except that
  • ${\cal C}$ is the challenger for the game produced by the $(|C|-1,1)$ compiler,
  • instead of $p_{q^{\prime}\tau}$, use $p_{C^{\prime}\tau}$, and
  • the construction of ${\cal A}_{2}$ is as in LABEL:Alg:Cminus11_A_2_reduction.
Let ${\cal A}$ be such that $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq 1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\epsilon$ and let ${\cal A}_{2,\epsilon}$ be as in LABEL:Lem:3PrecisionNonIssue. Then, it holds that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2,\epsilon},{\cal C}_{2}\right\rangle]\geq\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2},{\cal C}_{2}\right\rangle]-O(\epsilon^{3}).$
The following is the contrapositive of the soundness guarantee in LABEL:Thm:Cminus11_CompiledGameIsSecure.
Theorem 40 (Soundness condition restated from LABEL:Thm:Cminus11_CompiledGameIsSecure).
Suppose
  • ${\cal A}$ is any PPT algorithm that wins with probability $\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq 1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\epsilon$ for some non-negligible function $\epsilon$ and
  • the $\mathsf{OPad}$ used is secure, then
there is a PPT algorithm ${\cal A}_{2,\epsilon}$ that wins the $2$-IND security game of the $\mathsf{QFHE}$ scheme with probability
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2,\epsilon},C_{2}\right\rangle]\geq\frac{1}{2}+\mathsf{nonnegl},$
where $\mathsf{nonnegl}$ is a non-negligible function that depends on $\epsilon$.
In LABEL:Subsec:Cminus11_step1, we prove LABEL:Thm:Cminus11_MainSoundness assuming LABEL:Lem:Cminus11_1guessK=00003DgoodK,_Cminus11_2EstimateAssumingConsistent,_Cminus11_3PrecisionNonIssue. In LABEL:Subsec:Cminus11_step2, we prove the lemmas.
13.3 Proof assuming the lemmas (Step 1 of 2)
Proof.
From LABEL:Lem:Cminus11_1guessK=00003DgoodK, we have that $|\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]-\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}_{k\leftarrow K}\right\rangle]|\leq\mathsf{negl}$. Recall the definition of $p_{C^{\prime}\tau}$ and use LABEL:Lem:Cminus11_2EstimateAssumingConsistent to write
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]-\mathsf{negl}\leq\left(1-\frac{1}{|C|}\right)+\sum_{C\in C^{{\rm all}}}\Pr(C)\sum_{q_{\mathsf{skip}}\in C}\frac{1}{|C|}\sum_{\tau}p_{C^{\prime},\tau}\mathsf{pred}(\tau[C],C)$ (44)
where recall that $C^{\prime}:=C\backslash q_{\mathsf{skip}}$. Observe also that there exists $C^{\prime}_{*0}\neq C^{\prime}_{*1}$ such that
$\left\|p_{C^{\prime}_{*0}}-p_{C^{\prime}_{*1}}\right\|_{1}:=\sum_{\tau}\left|p_{C^{\prime}_{*0}\tau}-p_{C^{\prime}_{*1}\tau}\right|\geq\eta$ (45)
for some non-negligible function $\eta$. This is a consequence of the assumption that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}\right\rangle]\geq 1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\epsilon$ (46)
for some non-negligible function $\epsilon$. To see this, suppose for contradiction that for all $C^{\prime}_{*0}\neq C^{\prime}_{*1}$, it were the case that $\sum_{\tau}|p_{C^{\prime}_{*0}\tau}-p_{C^{\prime}_{*1}\tau}|\leq\mathsf{negl}$ for some negligible function, then one could write, using LABEL:Eq:Cminus11_Pr_acc_A_C and $p_{\tau}:=p_{C^{\prime}_{*0}\tau}$
[…]
$=$ $\left(1-\frac{1}{|C|}\right)+\sum_{\tau}p_{\tau}\sum_{C\in C^{{\rm all}}}\Pr(C)\mathsf{pred}(\tau[C],C)+\mathsf{negl}^{\prime}$
$\leq$ $1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\mathsf{negl}^{\prime}.$
But this contradicts LABEL:Eq:Cminus11_A_wins_against_C_w_non-negl and thus LABEL:Eq:Cminus11_p_q_star0tau_minus_p_q_star1tau holds for some $C^{\prime}_{*0}\neq C^{\prime}_{*1}$ as claimed.
The remaining analysis is the same as the $(1,1)$ case, briefly, note that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}_{2},{\cal C}_{2}\right\rangle]=$ $\frac{1}{2}\cdot\sum_{\tau\in T_{0}}\Pr[{\cal A}\text{ outputs }\tau|C^{\prime}_{*0}\text{ was encrypted}]+$
$\frac{1}{2}\cdot\sum_{\tau\in T_{1}}\Pr[{\cal A}\text{ outputs }\tau|C^{\prime}_{*1}\text{ was encrypted}]$
$=$ $\frac{1}{2}+\frac{1}{2}\sum_{\tau\in T_{0}}(p_{C^{\prime}_{*0}\tau}-p_{C^{\prime}_{*1}\tau})$
$\geq$ $\frac{1}{2}+\frac{\eta}{4}.$
Since $\eta$ is non-negligible, existence of a PPT ${\cal A}$ satisfying LABEL:Eq:Cminus11_A_wins_against_C_w_non-negl breaks the security of the underlying $\mathsf{QFHE}$ scheme. The precision issue is handled by invoking LABEL:Lem:Cminus11_3PrecisionNonIssue. This completes the proof. ∎
13.4 Proof of the lemmas (Step 2 of 2)
We only prove LABEL:Lem:Cminus11_2EstimateAssumingConsistent. The proofs of LABEL:Lem:Cminus11_1guessK=00003DgoodK and LABEL:Lem:Cminus11_3PrecisionNonIssue are analogous to those of LABEL:Lem:1guessK=00003DgoodK and LABEL:Lem:3PrecisionNonIssue, respectively.
Proof of LABEL:Lem:Cminus11_2EstimateAssumingConsistent.
Consider the adversary ${\cal A}^{\prime}$ in LABEL:Alg:Cminus11_red_for_lemma2 that uses ${\cal A}$ to interact with ${\cal C}_{k\leftarrow K}=:{\cal C}^{\prime}$ (not to be confused with $C^{\prime}$ which denotes a context with exactly one question removed).
${\cal A}$ ${\cal A}^{\prime}$ ${\cal C}_{k\leftarrow K}=:{\cal C}^{\prime}$
$\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$ $\xleftarrow{(c,\mathsf{OPad}.\mathsf{pk})}$ Recall: $c\leftarrow\mathsf{Enc}_{\mathsf{sk}}(C^{\prime})$ for $C^{\prime}=C\backslash q_{\mathsf{skip}}$.
$\xrightarrow{(c_{\mathbf{a}^{\prime}},\hat{k}^{\prime\prime},s^{\prime})}$
$\tilde{k}\leftarrow K$
$\overset{(\tilde{q},\tilde{k})}{\circlearrowleft}$ Rewinds ${\cal A}^{\prime}$ to learn the truth table $\tau$.
Evaluate $c_{\tau[C^{\prime}]}\leftarrow\mathsf{QFHE}.\mathsf{Eval}_{\tau}(c).$ $\xrightarrow{(c_{\tau(q^{\prime})},\hat{k}^{\prime\prime},s^{\prime})}$
$\xleftarrow{(q,k)}$
$\xrightarrow{a:=\tau(q)}$ Recall: $\mathbf{a}^{\prime}=\mathsf{Dec}_{\mathsf{sk}}(c_{\tau[C^{\prime}]})=\tau[C^{\prime}]$    If $q\in C^{\prime}$,    accept if $a=\mathbf{a}^{\prime}[q]$ If $q\notin C^{\prime}$ (i.e. $q=q_{\mathsf{skip}}$),    accept if ${\rm pred}(\mathbf{a}^{\prime}\cup a,C)=1$    where $\mathbf{a}^{\prime}\cup a$ denotes answers indexed by $C$.
Algorithm 20 ${\cal A}^{\prime}$ uses ${\cal A}$ (a PPT algorithm), to play the compiled contextuality game $\mathsf{G}^{\prime}$. Its winning probability upper bounds that of ${\cal A}$ and can be computed in terms of $p_{C^{\prime}\tau}$ for ${\cal A}$.
We show that
$\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A},{\cal C}^{\prime}\right\rangle]$ $\leq\Pr[\mathsf{accept}\leftarrow\left\langle{\cal A}^{\prime},{\cal C}^{\prime}\right\rangle]$ (47)
$=\left(1-\frac{1}{|C|}\right)+\sum_{C\in C^{{\rm all}}}\Pr(C)\sum_{q_{\mathsf{skip}}\in C}\frac{1}{|C|}\sum_{\tau}p_{C^{\prime},\tau}\mathsf{pred}(\tau[C],C)$ (48)
where recall that $C^{\prime}=C\backslash q_{\mathsf{skip}}$.
Note that for any given $C^{\prime}$, the challenger asks any specific $q$ with probability $1/|C|$, which in particular means that $q=q_{\mathsf{skip}}$ with probability $1/|C|$ and $q\neq q_{\mathsf{skip}}$ with probability $1-1/|C|$.
Let’s derive LABEL:Eq:Cminus11_Aprime_just_as_well. Consider the interactions $\left\langle{\cal A},{\cal C}^{\prime}\right\rangle$ and $\left\langle{\cal A}^{\prime},{\cal C}^{\prime}\right\rangle$. Conditioned on $C^{\prime}$, there are two cases: (1) ${\cal A}$ is consistent, in which case, the answers given by ${\cal A}^{\prime}$ and ${\cal A}$ are identical, or (2) ${\cal A}$ is inconsistent in which case ${\cal A}$ fails with probability at least $1/|C|$ (because the challenger spots the inconsistency with probability at least $1/|C|$), while ${\cal A}^{\prime}$ fails with probability at most $1/|C|$ (because it at most (potentially) fails the predicate evaluation, which happens with probability exactly $1/|C|$).
As for LABEL:Eq:Cminus11_Aprime_at_most, it follows because ${\cal C}^{\prime}$ selects a context $C$ with probability $\Pr(C)$, it asks $q\neq q_{\mathsf{skip}}$ with probability $1-1/|C|$ which corresponds to doing a consistency test—which ${\cal A}^{\prime}$ passes with probability $1$ by construction. Finally, note that $\mathbf{a}^{\prime}\cup a=\tau[C]$. Now, ${\cal C}^{\prime}$ asks $q=q_{\mathsf{skip}}$ with probability $1/|C|$ and in this case, ${\cal A}^{\prime}$ responds with $\tau[C]$ with probability $p_{C^{\prime},\tau}$. Thus, its success probability in this case is the weighted average of $\mathsf{pred}(\tau[C],C)=\mathsf{pred}(\mathbf{a}^{\prime}\cup a,C)$ where the weights are given by $p_{C^{\prime},\tau}$ (recall that $C^{\prime}=C\backslash q_{\mathsf{skip}}$). This completes the proof. ∎